### 10 Apr 2026 v3

# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

Abstract
Classical stochastic-approximation analyses treat the covariance of stochastic gradients as an exogenous modeling input. We show that under exchangeable mini-batch sampling this covariance is identified by the sampling mechanism itself: to leading order it is the projected covariance of per-sample gradients. In well-specified likelihood problems this reduces locally to projected Fisher information; for general M-estimation losses the same object is the projected gradient covariance $G^{\star}(\theta)$, which together with the Hessian induces sandwich/Godambe geometry.
This identification—not the subsequent diffusion or Lyapunov machinery, which is classical once the noise matrix is given—is the paper’s main contribution. It endogenizes the diffusion coefficient (with effective temperature $\tau=\eta/b$), determines the stationary covariance via a Lyapunov equation whose inputs are now structurally fixed, and selects the identified statistical geometry as the natural metric for convergence analysis. We prove matching upper and lower bounds of order $\Theta(1/N)$ for risk in this metric under an oracle budget $N$; the lower bound is established first via a van Trees argument in the parametric Fisher setting and then extended to adaptive oracle transcripts under a predictable-information condition and mild conditional likelihood regularity. Translating these bounds into oracle complexity yields $\varepsilon$-stationarity guarantees in the Fisher dual norm that depend on an intrinsic effective dimension $d_{\mathrm{eff}}$ and a statistical condition number $\kappa_{F}$, rather than ambient dimension or Euclidean conditioning. Numerical experiments confirm the Lyapunov predictions at both continuous-time and discrete-time levels and show that scalar temperature matching cannot reproduce directional noise structure.
Keywords: Stochastic gradient descent; Fisher information; Godambe information; information geometry; diffusion approximation; minimax rates; oracle complexity; batch-size control; simulation optimization
1 Introduction
Stochastic gradient descent (SGD) is the computational workhorse behind a large class of operations research and management science problems, including simulation optimization, large-scale stochastic programming, and data-driven estimation in service and supply-chain systems. In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget. Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and plain SGD frequently behaves as if it were curvature-aware despite using only first-order information.
Classical stochastic approximation (SA) theory provides the analytical machinery—diffusion limits, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—for analyzing such recursions once the gradient-noise covariance is specified. We do not claim novelty for that machinery. Our contribution is logically prior: we identify a structural property of the stochastic-gradient oracle that SA typically treats as exogenous, namely that the mini-batch gradient covariance is pinned down as $b^{-1}G^{\star}(\theta)$ under exchangeable sampling, where $G^{\star}(\theta)$ is the projected covariance of per-sample gradients. In likelihood settings this matrix reduces to Fisher information at the correctly specified point $\theta^{\star}$; for general losses it is the projected gradient covariance that, together with the Hessian $H^{\star}$, induces sandwich/Godambe geometry. Once $G^{\star}(\theta)$ is identified, the diffusion approximation for constant-step SGD is determined rather than modeled; the natural metric for rates is statistical rather than Euclidean; and batch size admits principled design rules under a fixed sampling budget. The subsequent Lyapunov analysis, rate bounds, and complexity guarantees follow by applying standard tools in this identified geometry.
1.1 Motivation: Simulation Optimization as Sampling Control
Consider the canonical simulation-optimization objective
$\min_{\theta\in\Theta}\;L(\theta)\;:=\;\mathbb{E}\big[h(\theta,\xi)\big],$ (1)
where $\xi$ denotes random inputs (arrivals, service times, demand shocks) and $h(\theta,\xi)$ is the induced cost. A standard stochastic-gradient estimator uses $b$ simulation replications,
$g_{B}(\theta)\;=\;\frac{1}{b}\sum_{i=1}^{b}\nabla_{\theta}h(\theta,\xi_{i}),\qquad\xi_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}P,$ (2)
followed by $\theta^{+}=\theta-\eta\,g_{B}(\theta)$. Classical variance-reduction intuition focuses on the scalar variance of a component or of a scalarized performance measure, suggesting that larger $b$ should always help. In practice, under fixed sampling budgets, small $b$ often wins.
The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar. Its covariance is a matrix that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment. Our results show that, to leading order, this covariance scales as $b^{-1}G^{*}(\theta)$, where $G^{*}(\theta)$ is the projected covariance of per-sample gradients. For correctly specified likelihood objectives, $G^{*}(\theta^{\star})$ reduces to projected Fisher information at the well-specified point; for general losses $G^{*}(\theta)$ is the projected gradient covariance that, together with $H^{\star}$, induces sandwich/Godambe geometry. This provides the statistical foundation for interpreting SGD as a controlled diffusion whose effective temperature is $\tau=\eta/b$.
This sampling-control perspective is not limited to simulation optimization. Three canonical MOR problem classes share the same structure: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency. In each case, sampling effort is explicit, $b$ is small relative to $d$, and the question “how many samples per iterate?” is operationally binding. The theory developed below applies uniformly to all three classes.
1.2 Core Insights
Before turning to formal statements, we summarize the main idea informally. When SGD computes a mini-batch gradient by averaging $b$ per-sample gradients, the resulting noise is not generic randomness. Its covariance—the matrix describing which directions are noisy and which are quiet—is determined by the loss function and the data-generating process.
This has three immediate consequences:
  1. 1.
    Mini-batch noise is anisotropic. Doubling the batch size reduces noise equally in all directions, but the shape of the noise ellipsoid is fixed by the identified covariance $G^{\star}$, with Fisher information recovered locally under correct specification.
  2. 2.
    The natural metric for local analysis is statistical, not Euclidean. Since the noise covariance is shaped by $G^{\star}$ near $\theta^{\star}$, measuring convergence in the corresponding statistical metric gives tighter and more interpretable bounds than the Euclidean metric.
  3. 3.
    Batch size is a design variable, not just a hyperparameter. Under a fixed oracle budget $N=Tb$, batch size controls the effective temperature $\tau=\eta/b$ of the induced diffusion. It governs the stationary covariance through the Lyapunov balance, making $b$ a design lever for effort allocation.
The remainder of this section makes these claims precise.
1.3 What We Show
(i) Identification from sampling theory.
Under exchangeable sampling (and simple random mini-batching), the mini-batch gradient satisfies
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mathcal{F}\bigr)\;=\;\frac{1}{b}\,G^{*}(\theta)\;+\;o(b^{-1}),$ (3)
where $G^{*}(\theta)$ is the projected gradient covariance (Section 4). For correctly specified likelihood losses, $G^{*}(\theta^{\star})=F^{\star}(\theta^{\star})$.
(ii) Diffusion and Lyapunov structure.
In the small-stepsize regime with $\tau=\eta/b$ held fixed, the SGD recursion admits a diffusion approximation whose volatility is determined by the identified covariance $G^{*}(\theta)$ (Section 5). Near a nondegenerate critical point, the limiting SDE linearizes to an OU process whose stationary covariance $\Sigma_{\infty}$ solves the Lyapunov equation
$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}\;=\;\tau\,G^{*}(\theta^{\star}),$ (4)
making explicit that steady-state fluctuations are governed by the ratio of curvature to identified noise.
(iii) Matching rates and minimax lower bounds.
We analyze SGD performance in the frozen statistical geometry at $\theta^{\star}$. We establish $1/T$ upper bounds for mean-square error in the Fisher metric and match them with information-theoretic lower bounds via van Trees, showing that the rate and its dependence on intrinsic geometry are tight within the stated model class (Sections 6 and 7).
(iv) Oracle complexity in statistical conditioning.
The oracle complexity bounds scale with a statistical condition number and an effective dimension, isolating statistical difficulty from Euclidean conditioning. This clarifies when classical Hessian-based condition numbers can be misleading for statistical objectives.
(v) Batch size as an OR design variable.
The framework above connects batch-size design to the identified statistical geometry; Section 9 discusses extensions to misspecification, dependence, and operational effort allocation.
Main Results (informal).
Theorem 4.3 — Identification of mini-batch covariance under exchangeable sampling: $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)+o(b^{-1})$. Theorem 5.4 — Diffusion approximation with volatility determined by $G^{\star}$. Thm. 5.6, Cor. 5.7 — OU linearization at $\theta^{\star}$; stationary covariance solves a Lyapunov equation. Thm. 6.5, Prop. 6.10 — Matching $1/(Tb)$ upper and lower bounds in the frozen Fisher metric. Theorem 7.3 — Oracle complexity $N\!=\!\widetilde{\Theta}(\kappa_{F}\,d_{\mathrm{eff}}/\varepsilon^{2})$ for Fisher-dual stationarity.
1.4 Organization
Section 2 reviews related work. Section 3 defines the oracle model and geometric objects. Section 4 derives the covariance identification from exchangeability. Section 5 builds the diffusion approximation, OU linearization, and discrete-time bridge. Section 6 proves matching upper and lower bounds in Fisher geometry, and Section 7 states oracle complexity. Section 8 provides numerical validation, and Section 9 discusses extensions and practical implications.
2 Literature Review and Positioning
We organize prior work along two axes: (i) whether the mini-batch noise covariance is assumed (exogenous) or derived from the sampling mechanism; and (ii) whether rates are stated in a Euclidean or in the identified statistical metric. Table 1 summarizes the resulting taxonomy.
Stochastic approximation and stationary covariance.
Classical SA theory studies recursions $\theta_{t+1}=\theta_{t}-\eta_{t}\{h(\theta_{t})+\xi_{t+1}\}$ and derives a $\sqrt{t}$-CLT whose asymptotic covariance is governed by a Lyapunov equation involving the Jacobian and a long-run noise covariance $\Gamma$ (Robbins and Monro, 1951; Dvoretzky, 1956; Kushner and Yin, 2003; Polyak and Juditsky, 1992). This framework takes $\Gamma$ as given; the OU linearization, Lyapunov balance, and $1/\sqrt{t}$ scaling are standard consequences. Our contribution is logically prior: we derive $\Gamma=G^{\star}(\theta)/b$ from the mini-batch sampling design (Theorem 4.3), converting the SA machinery from a conditional analysis into a structural one.
Diffusion approximations for SGD.
A substantial literature interprets constant-step SGD as a discretization of an SDE with temperature $\tau=\eta/b$ and various proposed diffusion matrices $\mathbf{C}(\theta)$ (Mandt et al., 2017; Li et al., 2017; Smith et al., 2021). In those works the diffusion covariance is either postulated, estimated from running iterates, or justified heuristically through moment matching. Our covariance identification removes this modeling degree of freedom: under exchangeable mini-batching, the diffusion matrix is determined as $G^{\star}(\theta)$ (Theorem 5.4), and the local OU regime inherits its Lyapunov balance at temperature $\tau$ (Corollary 5.7). Because the identification is logically prior to the discretization corrections of Li et al. (2017) or the Hessian-scaling regime of Smith et al. (2021), it is what enables the subsequent rate-matching bounds.
Scaling regimes.
Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2024). Our analysis is complementary and closer to OR operating constraints: we keep $d$ fixed and treat $b$ as a design variable subject to a budget, with particular attention to the common regime $b\ll d$ where the matrix geometry of $G^{\star}(\theta)$ controls transient behavior and risk.
Information geometry, natural gradients, and Fisher-shaped preconditioning.
Information geometry treats Fisher information as a Riemannian metric and motivates natural-gradient updates $\theta_{t+1}=\theta_{t}-\eta\,F(\theta_{t})^{-1}\nabla L(\theta_{t})$ (Amari, 1998; Martens, 2020). Related methods—including stochastic gradient Fisher scoring (Ahn et al., 2012) and KFAC (Martens and Grosse, 2015)—inject or exploit Fisher geometry exogenously in the drift or preconditioner. Our contribution is distinct: even plain SGD, without any preconditioning, inherits the identified statistical geometry endogenously through its noise covariance under mini-batch sampling, changing the interpretation of diffusion strength, stationary risk, and batch-size tuning without requiring natural-gradient updates.
Minimax rates and information-theoretic limits.
Classical oracle lower bounds are stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012), which may not reflect the intrinsic difficulty of statistical objectives. We state both upper and lower bounds in the identified statistical metric via a van Trees argument (Theorem 6.5 and Proposition 6.10), with complexity governed by statistical curvature and $d_{\operatorname{eff}}$ rather than ambient dimension.
OR context: sampling budgets.
In simulation optimization and stochastic programming, gradients are estimated from sampled scenarios and sampling cost is explicit (Fu, 2006; Shapiro et al., 2014; Chernoff, 1959). Because batching controls temperature $\tau=\eta/b$, it directly shapes diffusion amplitude and terminal risk; Section 9 outlines the resulting effort allocation problem.
Why small batches can work when $b\ll d$.
The empirical preference for small batches (Keskar et al., 2017; Shallue et al., 2019; McCandlish et al., 2018) is explained by our framework: smaller $b$ buys more contraction steps under fixed budgets, the identified covariance concentrates noise along statistically flat directions, and rates scale with $d_{\operatorname{eff}}\ll d$.
Table 1: Positioning: primitives, contrasts, and where each appears in the paper.
Primitive Status quo This paper Pointers
Mini-batch noise covariance Exogenous / isotropic / heuristic $\Gamma$ Identified from sampling: $\mathrm{Cov}(g_{B})\propto b^{-1}G^{\star}(\theta)$; Fisher is a special case Sec. 4; Thm. 4.3
Diffusion matrix for SGD Postulated $\mathbf{C}(\theta)$ in SDEs Diffusion coefficient determined by sampling theory: $G^{\star}(\theta)$ Sec. 5; Thm. 5.4
OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU linearization with Lyapunov balance driven by $G^{\star}$ at temperature $\tau=\eta/b$ Sec. 5; Cor. 5.7
Metric for rates Euclidean distance; $\kappa_{H}$ dominates Identified statistical metric is the natural risk Sec. 6; Thm. 6.5
Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.10
Dimension notion Ambient $d$ Effective dimension $d_{\operatorname{eff}}$ induced by $F^{\star}$ / $G^{\star}$ Sec. 6; Def. (36)
Oracle complexity Euclidean stationarity; $\kappa_{H}$ Fisher-dual stationarity; $\kappa_{F}$ and $d_{\operatorname{eff}}$ Sec. 7; Thm. 7.3
3 Problem Setup and Preliminaries
This section fixes notation, specifies the stochastic-gradient oracle, and records the geometric objects that govern the diffusion approximation and the rate/complexity bounds. The perspective is Mathematics of Operations Research: we make the sampling design explicit, separate intrinsic noise geometry from finite-population correction, and state the local regularity conditions used throughout.
3.1 Notation and macros used throughout
We work on a measurable space $(\mathcal{Z},\mathcal{A})$. Vectors use the Euclidean inner product $\langle u,v\rangle=u^{\top}v$ and norm $\|u\|_{2}$. For a symmetric matrix $A$, $\|A\|_{\operatorname{op}}$ denotes the operator norm and $\lambda_{\min}(A),\lambda_{\max}(A)$ its extreme eigenvalues.
3.2 Data model, objective, and stochastic-gradient oracle
Let $\Theta\subseteq\mathbb{R}^{d}$ be open and convex. Let $Q$ denote the (unknown) data distribution on $(\mathcal{Z},\mathcal{A})$, and let $\ell:\Theta\times\mathcal{Z}\to\mathbb{R}$ be a measurable loss. Define the population risk
$L(\theta)\;:=\;\mathbb{E}_{Z\sim Q}\bigl[\ell(\theta;Z)\bigr],$ (5)
and let $\theta^{\star}\in\arg\min_{\theta\in\Theta}L(\theta)$ be a (local) minimizer.
Finite-population sampling model.
We observe a dataset $z_{1:n}:=(z_{1},\dots,z_{n})$ drawn i.i.d. from $Q$ (or, in the exchangeability formulation of Section 4, conditionally i.i.d. given a directing measure). All lower bounds in Section 6 are stated for a general oracle transcript with bounded predictable information (Assumption 6.9); the i.i.d. sampling model is recovered as a special case. At iteration $t$, the algorithm selects a mini-batch $B_{t}\subseteq\{1,\dots,n\}$ of size $b_{t}$ uniformly without replacement and forms the mini-batch gradient
$g_{t}(\theta)\;:=\;\frac{1}{b_{t}}\sum_{i\in B_{t}}\psi(\theta;z_{i}),\qquad\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z).$ (6)
The SGD recursion is
$\theta_{t+1}\;=\;\theta_{t}-\eta_{t}\,g_{t}(\theta_{t}),$ (7)
with stepsize $\eta_{t}>0$.
Filtration and martingale structure.
Let $\mathcal{F}_{t}$ be the $\sigma$-field generated by $(\theta_{0},B_{0},\dots,B_{t-1},\theta_{t})$ and the dataset $z_{1:n}$. Then
$\xi_{t}(\theta_{t}):=g_{t}(\theta_{t})-\mathbb{E}[g_{t}(\theta_{t})\mid\mathcal{F}_{t}]$
is a martingale difference sequence with respect to $(\mathcal{F}_{t})$, and all diffusion limits are driven by the predictable quadratic variation of $\sum_{t}\eta_{t}\xi_{t}$.
Cost accounting (oracle calls).
We measure computational cost in sample-gradient evaluations. One iteration with batch size $b_{t}$ uses $b_{t}$ sample evaluations. After $T$ iterations, the total sampling budget is
$N_{T}\;:=\;\sum_{t=0}^{T-1}b_{t}.$ (8)
We report results both in iterations $T$ and in oracle calls $N_{T}$, since control/OR questions naturally optimize over batch allocation under a fixed $N_{T}$.
3.3 Likelihood losses: Fisher information and the induced metric
When $\ell(\theta;z)=-\log p_{\theta}(z)$ for a parametric family $\{P_{\theta}:\theta\in\Theta\}$ with density $p_{\theta}$, the score is $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ and the loss gradient is the negative score, $\nabla_{\theta}\ell(\theta;z)=-s_{\theta}(z)$. The intrinsic noise geometry reduces to Fisher information at the correctly specified point $\theta^{\star}$.
Definition 3.1 (Fisher information).
Assume $p_{\theta}$ is differentiable in $\theta$ and $s_{\theta}(Z)$ is square integrable under $P_{\theta}$. The Fisher information is
$F(\theta)\;:=\;\mathbb{E}_{\theta}\!\bigl[s_{\theta}(Z)s_{\theta}(Z)^{\top}\bigr].$ (9)
Lemma 3.2 (Score identities (Lehmann and Casella, 1998, Ch. 7)).
Assume $\int p_{\theta}(z)\,dz=1$ and differentiation under the integral is valid. Then (i) $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$ and (ii)
$F(\theta)\;=\;-\mathbb{E}_{\theta}\!\bigl[\nabla_{\theta}^{2}\log p_{\theta}(Z)\bigr].$ (10)
Proof.
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
Fisher metric.
If $F(\theta)\succ 0$, it defines the local inner product $\langle u,v\rangle_{F(\theta)}:=u^{\top}F(\theta)v$ and norm $\|u\|_{F(\theta)}:=\sqrt{u^{\top}F(\theta)u}$. On compact sets where $F(\theta)$ is bounded and bounded away from 0, Fisher and Euclidean norms are equivalent.
3.4 General losses: sandwich / Godambe geometry
For a general loss, information equality need not hold. The relevant local geometry splits into (i) curvature of the drift and (ii) covariance of the estimating equation (gradient).
Definition 3.3 (Noise covariance, sensitivity, and Godambe information).
Let $\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z)$ and define
$G(\theta)$ $:=\mathrm{Cov}\bigl(\psi(\theta;Z)\bigr),$ (11)
$H(\theta)$ $:=\nabla^{2}L(\theta)\;=\;\mathbb{E}\!\left[\nabla_{\theta}\psi(\theta;Z)\right],$ (12)
where expectations are under $Z\sim Q$. When $H(\theta)$ is invertible on the identifiable directions, the classical sandwich covariance and Godambe information are
$\Sigma_{\mathrm{sand}}(\theta):=H(\theta)^{-1}G(\theta)H(\theta)^{-1},\qquad J_{\mathrm{God}}(\theta):=\Sigma_{\mathrm{sand}}(\theta)^{-1}=H(\theta)^{\top}G(\theta)^{-1}H(\theta).$ (13)
Remark 3.4 (Likelihood as a special case).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $\psi=s_{\theta}$ and $G(\theta)=H(\theta)=F(\theta)$ at $\theta^{\star}$, so the sandwich collapses and the Godambe geometry equals Fisher.
3.5 Projection onto identifiable directions
Projection is only needed when the parametrization is redundant or when only a lower-dimensional tangent space is identifiable.
Definition 3.5 (Tangent-space projection and projected geometry).
Let $T_{\theta}\subseteq\mathbb{R}^{d}$ denote the identifiable tangent space at $\theta$ (i.e., the linear span of directions along which the loss varies at first order; formally $T_{\theta}:=\mathrm{range}(H(\theta))$, or equivalently $\mathrm{supp}(F(\theta))$ in the likelihood case), and let $\Pi_{\theta}$ be the Euclidean-orthogonal projector onto $T_{\theta}$. Define the projected noise covariance
$G^{\star}(\theta)\;:=\;\Pi_{\theta}\,G(\theta)\,\Pi_{\theta}^{\top},$ (14)
and, in the likelihood case, the projected Fisher information
$F^{\star}(\theta)\;:=\;\Pi_{\theta}\,F(\theta)\,\Pi_{\theta}^{\top}.$ (15)
When $T_{\theta}=\mathbb{R}^{d}$, $\Pi_{\theta}=I$ and $G^{\star}=G$, $F^{\star}=F$.
Remark 3.6 (Terminological convention: “Fisher/Godambe”).
Throughout the paper we write “Fisher/Godambe” as shorthand for the following identification: in the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ (Fisher information); for general losses, $G^{\star}(\theta^{\star})$ is the projected gradient covariance whose inverse is the Godambe information. Every theorem statement holds with $G^{\star}$ in the general case and specializes to $F^{\star}$ under correct specification. When we write “Fisher metric” or “Fisher-dual norm,” the general-loss reader should substitute the corresponding Godambe/sandwich object.
Remark 3.7 (What enters later sections).
Sections 45 show that the mini-batch noise covariance is (up to a sampling-design factor) exactly $G^{\star}(\theta)/b_{t}$ and reduces to $F^{\star}(\theta)/b_{t}$ under correct specification. This is the only matrix that enters the diffusion coefficient and Lyapunov balance.
3.6 Local regularity conditions
All results are local near $\theta^{\star}$ and use standard smoothness/moment and nondegeneracy conditions.
Assumption 3.8 (Smoothness and moments).
The loss $\ell(\theta;z)$ is twice continuously differentiable in $\theta$. There exist constants $M<\infty$ and $m_{4}<\infty$ such that for all $\theta$ in a neighborhood $\mathcal{N}$ of $\theta^{\star}$:
  1. (a)
    $\|\nabla_{\theta}^{2}\ell(\theta;z)\|_{\operatorname{op}}\leq M$ almost surely;
  2. (b)
    $\mathbb{E}\bigl[\|\psi(\theta;Z)\|_{2}^{4}\bigr]\leq m_{4}$.
Assumption 3.9 (Local stability).
The population risk is locally strongly convex at $\theta^{\star}$:
$H^{\star}:=\nabla^{2}L(\theta^{\star})\succeq\mu I\quad\text{for some }\mu>0.$ (16)
Assumption 3.10 (Geometric nondegeneracy).
The intrinsic noise geometry is positive definite on identifiable directions:
$G^{\star}(\theta^{\star})\succeq\lambda_{\min}I\quad\text{for some }\lambda_{\min}>0.$ (17)
(In the correctly specified likelihood case this is $F^{\star}(\theta^{\star})\succ 0$.)
Remark 3.11 (Role of assumptions).
Assumptions 3.83.10 provide Taylor/moment control, a locally stable equilibrium, and nondegeneracy of the diffusion coefficient, respectively.
3.7 Key geometric parameters (preview)
Several quantities derived from the local geometry at $\theta^{\star}$ govern the rate and complexity results of Sections 67. We collect their definitions here so the reader can refer to them before encountering the formal theorems.
Definition 3.12 (Fisher-strong-convexity constant).
Let $H^{\star}:=\nabla^{2}L(\theta^{\star})$ and $F^{\star}:=F(\theta^{\star})$ (or $G^{\star}(\theta^{\star})$ for general losses). Define
$\mu_{F}\;:=\;\lambda_{\min}\!\left(\frac{(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}+(F^{\star})^{-1/2}H^{\star}(F^{\star})^{1/2}}{2}\right).$ (18)
Equivalently, $\mu_{F}$ is the smallest eigenvalue of the symmetrized curvature–information interaction $\mathrm{Sym}((F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2})$.
Definition 3.13 (Fisher condition number and effective dimension).
$\kappa_{F}\;:=\;\frac{\lambda_{\max}(F^{\star})}{\lambda_{\min}(F^{\star})},\qquad d_{\mathrm{eff}}(F^{\star})\;:=\;\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}\;\in\;(0,d].$ (19)
The quantity $\kappa_{F}$ measures statistical anisotropy rather than Euclidean ill-conditioning; a problem can be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and information share eigenstructure. The effective dimension $d_{\mathrm{eff}}$ is the stable rank of $F^{\star}$: it equals $d$ when $F^{\star}$ is isotropic and approaches $1$ when a single direction dominates.
Remark 3.14 (Generalized eigenvalue interpretation).
The constant $\mu_{F}$ is linked to the generalized eigenvalues of the pair $(H^{\star},F^{\star})$. If $H^{\star}$ and $F^{\star}$ commute (i.e., share an eigenbasis $\{v_{j}\}$ with respective eigenvalues $h_{j}$ and $f_{j}$), then $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$ has eigenvalues $f_{j}^{1/2}\,h_{j}\,f_{j}^{-1/2}=h_{j}$, so the symmetrization is trivial and $\mu_{F}=\min_{j}h_{j}=\lambda_{\min}(H^{\star})$: the Fisher-strong-convexity constant reduces to the ordinary strong-convexity constant. In the non-commuting case, $(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}$ is non-normal (its eigenvalues are still those of $H^{\star}$, but it is no longer symmetric), and the minimum of its quadratic form—i.e., $\mu_{F}$—can drop strictly below $\lambda_{\min}(H^{\star})$. Geometrically, this occurs when the Fisher metric amplifies directions along which curvature is weak, making the effective contraction rate slower than the Euclidean strong-convexity constant. The product $\kappa_{F}\cdot d_{\mathrm{eff}}$ that enters the oracle complexity (Theorem 7.3) therefore measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale.
Table 2: Key notation. All quantities are evaluated locally at or near $\theta^{\star}$ unless noted otherwise.
Symbol Meaning
$L(\theta)=\mathbb{E}[\ell(\theta;Z)]$ Population risk (loss)
$\psi(\theta;z)=\nabla_{\theta}\ell(\theta;z)$ Per-sample estimating equation (gradient)
$g_{B}(\theta)$ Mini-batch gradient estimator
$G(\theta)=\mathrm{Cov}(\psi(\theta;Z))$ Gradient covariance; $G^{\star}=\Pi G\Pi^{\top}$ projected
$F(\theta)=\mathbb{E}[s_{\theta}s_{\theta}^{\top}]$ Fisher information (likelihood case)
$F^{\star}(\theta^{\star})$ Projected Fisher; equals $G^{\star}(\theta^{\star})$ under correct specification
$H^{\star}=\nabla^{2}L(\theta^{\star})$ Hessian at optimum (drift curvature)
$J_{\mathrm{God}}=H^{\top}G^{-1}H$ Godambe (sandwich) information
$\tau=\eta/b$ Effective temperature
$\mu_{F}$ Fisher-strong-convexity constant (Def. 3.12)
$\kappa_{F}$ Fisher condition number (Def. 3.13)
$d_{\operatorname{eff}}(F^{\star})$ Effective dimension / stable rank of $F^{\star}$
4 Fisher/Godambe Alignment of Mini-Batch Noise
Mini-batch gradient noise is not an arbitrary covariance: it is determined by (i) the sampling design used to form mini-batches and (ii) the intrinsic covariance of the per-sample estimating equation $\psi(\theta;Z)$ (Fisher information in the well-specified likelihood case, the projected gradient covariance more generally). We make this precise in two steps:
  • an exact finite-population identity conditional on the realized dataset (classical survey-sampling algebra), and
  • an exchangeable/i.i.d. limit that converts the dataset-dependent quantity into the population covariance $G(\theta):=\mathrm{Cov}(\psi(\theta;Z))$ (or its projected analogue).
4.1 Finite-population covariance identity (exact, conditional on data)
Fix $\theta$ and condition on the realized dataset $\{z_{1},\dots,z_{n}\}$. Write $\psi_{i}:=\psi(\theta;z_{i})$ and define the finite-population mean and covariance
$\bar{\psi}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\psi_{i},\qquad S_{n}(\theta):=\frac{1}{n-1}\sum_{i=1}^{n}(\psi_{i}-\bar{\psi}_{n})(\psi_{i}-\bar{\psi}_{n})^{\top}.$
Let $B$ be a simple random sample without replacement of size $b$ from $\{1,\dots,n\}$ and define the mini-batch average
$g_{B}(\theta):=\frac{1}{b}\sum_{i\in B}\psi_{i}.$
Lemma 4.1 (Exact finite-population covariance of the mini-batch mean).
Conditionally on the dataset,
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid z_{1:n}\bigr)\;=\;\frac{1}{b}\Bigl(1-\frac{b}{n}\Bigr)\,S_{n}(\theta).$ (20)
Proof.
Write $g_{B}(\theta)=\frac{1}{b}\sum_{i=1}^{n}I_{i}\,\psi_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$. Under simple random sampling without replacement, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and for $i\neq j$, $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$. Expanding $\mathrm{Cov}(\sum_{i}I_{i}\psi_{i}\mid z_{1:n})$ and collecting terms yields (20). ∎
Remark 4.2 (Equivalent correction-factor forms).
Using $S_{n}$ with the 1/(n1)1/(n-1) normalization gives the factor (1b/n)(1-b/n). If instead one uses the $1/n$ covariance normalization, the factor becomes 1(b1)/(n1)1-(b-1)/(n-1). The two are algebraically equivalent up to the change of normalization.
Lemma 4.1 is the clean MOR-style decomposition: sampling design contributes the scalar finite-population correction 1b/n1-b/n, while the geometry lives in $S_{n}(\theta)$.
4.2 From exchangeability to Godambe/Fisher alignment (population geometry)
We now convert the dataset-dependent covariance $S_{n}(\theta)$ into a population-level object.
Let $(Z_{i})_{i\geq 1}$ be exchangeable with de Finetti directing measure $\mu$ (Hewitt and Savage, 1955). Conditional on $\mu$, the sequence is i.i.d. with law $P_{\mu}$. Define the population covariance (conditional on $\mu$)
$G_{\mu}(\theta)\;:=\;\mathrm{Cov}_{\mu}\bigl(\psi(\theta;Z)\bigr),\qquad Z\sim P_{\mu},$ (21)
and, when identifiability requires it, the projected version $G_{\mu}^{\star}(\theta):=\Pi_{\theta}G_{\mu}(\theta)\Pi_{\theta}^{\top}$.
Theorem 4.3 (Godambe alignment under exchangeability).
Assume $\mathbb{E}_{\mu}\|\psi(\theta;Z)\|_{2}^{2}<\infty$. Let $g_{B}(\theta)=\frac{1}{b}\sum_{i\in B}\psi(\theta;Z_{i})$ where $B$ is a uniform subset of $\{1,\dots,n\}$ of size $b$ (sampling without replacement). Then, conditional on $\mu$,
$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mu\bigr)\;=\;\frac{1}{b}\,G_{\mu}(\theta).$ (22)
Interpretation. Equation (22) is exact (not an approximation) when one marginalizes over both the mini-batch index $B$ and the dataset realization conditional on $\mu$. The finite-population correction (1b/n)(1-b/n) of Lemma 4.1 is present when one conditions on a fixed dataset, but it disappears at the population level because the randomness of the dataset itself contributes a complementary term through the total-covariance decomposition. Equation (23) below makes this bookkeeping explicit.
Moreover,
$\mathbb{E}\!\Big[\mathrm{Cov}(g_{B}(\theta)\mid Z_{1:n},\mu)\,\big|\,\mu\Big]\;=\;\frac{1}{b}\Bigl(1-\frac{b}{n}\Bigr)\,\mathbb{E}[S_{n}(\theta)\mid\mu]\;=\;\frac{1}{b}\Bigl(1-\frac{b}{n}\Bigr)\,G_{\mu}(\theta),$ (23)
and the remaining part of (22) comes from the randomness of the dataset through the total-covariance decomposition.
Proof.
Conditional on $\mu$, the sequence $(Z_{i})_{i\geq 1}$ is i.i.d. with law $P_{\mu}$; therefore any uniformly sampled subset of size $b$ has the same joint distribution as the first $b$ coordinates. Hence selecting indices $B$ without replacement from $\{1,\dots,n\}$ is distributionally equivalent to taking the first $b$ terms, and $g_{B}(\theta)$ is the average of $b$ i.i.d. draws from $P_{\mu}$. This yields $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=\frac{1}{b}G_{\mu}(\theta)$, proving (22).
For (23), apply the exact identity Lemma 4.1 conditional on $(Z_{1:n},\mu)$ and take $\mathbb{E}[\cdot\mid\mu]$. Under $\mathbb{E}_{\mu}\|\psi\|^{2}<\infty$, $\mathbb{E}[S_{n}(\theta)\mid\mu]=G_{\mu}(\theta)$ (the 1/(n1)1/(n-1) sample-covariance normalization is unbiased). ∎
Remark 4.4 (What “alignment” means operationally).
Equation (22) is the alignment statement used by the diffusion/Lyapunov analysis: mini-batch noise has covariance $\frac{1}{b}$ times a population matrix determined by the loss and the data law. The finite-population correction in Lemma 4.1 matters when one conditions on a fixed dataset and reuses it many times; in the standard stochastic-approximation regime with fresh randomness across iterations, the population covariance (22) is the relevant object.
Proposition 4.5 (Finite-population correction under dataset reuse).
Under the notation of Theorem 4.3, three conditioning levels yield distinct covariance expressions:
  1. (i)
    Conditional on the dataset $z_{1:n}$ (fixed pool, sampling without replacement): $\mathrm{Cov}(g_{B}(\theta)\mid z_{1:n})=\frac{1}{b}(1-\frac{b}{n})\,S_{n}(\theta)$ (Lemma 4.1).
  2. (ii)
    Conditional on the directing measure $\mu$ (fresh data per iteration): $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=\frac{1}{b}\,G_{\mu}(\theta)$ (Theorem 4.3), with no finite-population correction.
  3. (iii)
    Population level (unconditional): $\mathrm{Cov}(g_{B}(\theta))=\frac{1}{b}\,\mathbb{E}_{\mu}[G_{\mu}(\theta)]+\mathrm{Cov}_{\mu}(\mathbb{E}[g_{B}(\theta)\mid\mu])$. This reduces to $\frac{1}{b}\,G(\theta)$ when the conditional mean $\mathbb{E}[\psi(\theta;Z)\mid\mu]$ does not vary with $\mu$ (e.g., under identifiability with a single data-generating distribution).
When a finite dataset is reused across many iterations (i.e., without-replacement sampling from the same $z_{1:n}$ at every step), the conditional covariance carries the factor (1b/n)(1-b/n), which is non-negligible unless $b=o(n)$. The main diffusion, rate, and complexity results (Sections 57) are population-level or fresh-sampling statements that use level (ii).
Proof.
Level (i) is Lemma 4.1. Level (ii) is Theorem 4.3. Level (iii) follows by integrating level (ii) over $\mu$ under identifiability. The total-covariance decomposition $\mathrm{Cov}(g_{B})=\mathbb{E}_{\mu}[\mathrm{Cov}(g_{B}\mid\mu)]+\mathrm{Cov}_{\mu}(\mathbb{E}[g_{B}\mid\mu])$ reconciles levels (ii) and (iii), with the second term contributing when $\mathbb{E}[g_{B}\mid\mu]$ varies with $\mu$. ∎
Remark 4.6 (Batch size rescales temperature, not geometry).
Equation (22) implies that increasing the batch size $b$ rescales the effective temperature $\tau=\eta/b$ (Section 5) but does not alter the noise ellipsoid: its shape is entirely determined by the statistical geometry $G_{\mu}(\theta)$. This decoupling—scale from the sampling budget, shape from the loss—underlies the subsequent rate and complexity results.
[…]
Figure 1: Isotropic noise (left) vs. Fisher-aligned noise (right) at the same total variance. Batch size rescales temperature but not shape.
Remark 4.7 (Specialization to Fisher alignment).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $\psi(\theta;z)=s_{\theta}(z)$ is the score and the Bartlett identity yields $\mathrm{Cov}(s_{\theta}(Z))=F(\theta)$ at $\theta=\theta^{\star}$. Projecting onto the identifiable tangent space gives $G_{\mu}^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$. Thus (22) becomes $\mathrm{Cov}(g_{B}(\theta^{\star})\mid\mu)=\frac{1}{b}F^{\star}(\theta^{\star})$, i.e., Fisher alignment.
Remark 4.8 (No Gaussian assumption).
Neither Lemma 4.1 nor Theorem 4.3 assumes Gaussian noise. Gaussianity enters later only as a diffusion approximation (martingale CLT) when we pass from the discrete recursion to an OU limit.
5 Continuous-Time Diffusion Approximation and the OU Regime
This section links the discrete-time SGD recursion to a continuous-time SDE whose diffusion covariance is determined by the mini-batch sampling model and the intrinsic noise geometry derived in Section 4. Specifically, $G^{\star}(\theta)$ was defined in Section 3 and shown in Theorem 4.3 to govern the leading-order mini-batch covariance. Near a nondegenerate critical point, the SDE linearizes to an Ornstein–Uhlenbeck (OU) process, and its stationary covariance is characterized by a Lyapunov equation. The resulting “temperature” parameter $\tau=\eta/b$ is the scalar that converts sampling effort into diffusion strength. Diffusion approximations are a standard tool in OR for reducing complex stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
The stationary covariance and risk expressions derived below depend only on the local second-moment structure of the oracle noise and do not require independence across oracle calls; a martingale-difference oracle (Assumption 6.9) suffices.
5.1 SGD decomposition and effective temperature
Fix constant stepsize $\eta>0$ and batch size $b$. Let
$h(\theta):=\nabla L(\theta)=\mathbb{E}[\psi(\theta;Z)],\qquad\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z),$
and write the mini-batch gradient estimator as
$g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t}(\theta_{t}),\qquad\mathbb{E}[\xi_{t}(\theta_{t})\mid\mathcal{F}_{t}]=0,$
where $\mathcal{F}_{t}$ is the natural filtration. Then SGD is
$\theta_{t+1}=\theta_{t}-\eta\,h(\theta_{t})-\eta\,\xi_{t}(\theta_{t}).$ (24)
By Theorem 4.3 (and its Fisher specialization in Remark 4.7), the conditional covariance of the noise obeys
$\mathrm{Cov}\!\bigl(\xi_{t}(\theta)\mid\mathcal{F}_{t}\bigr)=\frac{1}{b}\,G^{\star}(\theta)+r_{b,n}(\theta),\qquad\|r_{b,n}(\theta)\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1}),$ (25)
uniformly in a local neighborhood (the $O(n^{-1})$ term accounts for finite-population corrections and covariance-estimation bias; it is negligible whenever $b=o(n)$).
We call $\tau:=\eta/b$ the effective temperature:
$\tau:=\eta/b,$ (26)
which will appear as the single scalar factor multiplying the diffusion covariance in the continuous-time limit.
5.2 A diffusion approximation (Euler–Maruyama viewpoint)
Let $s=t\eta$ denote algorithmic time and define the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ and linearity on $t\eta,(t+1)\eta]$.
Assumption 5.1 (Local regularity and localization).
On a neighborhood $\mathcal{U}$ of $\theta^{\star}$:
  1. (i)
    $h(\theta):=\nabla L(\theta)$ is locally Lipschitz;
  2. (ii)
    $\nabla^{2}L(\theta)$ is locally Lipschitz (i.e., there exists $L_{H}<\infty$ such that $\|\nabla^{2}L(\theta)-\nabla^{2}L(\theta^{\prime})\|_{\operatorname{op}}\leq L_{H}\|\theta-\theta^{\prime}\|_{2}$ for all $\theta,\theta^{\prime}\in\mathcal{U}$);
  3. (iii)
    $\sup_{\theta\in\mathcal{U}}\mathbb{E}\|\psi(\theta;Z)\|^{4}<\infty$ (bounded fourth moment; implied by Assumption 3.8(b));
  4. (iv)
    $H^{\star}:=\nabla^{2}L(\theta^{\star})\succ 0$ (local strong convexity; Assumption 3.9);
  5. (v)
    for each fixed horizon $T>0$, $\Pr(\theta_{t}\in\mathcal{U}\ \forall\,t\leq T/\eta)\to 1$ as $\eta\downarrow 0$ (localization; standard in stochastic approximation).
Assumption 5.2 (Asymptotic conditional covariance and Lindeberg).
Uniformly on $\mathcal{U}$, $\mathrm{Cov}(\xi_{t}(\theta)\mid\mathcal{F}_{t})=b^{-1}G^{\star}(\theta)+o(b^{-1})$ and the conditional Lindeberg condition holds for the martingale increments $\eta\xi_{t}(\theta_{t})$ (e.g., via a conditional CLT under exchangeability as in Section 4).
Remark 5.3 (Verifiable sufficient conditions).
Assumption 5.2 holds if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|_{2}^{2+\delta}]<\infty$ for some $\delta>0$ (guaranteed by Assumption 3.8(b)), $G^{\star}(\theta)$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence.
Theorem 5.4 (Diffusion approximation).
Suppose Assumptions 5.15.2 hold. For fixed batch size $b\geq 1$ and constant step size $\eta>0$, define the effective temperature $\tau:=\eta/b$. For sufficiently small $\eta$, the SGD recursion admits a diffusion approximation: the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ converges weakly on $D([0,T];\mathbb{R}^{d})$ as $\eta\downarrow 0$ to the solution $\theta(\cdot)$ of the SDE
$d\theta_{s}=-h(\theta_{s})\,ds+\sqrt{\tau}\,C^{\star}(\theta_{s})\,dW_{s},\qquad C^{\star}(\theta)\,{C^{\star}(\theta)}^{\top}=G^{\star}(\theta),$ (27)
where $C^{\star}$ is, e.g., the symmetric PSD square root of $G^{\star}$,
where $W_{s}$ is standard $d$-dimensional Brownian motion. In particular, the diffusion has instantaneous covariance $\tau\,G^{\star}(\theta_{s})\,ds$. In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ at $\theta^{\star}$, so the diffusion covariance is locally Fisher-structured near the optimum.
Proof.
Step 1 (scaled interpolation). Write the recursion in signed-noise form $\theta_{t+1}=\theta_{t}-\eta h(\theta_{t})+\eta\,\zeta_{t+1}$ with $\zeta_{t+1}:=-\xi_{t}$, so $(\zeta_{t+1},\mathcal{F}_{t+1})$ is a martingale difference. Define the piecewise-linear interpolation $\bar{\theta}^{\eta}(s)$ on algorithmic time $s=t\eta$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ with linear interpolation.
Step 2 (drift convergence). By local Lipschitzness of $h$ (Assumption 5.1) and the localization condition, $\sum_{k<s/\eta}\eta\,h(\theta_{k})\to\int_{0}^{s}h(\theta(u))\,du$ uniformly on compacts, by standard ODE-approximation theory (Kushner and Yin, 2003, Chapter 5).
Step 3 (predictable quadratic variation). Define $M^{\eta}(s):=\sum_{k<s/\eta}\eta\,\zeta_{k+1}$. By the alignment theorem (Theorem 4.3), $\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\frac{1}{b}\,G^{\star}(\theta_{k})+R_{k}$ with $\|R_{k}\|_{\operatorname{op}}\leq(c/b)\rho_{k}$ and $\rho_{k}\to 0$ locally. Using $\eta^{2}/b=\eta\tau$:
$\langle M^{\eta}\rangle(s)=\sum_{k<s/\eta}\eta^{2}\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\tau\sum_{k<s/\eta}\eta\,G^{\star}(\theta_{k})+o(1)\;\to\;\tau\!\int_{0}^{s}\!G^{\star}(\theta(u))\,du$
in probability, uniformly on $[0,S]$.
Step 4 (Lindeberg and tightness). The fourth moment bound (Assumption 3.8(b)) yields $\sum_{k<S/\eta}\mathbb{E}[\|\eta\zeta_{k+1}\|^{2}\mathbf{1}\{\|\eta\zeta_{k+1}\|>\varepsilon\}\mid\mathcal{F}_{k}]\leq CS\eta^{3}m_{4}/(\varepsilon^{2}b^{2})\to 0$, verifying the conditional Lindeberg condition. Tightness of $\{\bar{\theta}^{\eta}\}$ in $D([0,S];\mathbb{R}^{d})$ follows from Aldous’ criterion and the Lipschitz-controlled drift.
Step 5 (martingale FCLT). By the martingale central limit theorem (Ethier and Kurtz, 1986, Theorem 7.1.4), $M^{\eta}(\cdot)\Rightarrow M(\cdot)$ weakly in $D([0,S];\mathbb{R}^{d})$ where $M$ is a continuous martingale with quadratic variation $\tau\int_{0}^{s}G^{\star}(\theta(u))\,du$. Lévy’s characterization gives a Brownian motion $W$ such that $M(s)=\int_{0}^{s}\sqrt{\tau}\,C^{\star}(\theta(u))\,dW_{u}$. Combining drift and martingale convergence identifies the limit as the unique solution of (27). ∎
5.3 The OU regime near a nondegenerate critical point
Assume $\theta^{\star}$ is a stationary point: $h(\theta^{\star})=0$, and define the local curvature $H^{\star}:=\nabla h(\theta^{\star})=\nabla^{2}L(\theta^{\star})$. Let $\Delta_{s}:=\theta_{s}-\theta^{\star}$ denote the centered diffusion.
Assumption 5.5 (Nondegeneracy at $\theta^{\star}$).
$H^{\star}\succ 0$ and $G^{\star}(\theta^{\star})\succ 0$.
Theorem 5.6 (OU linearization).
Under Assumption 5.5 and the conditions of Theorem 5.4, the diffusion (27) admits the local expansion
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s}+\rho_{1}(\Delta_{s})\,ds+\sqrt{\tau}\,\rho_{2}(\Delta_{s})\,dW_{s},$ (28)
where $\|\rho_{1}(\Delta)\|=o(\|\Delta\|)$ and $\|\rho_{2}(\Delta)\|_{\operatorname{op}}=o(1)$ as $\Delta\to 0$. Neglecting the higher-order remainder terms yields the OU SDE
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s},\qquad C^{\star}(\theta^{\star}){C^{\star}(\theta^{\star})}^{\top}=G^{\star}(\theta^{\star}).$ (29)
Proof.
A Taylor expansion gives $h(\theta^{\star}+\Delta)=H^{\star}\Delta+r(\Delta)$ with $\|r(\Delta)\|=o(\|\Delta\|)$. Continuity of $G^{\star}$ implies we can choose a continuous square root $C^{\star}(\theta)$ locally (e.g., the symmetric PSD square root), so
$C^{\star}(\theta^{\star}+\Delta)=C^{\star}(\theta^{\star})+R_{C}(\Delta),\qquad\|R_{C}(\Delta)\|_{\operatorname{op}}=o(1)\quad(\Delta\to 0).$
Substituting into (27) and collecting terms yields (28) with $\rho_{1}(\Delta):=-r(\Delta)$ and $\rho_{2}(\Delta):=R_{C}(\Delta)$. ∎
Corollary 5.7 (Stationary covariance and the Lyapunov equation).
If $H^{\star}\succ 0$, the OU process (29) is ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$. Since $H^{\star}\succ 0$, the matrix $H^{\star}$ is Hurwitz, so the continuous Lyapunov equation
$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}=\tau\,G^{\star}(\theta^{\star})$ (30)
admits a unique symmetric positive definite solution $\Sigma_{\infty}$. In particular, the stationary covariance scales linearly with temperature $\tau=\eta/b$.
Because $H^{\star}$ is symmetric positive definite, left-multiplying (30) by $(H^{\star})^{-1}$ and taking traces gives the trace identity
$\mathrm{Tr}(\Sigma_{\infty})+\mathrm{Tr}\!\bigl((H^{\star})^{-1}\Sigma_{\infty}H^{\star}\bigr)=\tau\,\mathrm{Tr}\!\bigl((H^{\star})^{-1}G^{\star}(\theta^{\star})\bigr),$ (31)
which reduces to $2\,\mathrm{Tr}(\Sigma_{\infty})=\tau\,\mathrm{Tr}(G^{\star}H^{\star-1})$ when $H^{\star}$ and $\Sigma_{\infty}$ commute (e.g., when the eigenbases of curvature and noise align).
Proof.
Write $B:=\sqrt{\tau}\,(G^{\star}(\theta^{\star}))^{1/2}$, so $BB^{\top}=\tau G^{\star}(\theta^{\star})$. Since $^{\star}\succ 0, H-H^{\star}$ is Hurwitz and the OU process is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$. Apply Itô’s formula to $\Delta_{s}\Delta_{s}^{\top}$: the drift contributes $(H^{\star}\Sigma_{s}+\Sigma_{s}H^{\star\top})$ and the quadratic variation contributes $BB^{\top}$, giving $\frac{d}{ds}\Sigma_{s}=-H^{\star}\Sigma_{s}-\Sigma_{s}H^{\star\top}+BB^{\top}$. Setting $\frac{d}{ds}\Sigma_{s}=0$ yields (30). Uniqueness follows because for Hurwitz $H^{\star}$ the integral $\Sigma_{\infty}=\int_{0}^{\infty}e^{-H^{\star}s}\,(\tau G^{\star})\,e^{-H^{\star\top}s}\,ds$ converges and is the unique solution (Lemma A.3). ∎
Remark 5.8 (Likelihood case (Fisher-aligned diffusion)).
If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ and the OU diffusion reads
$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C_{F}^{\star}\,dW_{s},\qquad C_{F}^{\star}{C_{F}^{\star}}^{\top}=F^{\star}(\theta^{\star}),$
making Fisher geometry explicit in both transient fluctuations and in the stationary Lyapunov balance (30).
5.4 Discrete-time stationary covariance: an exact bridge
The OU viewpoint is not the only consequence of the covariance identification. The same $G^{\star}$ governs the exact linearized discrete-time stationary covariance; the continuous-time Lyapunov equation emerges as the small-stepsize asymptotic reduction. This subsection makes the bridge explicit, so that the paper’s conclusions do not rest on the diffusion approximation alone.
Proposition 5.9 (Discrete-time linearized stationary covariance).
Consider the linearized discrete-time recursion around $\theta^{\star}$ with constant stepsize $\eta$ and batch size $b$:
$\Delta_{t+1}=(I-\eta H^{\star})\,\Delta_{t}+\eta\,\zeta_{t+1},\qquad\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]=\frac{1}{b}\,G^{\star}(\theta^{\star})+r_{t},$ (32)
where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1})$ as in (25). If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), then the iterates are geometrically ergodic and the stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfies the discrete Lyapunov equation
$\Sigma_{\eta}=(I-\eta H^{\star})\,\Sigma_{\eta}\,(I-\eta H^{\star})^{\top}+\frac{\eta^{2}}{b}\,G^{\star}(\theta^{\star}),$ (33)
up to the same local remainder terms already used in the diffusion analysis.
Proof.
From (32), compute $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$, using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$. Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the iteration converges to the unique fixed point (33). The spectral radius condition $\rho(I-\eta H^{\star})<1$ is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$. ∎
Remark 5.10 (Continuous-time reduction and discrete-time corrections).
Expanding (33) and dividing by $\eta$, one recovers $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=\frac{\eta}{b}G^{\star}+\eta H^{\star}\Sigma_{\eta}H^{\star\top}$. Sending $\eta\downarrow 0$ with $\tau=\eta/b$ fixed drops the $O(\eta)$ correction term and yields the continuous Lyapunov equation (30). The correction is governed by $\eta\lambda_{\max}(H^{\star})$: when this product is small the two benchmarks nearly coincide; for moderate stepsizes the discrete equation is the correct comparator (Section 8). The covariance identification $\mathrm{Cov}(\xi_{t})\approx b^{-1}G^{\star}$ is logically prior to both analyses: the discrete Lyapunov equation is its exact linearized consequence, and the OU equation is the small-$\eta$ reduction.
6 Convergence Rates in Fisher Geometry
This section establishes matching upper and lower bounds for SGD when the error is measured in a frozen Fisher metric at a local optimum $\theta^{\star}$. Once the mini-batch noise covariance is identified (Section 4), the remaining work is a localized stochastic-approximation argument in a fixed geometry, followed by an information-theoretic lower bound expressed in the same metric. Convention. All rates in this section are stated in Fisher/Godambe risk $\mathbb{E}[\theta_{T}^{\top}F^{\star}\theta_{T}]$ unless explicitly noted otherwise; any Euclidean norm that appears is immediately converted.
6.1 Local Fisher Geometry and a Frozen-Metric Regime
Work in a neighborhood of $\theta^{\star}$ where the Fisher information varies slowly. For clean rate statements we freeze the metric
$F^{\star}:=F(\theta^{\star}),\qquad\|v\|_{F^{\star}}^{2}:=v^{\top}F^{\star}v.$
(Allowing a slowly varying metric $F(\theta)$ is standard via localization; the frozen-metric bounds below are the local core and are what we use later for oracle complexity.)
Lemma 6.1 (Sufficient condition for the frozen-metric regime).
Suppose $F(\theta)$ is Lipschitz on a neighborhood $\mathcal{U}$ of $\theta^{\star}$: $\|F(\theta)-F(\theta^{\star})\|_{\operatorname{op}}\leq L_{F}\|\theta-\theta^{\star}\|_{2}$ for all $\theta\in\mathcal{U}$. If the SGD iterates satisfy $\|\theta_{t}-\theta^{\star}\|_{2}=o(1)$ almost surely, then for every $\epsilon>0$ there exists $t_{0}$ such that for $t\geq t_{0}$, $(1-\epsilon)F^{\star}\preceq F(\theta_{t})\preceq(1+\epsilon)F^{\star}$.
Let $h(\theta):=\nabla L(\theta)$ and decompose the mini-batch estimator as
$g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t},\qquad\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0,$
where $\mathcal{F}_{t}$ is the natural filtration. Under Fisher alignment (Section 4), the conditional covariance satisfies the local approximation
$\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\;=\;\frac{1}{b}\,F^{\star}\;+\;r_{t},\qquad\|r_{t}\|_{\operatorname{op}}\leq\frac{c_{r}}{b}\,\rho_{t},$ (34)
with $\rho_{t}\to 0$ in the local regime.
Define the Fisher-strong-convexity constant at $\theta^{\star}$:
$\mu_{F}\;:=\;\lambda_{\min}\!\left(\frac{(F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}+(F^{\star})^{-1/2}H^{\star}(F^{\star})^{1/2}}{2}\right),\qquad H^{\star}:=\nabla^{2}L(\theta^{\star}).$ (35)
Assumption 6.2 (Local coercivity in the frozen Fisher geometry).
The constant $\mu_{F}$ defined in (35) is strictly positive. This is not automatic from $H^{\star}\succ 0$ and $F^{\star}\succ 0$ alone; it is an explicit compatibility condition between the loss curvature and noise geometry. It holds whenever $H^{\star}$ and $F^{\star}$ are jointly diagonalizable, and more generally whenever $H^{\star}$ and $F^{\star}$ are sufficiently aligned.
Define the effective dimension
$d_{\operatorname{eff}}(F^{\star})\;:=\;\frac{\mathrm{Tr}(F^{\star})}{\|F^{\star}\|_{\operatorname{op}}}\;=\;\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}\;\in(0,d].$ (36)
Remark 6.3 (Effective dimension as stable rank).
The quantity $d_{\operatorname{eff}}(F^{\star})$ is exactly the stable rank of $F^{\star}$ (Tropp, 2015): $d_{\operatorname{eff}}\leq d$ always, with equality iff $F^{\star}$ is isotropic, and $d_{\operatorname{eff}}\to 1$ when a single direction dominates. It replaces $d$ in both the upper bound (Theorem 6.5) and the oracle complexity (Theorem 7.3), so that when Fisher information is concentrated ($d_{\operatorname{eff}}\ll d$), the guarantees are correspondingly tighter than classical Euclidean bounds.
6.2 Upper Bound
Lemma 6.4 (One-step Fisher Lyapunov drift).
Let $\Delta_{t}:=\theta_{t}-\theta^{\star}$ and $V_{t}:=\|\Delta_{t}\|_{F^{\star}}^{2}$. Assume: (i) $L$ is locally $M$-smooth near $\theta^{\star}$; (ii) $H^{\star}\succ 0$; (iii) (34) holds. Then, as long as $\theta_{t}$ stays in the local neighborhood,
$\mathbb{E}[V_{t+1}\mid\mathcal{F}_{t}]\;\leq\;\Bigl(1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}\Bigr)V_{t}\;+\;\eta_{t}^{2}\,\frac{1}{b}\,\mathrm{Tr}\!\bigl((F^{\star})^{2}\bigr)\;+\;\eta_{t}^{2}\,\frac{c_{1}}{b}\,\mathrm{Tr}(F^{\star})\,\rho_{t},$ (37)
where $c_{0},c_{1}$ depend only on $(M,\|F^{\star}\|_{\operatorname{op}},c_{r})$.
Proof.
From $\theta_{t+1}=\theta_{t}-\eta_{t}(h(\theta_{t})+\xi_{t})$, $\Delta_{t+1}=\Delta_{t}-\eta_{t}h(\theta_{t})-\eta_{t}\xi_{t}$. Expanding $V_{t+1}=\Delta_{t+1}^{\top}F^{\star}\Delta_{t+1}$ and taking conditional expectation (using $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$) yields
$\mathbb{E}[V_{t+1}\mid\mathcal{F}_{t}]=V_{t}-2\eta_{t}\Delta_{t}^{\top}F^{\star}h(\theta_{t})+\eta_{t}^{2}\,h(\theta_{t})^{\top}F^{\star}h(\theta_{t})+\eta_{t}^{2}\,\mathrm{Tr}\!\Bigl(F^{\star}\,\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\Bigr).$
Using $h(\theta^{\star}+\Delta)=H^{\star}\Delta+o(\|\Delta\|)$, we have $\Delta_{t}^{\top}F^{\star}h(\theta_{t})=\Delta_{t}^{\top}F^{\star}H^{\star}\Delta_{t}+o(V_{t})$. Writing $x_{t}:=(F^{\star})^{1/2}\Delta_{t}$ gives $\Delta_{t}^{\top}F^{\star}H^{\star}\Delta_{t}=x_{t}^{\top}\,\operatorname{Sym}\bigl((F^{\star})^{1/2}H^{\star}(F^{\star})^{-1/2}\bigr)\,x_{t}\geq\mu_{F}V_{t}$. Local $M$-smoothness implies $h(\theta_{t})^{\top}F^{\star}h(\theta_{t})\leq c_{0}V_{t}$. For the noise term, plug (34):
$\mathrm{Tr}\!\Bigl(F^{\star}\,\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\Bigr)=\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\mathrm{Tr}(F^{\star}r_{t}).$
Finally $|\mathrm{Tr}(F^{\star}r_{t})|\lesssim\mathrm{Tr}(F^{\star})\|r_{t}\|_{\operatorname{op}}\leq(c_{1}/b)\mathrm{Tr}(F^{\star})\rho_{t}$, giving (37). ∎
Theorem 6.5 (Upper bound in frozen Fisher geometry).
Assume Lemma 6.4. Take $\eta_{t}=\eta_{0}/t$ with $\eta_{0}>1/(2\mu_{F})$, and assume $\sum_{t\geq 1}\eta_{t}^{2}\rho_{t}<\infty$. Then for all $T\geq 2$,
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]\;\leq\;\frac{C_{1}}{b}\cdot\frac{\mathrm{Tr}\!\bigl((F^{\star})^{2}\bigr)}{\mu_{F}}\cdot\frac{1}{T}\;+\;C_{2}\cdot\frac{\|\theta_{1}-\theta^{\star}\|_{F^{\star}}^{2}}{T^{2\mu_{F}\eta_{0}}}\;+\;\frac{C_{3}}{b}\cdot\frac{\mathrm{Tr}(F^{\star})}{T},$ (38)
where $C_{1},C_{2},C_{3}$ depend only on $(\eta_{0},\mu_{F},c_{0},c_{1})$.
Moreover, using $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\,\mathrm{Tr}(F^{\star})$, we obtain
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]$ $\;\leq\;\frac{\widetilde{C}_{1}}{b}\cdot\frac{\lambda_{\max}(F^{\star})\,\mathrm{Tr}(F^{\star})}{\mu_{F}}\cdot\frac{1}{T}\;+\;\widetilde{C}_{2}\cdot\frac{\|\theta_{1}-\theta^{\star}\|_{F^{\star}}^{2}}{T^{2\mu_{F}\eta_{0}}}$
$\;=\;\frac{\widetilde{C}_{1}}{b}\cdot\frac{\lambda_{\max}(F^{\star})^{2}}{\mu_{F}}\cdot\frac{d_{\operatorname{eff}}(F^{\star})}{T}\;+\;\widetilde{C}_{2}\cdot\frac{\|\theta_{1}-\theta^{\star}\|_{F^{\star}}^{2}}{T^{2\mu_{F}\eta_{0}}}.$ (39)
Proof.
Take total expectations in (37) to obtain $\mathbb{E}[V_{t+1}]\leq a_{t}\,\mathbb{E}[V_{t}]+u_{t}$ with $a_{t}:=1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}$ and $u_{t}:=\eta_{t}^{2}\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\eta_{t}^{2}\frac{c_{1}}{b}\mathrm{Tr}(F^{\star})\rho_{t}$. For $\eta_{t}=\eta_{0}/t$, there exists $t_{0}$ such that for $t\geq t_{0}$, $a_{t}\leq 1-\mu_{F}\eta_{0}/t$. Define $A_{t,T}:=\prod_{s=t}^{T-1}a_{s}$. Unrolling gives $\mathbb{E}[V_{T}]\leq A_{1,T}V_{1}+\sum_{t=1}^{T-1}A_{t+1,T}u_{t}$. Standard product bounds yield $A_{1,T}\leq CT^{-2\mu_{F}\eta_{0}}$ and $A_{t+1,T}\leq C(t/T)^{\mu_{F}\eta_{0}}$. Then, since $\mu_{F}\eta_{0}>1/2$, $\sum_{t\leq T}A_{t+1,T}\eta_{t}^{2}=O(1/T)$, giving the leading $(\mathrm{Tr}((F^{\star})^{2})/b)(1/T)$ term. The remainder term is controlled because $\sum_{t}\eta_{t}^{2}\rho_{t}<\infty$ and $A_{t+1,T}\leq 1$, yielding an additional $O((\mathrm{Tr}(F^{\star})/b)(1/T))$ contribution. This proves (38). The reduction to (6.5) follows from $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\mathrm{Tr}(F^{\star})$. ∎
Remark 6.6 (Consistency with the OU viewpoint).
The $1/(Tb)$ scaling in the frozen Fisher norm is consistent with the OU/Lyapunov stationary analysis: both predict the same leading-order dependence on temperature $\tau=\eta/b$ and intrinsic curvature $\mu_{F}$.
6.3 Lower Bound
The lower bound proceeds in three layers of increasing generality. We first state the cleanest result in the parametric likelihood/Fisher case, then interpret it through oracle information, and finally discuss the extension to general losses.
6.3.1 Layer 1: Parametric Fisher-case lower bound.
Assumption 6.7 (Local parametric regularity for van Trees).
Let $\{P_{\theta}:\theta\in\Theta_{\operatorname{loc}}\}$ be a parametric family on a neighborhood $\Theta_{\operatorname{loc}}$ of $\theta^{\star}$ with densities $p_{\theta}$. Assume:
  1. (i)
    $\theta\mapsto p_{\theta}(z)$ is twice continuously differentiable for a.e. $z$;
  2. (ii)
    the Fisher information $F(\theta):=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}]$ is continuous and positive definite on $\Theta_{\operatorname{loc}}$;
  3. (iii)
    $\int\|s_{\theta}(z)\|^{2}\,p_{\theta}(z)\,dz$ is uniformly bounded on $\Theta_{\operatorname{loc}}$;
  4. (iv)
    there exists a smooth prior density $\pi$ supported on $\Theta_{\operatorname{loc}}$ with finite Fisher information $J_{\pi}:=\int\|\nabla\log\pi(\theta)\|^{2}\pi(\theta)\,d\theta<\infty$.
Theorem 6.8 (Fisher-case lower bound via van Trees).
Suppose Assumption 6.7 holds and the estimator $\hat{\theta}_{N}$ is based on $N$ i.i.d. observations from $P_{\theta^{\star}}$. Then
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (40)
for a universal constant $c>0$.
Proof.
This is the special case of Proposition 6.10 below with i.i.d. observations (so the oracle information bound holds with equality). We give the self-contained argument in the proof of Proposition 6.10. ∎
6.3.2 Layer 2: Oracle-information interpretation.
The heart of the lower bound is not i.i.d. sampling per se, but bounded predictable information growth. We now restate the result in terms of an abstract oracle transcript.
Assumption 6.9 (Oracle–information structure).
The oracle produces observations $Y_{t}=\psi(\theta_{t};Z_{t})$, where:
  1. (i)
    $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, with filtration $\mathcal{F}_{t}:=\sigma(Y_{1},\dots,Y_{t})$;
  2. (ii)
    Conditional unbiasedness: $\mathbb{E}_{\theta^{\star}}[Y_{t}\mid\mathcal{F}_{t-1}]=0$;
  3. (iii)
    Predictable information bound: $\mathbb{E}_{\theta^{\star}}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq\mathcal{I}_{1}(\theta^{\star})$, a.s., where $\mathcal{I}_{1}(\theta^{\star}):=F(\theta^{\star})=F^{\star}$ in the likelihood case;
  4. (iv)
    Uniform moment control: $\sup_{t}\mathbb{E}_{\theta^{\star}}\|Y_{t}\|^{4}<\infty$.
Under Assumption 6.9, the predictable quadratic variation of the oracle transcript satisfies $\sum_{t=1}^{N}\mathbb{E}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq N\mathcal{I}_{1}(\theta^{\star})$, so the total information available to any $\mathcal{F}_{N}$-measurable estimator is bounded by $N\mathcal{I}_{1}$.
Proposition 6.10 (Lower bound under predictable information budget).
Suppose Assumption 6.7 holds for the parametric family $\{P_{\theta}\}$ and the oracle satisfies Assumption 6.9 with $\mathcal{I}_{1}(\theta^{\star})=F^{\star}$. Additionally, assume that each conditional law of $Z_{t}$ given $\mathcal{F}_{t-1}$ has a density differentiable in $\theta$, so that the conditional Fisher information $I_{t}(\theta\mid\mathcal{F}_{t-1})$ exists and satisfies $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$ a.s. Then every $\mathcal{F}_{N}$-measurable estimator $\hat{\theta}_{N}$ satisfies
$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (41)
for a universal constant $c>0$.
In particular, if one SGD iteration uses a mini-batch of size $b$ (so $N=Tb$ oracle calls), then $\inf_{\hat{\theta}_{T}}\sup_{\theta}\mathbb{E}[\|\hat{\theta}_{T}-\theta\|_{F(\theta)}^{2}]\geq c\,d/(Tb)$.
Proof.
The argument extends Theorem 6.8 from i.i.d. observations to the adaptive oracle transcript. The van Trees bound requires an upper bound on the Fisher information $I_{N}(\theta)$ of the joint likelihood of $(Y_{1},\ldots,Y_{N})$ with respect to $\theta$. Three ingredients deliver this bound.
Step 1 (Adaptivity does not inflate per-step information). At step $t$, the query $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, but $Z_{t}\sim P_{\theta}$ is drawn independently of $\mathcal{F}_{t-1}$ given $\theta$. Hence the conditional distribution of $Z_{t}$ given $\mathcal{F}_{t-1}$ depends on $\theta$ only through $\{P_{\theta}\}$, not through the past transcript.
Step 2 (Second-moment bound == per-step Fisher bound). Under correct specification, $Y_{t}=-s_{\theta}(Z_{t})$ is the negative score. The conditional Fisher information of $Z_{t}$ given $\mathcal{F}_{t-1}$ about $\theta$ is $I_{t}(\theta\mid\mathcal{F}_{t-1})=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}\mid\mathcal{F}_{t-1}]=F(\theta)$—exactly the score second-moment matrix. So Assumption 6.9(iii) directly bounds $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$. To extend from $\theta^{\star}$ to all $\theta\in\Theta_{\operatorname{loc}}$, continuity of $F(\theta)$ (Assumption 6.7(ii)) gives $F(\theta)\preceq(1+\epsilon)F^{\star}$ on a sufficiently small $\Theta_{\operatorname{loc}}$; the factor $(1+\epsilon)$ is absorbed into the universal constant $c$.
Step 3 (Chain rule and conclusion). The chain rule for conditional Fisher information applied to the sequential factorization of the joint likelihood gives
$I_{N}(\theta)=\sum_{t=1}^{N}\mathbb{E}_{\theta}\!\bigl[I_{t}(\theta\mid\mathcal{F}_{t-1})\bigr]\preceq N\,(1{+}\epsilon)\,F^{\star}.$
Substituting into van Trees with prior $\pi$ and sending $\epsilon\to 0$ yields (41). ∎
Remark 6.11 (Beyond i.i.d. oracles).
The classical i.i.d. setting satisfies Assumption 6.9 with equality in (iii). The assumption also accommodates exchangeable or conditionally i.i.d. sampling schemes (e.g., de Finetti mixtures) provided the per-step conditional information is bounded by $\mathcal{I}_{1}$. What is required is a predictable information bound on the oracle transcript, not independence across calls. The critical logical step is that the van Trees inequality applies to the joint transcript likelihood, whose total Fisher information is controlled by summing the per-step conditional bounds.
Remark 6.12 (What enters the lower bound).
The lower bound (41) involves $d$ (ambient dimension) and $N$ (oracle budget). The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ enter the upper bound and oracle complexity (Theorems 6.5 and 7.3) but not the lower bound, which is stated in terms of $d$ to give the sharpest minimax comparison. When $d_{\mathrm{eff}}\ll d$, the gap between upper and lower bounds tightens correspondingly.
6.3.3 Layer 3: General-loss / Godambe extension.
Remark 6.13 (Extension to general losses).
In the general-loss case, the same lower-bound form $\Omega(d/N)$ is recovered under local asymptotic conditions that identify the relevant sandwich geometry. Specifically, if the loss satisfies a local asymptotic normality (LAN) condition at $\theta^{\star}$ with information matrix $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$, then the van Trees argument carries through with $F^{\star}$ replaced by $G^{\star}(\theta^{\star})$ and the dual norm adjusted accordingly. The LAN condition is standard for smooth M-estimators under regularity conditions comparable to those already imposed in Section 3; see Le Cam (1986) and van der Vaart (1998). We state this extension as a remark rather than a theorem because the full verification of LAN for the adaptive oracle transcript of Assumption 6.9 requires additional technical conditions (uniform local asymptotic equicontinuity) that we do not develop here.
Corollary 6.14 (Matching rates (up to constants)).
Under Theorem 6.5 and the oracle model of Assumption 6.9,
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]=\Theta\!\left(\frac{1}{Tb}\right)\quad\text{equivalently}\quad\Theta\!\left(\frac{1}{N}\right)\text{ per oracle call},\quad N=Tb.$
7 Oracle Complexity
This section turns the Fisher-geometry mean-square bounds of Section 6 into an explicit oracle-complexity guarantee for achieving $\varepsilon$-stationarity. The steps are standard—smoothness descent, averaged-gradient bound, martingale concentration (cf. Nesterov, 2004, Chapter 2)—but we express the result in the natural Fisher dual norm and keep the statistical conditioning explicit.
Throughout, we work locally near $\theta^{\star}$ and freeze the metric $F^{\star}:=F(\theta^{\star})\succ 0$ (for general losses, replace $F^{\star}$ by $G^{\star}(\theta^{\star})$; see Remark 7.5).
7.1 Stationarity in the Fisher dual norm
Definition 7.1 ($\varepsilon$-stationarity in Fisher dual norm).
Let $F^{\star}\succ 0$. We say $\theta$ is $\varepsilon$-stationary (in the Fisher metric) if
$\|\nabla L(\theta)\|_{(F^{\star})^{-1}}\;:=\;\sqrt{\nabla L(\theta)^{\top}(F^{\star})^{-1}\nabla L(\theta)}\;\leq\;\varepsilon.$
7.2 From Fisher distance to Fisher-dual stationarity
Lemma 7.2 (Distance-to-stationarity conversion).
Assume $L$ is $M$-smooth on a neighborhood $\mathcal{U}$ of $\theta^{\star}$ and $F^{\star}\succ 0$. Then for all $\theta\in\mathcal{U}$,
$\|\nabla L(\theta)\|_{(F^{\star})^{-1}}^{2}\;\leq\;\frac{M^{2}}{\lambda_{\min}(F^{\star})^{2}}\,\|\theta-\theta^{\star}\|_{F^{\star}}^{2}.$ (42)
Moreover, if $\theta\to\theta^{\star}$ and $\nabla L(\theta)=H^{\star}(\theta-\theta^{\star})+o(\|\theta-\theta^{\star}\|)$, then locally
$\|\nabla L(\theta)\|_{(F^{\star})^{-1}}^{2}\;\leq\;(1+o(1))\,\lambda_{\max}\!\bigl(F^{\star-1/2}H^{\star}F^{\star-1/2}\bigr)^{2}\,\|\theta-\theta^{\star}\|_{F^{\star}}^{2}.$ (43)
Proof.
By $M$-smoothness and $\nabla L(\theta^{\star})=0$, $\|\nabla L(\theta)\|_{2}\leq M\|\theta-\theta^{\star}\|_{2}$. The dual-norm bound $\|v\|_{(F^{\star})^{-1}}^{2}\leq\lambda_{\min}(F^{\star})^{-1}\|v\|_{2}^{2}$ converts Euclidean gradient norm to Fisher dual norm, and $\|\theta-\theta^{\star}\|_{2}^{2}\leq\lambda_{\min}(F^{\star})^{-1}\|\theta-\theta^{\star}\|_{F^{\star}}^{2}$ converts Fisher distance to Euclidean. Combining: $\|\nabla L\|_{(F^{\star})^{-1}}^{2}\leq\lambda_{\min}(F^{\star})^{-1}M^{2}\lambda_{\min}(F^{\star})^{-1}\|\theta-\theta^{\star}\|_{F^{\star}}^{2}=M^{2}/\lambda_{\min}(F^{\star})^{2}\cdot\|\theta-\theta^{\star}\|_{F^{\star}}^{2}$. The local refinement follows by replacing $M$ by $\|H^{\star}\|_{\operatorname{op}}$ up to $1+o(1)$ via the linearization. ∎
7.3 High-probability oracle complexity
Theorem 7.3 (Oracle complexity for Fisher-dual stationarity).
Assume the local regularity conditions of Section 3 hold on a neighborhood $\mathcal{U}$ of $\theta^{\star}$, and suppose that on $\mathcal{U}$ the mini-batch noise is Fisher-aligned in the sense that, for some batch size $b\geq 1$,
$\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0,\qquad\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\preceq\frac{1}{b}\,F^{\star}.$ (44)
Additionally, assume conditional sub-Gaussianity of the noise: for all $v\in\mathbb{R}^{d}$ with $\|v\|=1$ and all $\lambda\in\mathbb{R}$, $\mathbb{E}[\exp(\lambda\,v^{\top}\xi_{t})\mid\mathcal{F}_{t}]\leq\exp(\lambda^{2}\|F^{\star}\|_{\operatorname{op}}/(2b))$ a.s. (This is satisfied, for example, when per-sample gradients are almost surely bounded on $\mathcal{U}$.) Run SGD with constant batch size $b$ and constant stepsize $\eta\leq 1/(4M)$:
$\theta_{t+1}=\theta_{t}-\eta\bigl(\nabla L(\theta_{t})+\xi_{t}\bigr).$
Define the Fisher condition number and intrinsic dimension
$\kappa_{F}:=\frac{\lambda_{\max}(F^{\star})}{\lambda_{\min}(F^{\star})},\qquad d_{\mathrm{eff}}(F^{\star}):=\frac{\mathrm{Tr}(F^{\star})}{\lambda_{\max}(F^{\star})}.$
Then there exists a universal constant $C>0$ such that, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\min_{0\leq t\leq T-1}\;\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C\left(\frac{L(\theta_{0})-L(\theta^{\star})}{\eta\,T\,\lambda_{\min}(F^{\star})}\;+\;\frac{\eta}{b}\,\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\right)\log\!\frac{1}{\delta}.$ (45)
In particular, choosing $(\eta,T)$ so that the right-hand side is $\leq\varepsilon^{2}$ implies $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}\leq\varepsilon$ with probability at least $1-\delta$.
Equivalently, in terms of oracle calls $N:=Tb$ (sample-gradient evaluations), one sufficient scaling (up to constants and ignoring the transient term $\frac{L(\theta_{0})-L(\theta^{\star})}{\eta T\lambda_{\min}(F^{\star})}$) is
$N\;=\;\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})}{\varepsilon^{2}}\,\log\!\frac{1}{\delta}\right).$ (46)
Proof.
Step 1 (descent inequality). By $M$-smoothness and $\eta\leq 1/(4M)$, expand $L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L,\theta_{t+1}-\theta_{t}\rangle+\frac{M}{2}\|\theta_{t+1}-\theta_{t}\|^{2}$ with $\theta_{t+1}-\theta_{t}=-\eta(\nabla L(\theta_{t})+\xi_{t})$. After expanding and using $\eta\leq 1/(4M)$, so $1-M\eta/2\geq 1-1/8=7/8$ and $\eta M\leq 1/4$:
$L(\theta_{t+1})\leq L(\theta_{t})-\tfrac{3\eta}{4}\|\nabla L(\theta_{t})\|_{2}^{2}+\eta(1+M\eta)\langle\nabla L(\theta_{t}),\xi_{t}\rangle+\tfrac{M\eta^{2}}{2}\|\xi_{t}\|_{2}^{2}.$
Step 2 (define martingale). Let $\Delta M_{t+1}:=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ and $M_{T}:=\sum_{t=0}^{T-1}\Delta M_{t+1}$; since $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$, this is a martingale.
Step 3 (sum the descent). Summing from $t=0$ to T1T-1:
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}}{2}\,Q_{T},\quad\Delta L:=L(\theta_{0})-L(\theta^{\star}),$
where $Q_{T}:=\sum_{t<T}\|\xi_{t}\|_{2}^{2}$. To replace the random $Q_{T}$ by a deterministic bound we need concentration, not just conditional moments. Under conditional sub-Gaussianity, $\|\xi_{t}\|_{2}^{2}$ is conditionally sub-exponential with $\mathbb{E}[\|\xi_{t}\|_{2}^{2}\mid\mathcal{F}_{t}]\leq b^{-1}\mathrm{Tr}(F^{\star})$; Bernstein’s inequality for sums of conditionally sub-exponential random variables (see, e.g., Wainwright 2019, Cor. 2.21) gives, with probability at least $1-\delta$, $Q_{T}\leq(T/b)\mathrm{Tr}(F^{\star})+c\sqrt{T\,d}\,(\|F^{\star}\|_{\operatorname{op}}/b)\sqrt{\log(1/\delta)}.$ Since $M\eta^{2}/2\leq\eta/8$ (as $\eta\leq 1/(4M)$), the deviation is $O(\eta\sqrt{T/b}\,\sqrt{d}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)}\,/b)$, which is lower-order relative to the bias term $\Delta L/(\eta T)$ in the final bound. Absorbing it into the constant $C$ (and taking a union bound over the two $\delta$-events, i.e. replacing $\delta$ by $\delta/2$ throughout):
$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})\log\!\tfrac{2}{\delta},$
with probability at least $1-\delta$.
Step 4 (sub-Gaussian martingale concentration). Each increment satisfies $\Delta M_{t+1}=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ where $\nabla L(\theta_{t})$ is $\mathcal{F}_{t}$-measurable. By the conditional sub-Gaussianity assumption, the scalar $\langle v,\xi_{t}\rangle$ with $v:=\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|_{2}$ has sub-Gaussian parameter $\sigma_{t}^{2}\leq\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $c_{t}^{2}:=\eta^{2}\|\nabla L(\theta_{t})\|_{2}^{2}\|F^{\star}\|_{\operatorname{op}}/b$. By the standard exponential supermartingale bound for sums of conditionally sub-Gaussian increments (see, e.g., Wainwright 2019, Theorem 2.19), with probability at least $1-\delta$, $M_{T}\leq\sqrt{2\sum_{t<T}c_{t}^{2}\,\log(1/\delta)}=\eta\sqrt{2(\|F^{\star}\|_{\operatorname{op}}/b)\,S_{T}\,\log(1/\delta)}$, where $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$.
Step 5 (solve for gradient bound). Substituting the Step 4 tail bound into Step 3 gives a quadratic inequality in $\sqrt{S_{T}}$; solving:
$\min_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}\;\leq\;\frac{S_{T}}{T}\;\leq\;C\!\left(\frac{\Delta L}{\eta T}+\frac{\eta}{b}\mathrm{Tr}(F^{\star})\right)\!\log\!\frac{1}{\delta}.$
Step 6 (Fisher dual norm). Use $\|v\|_{(F^{\star})^{-1}}^{2}\leq\lambda_{\min}(F^{\star})^{-1}\|v\|_{2}^{2}$ and $\mathrm{Tr}(F^{\star})/\lambda_{\min}(F^{\star})=\kappa_{F}\,d_{\operatorname{eff}}(F^{\star})$:
$\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C\!\left(\frac{\Delta L}{\eta T\lambda_{\min}(F^{\star})}+\frac{\eta}{b}\,\kappa_{F}\,d_{\operatorname{eff}}(F^{\star})\right)\!\log\!\frac{1}{\delta},$
establishing (45). The oracle complexity (46) follows by balancing the two terms and setting $N=Tb$. ∎
Remark 7.4 (Statistical vs. Euclidean conditioning).
Classical oracle-complexity results use the Hessian condition number $\kappa_{H}$ and ambient dimension $d$; here they are replaced by the Fisher counterparts $\kappa_{F}$ and $d_{\operatorname{eff}}$. The substitution is not cosmetic: a problem may be Euclidean-stiff ($\kappa_{H}\gg 1$) yet statistically well-conditioned ($\kappa_{F}=O(1)$) when curvature and information share the same eigenstructure; conversely, isotropic curvature does not guarantee favorable statistical complexity.
Remark 7.5 (General losses / Godambe geometry).
For general losses, replace $F^{\star}$ by $G^{\star}(\theta^{\star})$ throughout and interpret the dual norm accordingly. The alignment condition (44) is then exactly the local conclusion of Theorem 4.3.
Remark 7.6 (Iteration vs. oracle complexity).
Equation (46) counts oracle calls $N=Tb$. If reporting iteration complexity, keep $b$ explicit and interpret $T=N/b$.
Remark 7.7 (Interpreting the complexity constants).
The product $\kappa_{F}\,d_{\operatorname{eff}}$ in (46) measures the total cost of resolving all statistically relevant directions at the hardest direction’s scale (see Remark 6.3 for the role of $d_{\operatorname{eff}}$ as the stable rank of $F^{\star}$). Equation (46) links the sampling budget NN to achievable stationarity through intrinsic problem geometry.
Corollary 7.8 (Batch-size scaling under a fixed sample budget).
Fix a total sampling budget $N$ and run $T=N/b$ iterations of constant-stepsize SGD with batch size $b$. Equating the bias term $\sim(L(\theta_{0})-L^{\star})/(\eta T\lambda_{\min}(F^{\star}))$ and the variance term $\sim\eta\kappa_{F}d_{\operatorname{eff}}/b$ in (45) yields the balanced stepsize $\eta^{\star}\propto\sqrt{b/(T\kappa_{F}d_{\operatorname{eff}})}$. Substituting $T=N/b$, the resulting Fisher-dual stationarity gap satisfies
$\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;=\;\mathcal{O}\!\left(\sqrt{\frac{\kappa_{F}\,d_{\operatorname{eff}}}{N}}\right),$
which is independent of $b$ at leading order. The practical implication for OR settings is that, under this budget-constrained regime, small batches lose nothing in worst-case guarantee while gaining more frequent iterates; the Fisher/Godambe geometry determines the constant but not the scaling in $N$.
8 Numerical Validation
We validate the theory through six diagnostic experiments summarized in Table 3. The experiments probe downstream consequences of the identified covariance $\mathrm{Cov}(g_{B})\approx b^{-1}G^{\star}$ (Theorem 4.3): if the identification is correct, the Lyapunov equilibrium, convergence rates, and directional variance structure must all agree with the predictions derived from $G^{\star}$. Appendix B additionally validates the covariance identity itself under misspecification and approximate exchangeability.
Table 3: Reproducibility capsule: key experimental parameters.
Experiment $d$ Replicates $\eta$ Steps Sweep variable
1 (Batch scaling) 2 500 0.05 200 000 $b\in\{1,2,4,\ldots,512\}$
2 (Angle sweep) 2 500 0.05 200 000 $\theta\in[0,\pi/2]$
3 (Rotation diagnostic) 2 $\phi\in[0,\pi/2]$ (theory)
4 (OU plateau, d=10d\!=\!10) 10 2 000 $0.5/\lambda_{\max}$ 300–10 000 $T$ (horizon)
5 ($1/N$ rate) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 $N=Tb$
6 (Directional var.) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 eigenvectors of $G^{\star}$
Experiments 1–3: Euler–Maruyama, $d=2$. Experiments 4–6: direct SGD recursion, $d=10$, $b=64$.
8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
The first claim is that the Fisher–Lyapunov balance (Corollary 5.7) is not merely a qualitative insight but a quantitatively accurate prediction of steady-state risk, and that the shape of the noise covariance—not just its scale—governs which directions receive noise energy at equilibrium. Three experiments validate this at increasing dimensionality.
Throughout, when the simulator uses finite-step (Euler–Maruyama or direct SGD) dynamics, the mathematically correct benchmark is the discrete Lyapunov equation (Proposition 5.9); the continuous-time OU prediction is its small-$\eta$ asymptotic reduction. The regime of approximation accuracy is governed by the product $\eta\lambda_{\max}(H^{\star})$: when this quantity is small (say $\leq 0.1$), the continuous and discrete benchmarks are nearly indistinguishable; for moderate stepsizes the discrete equation is the appropriate comparator.
8.1.1 Experiment 1: Fisher–Lyapunov Equilibrium and $1/b$ Scaling (OU Benchmark).
We begin with the local OU surrogate implied by the diffusion limit near a nondegenerate optimum:
$d\theta_{t}\;=\;-\eta\,H\,\theta_{t}\,dt\;+\;\sqrt{\frac{2\eta}{b}}\;F^{1/2}\,dW_{t},$ (47)
with a two-dimensional sharp/flat curvature $H=\mathrm{diag}(1,0.1)$ and isotropic noise $F=I_{2}$, sweeping $b$ over a wide range. At stationarity, the covariance $\Sigma(b)$ solves the Lyapunov equation
$H\Sigma+\Sigma H\;=\;\frac{2}{b}\,F,$ (48)
so every stationary marginal variance scales as $1/b$; $\eta$ governs mixing time but not the fixed point. For the Euler–Maruyama simulation used here, the discrete Lyapunov equation (49) is the exact benchmark; the continuous equation (48) coincides at leading order because the stepsize satisfies $\eta\lambda_{\max}(H)\leq 0.05$.
Figure 2 overlays simulated stationary variances with the Lyapunov prediction and shows clean $1/b$ scaling across the batch sweep. The flat coordinate mixes more slowly (as expected under weak curvature) but converges to the same equilibrium law. This serves as a baseline check: the Fisher–Lyapunov equilibrium is quantitatively predictive, not merely qualitative. We observed early entry into the basin of Fisher geometry even from random starts; see Section 9.3 for a diagnostic based on the stability ratios $\delta_{F}$ and $\delta_{H}$.
Figure 2: OU benchmark: stationary variances vs. batch size $b$ (log scales). Solid: simulation; dashed: Lyapunov prediction from (48).
$d=2$, $H=\mathrm{diag}(1,0.1)$, $F=I_{2}$, 500 replicates, 200k steps.
8.1.2 Experiment 2: Anisotropic Fisher–Hessian Misalignment (Angle Sweep).
We keep the same curvature $H=\mathrm{diag}(1,0.1)$ but rotate an anisotropic noise-shape matrix relative to the curvature basis:
$F(\varphi)=R(\varphi)\begin{pmatrix}1.5&0\\[2.0pt] 0&0.5\end{pmatrix}R(\varphi)^{\top},\qquad\varphi\in[0,\pi/3],$
with fixed batch size, isolating alignment as an operational parameter. Because the simulator is Euler–Maruyama with finite time step, the appropriate comparator is the discrete Lyapunov equation: writing $A=I-\eta H\Delta t$ and $Q=(2\eta/b)F(\theta)\Delta t$, the stationary covariance solves
$\Sigma\;=\;A\Sigma A^{\top}+Q.$ (49)
Figure 3 shows a clear transfer of stationary variance across coordinates as $\varphi$ varies: when the noise eigenbasis rotates away from the curvature eigenbasis, equilibrium variance is redistributed in a predictable way, and the discrete Lyapunov benchmark tracks the simulation closely. The takeaway is structural: alignment governs which directions receive noise energy at equilibrium, even when scalar temperature $\eta/b$ is held fixed.
Figure 3: Angle sweep: stationary variances vs. misalignment angle. Bands show replicate variability; dashed curves are the discrete Lyapunov prediction from (49).
$d=2$, $H=\mathrm{diag}(1,0.1)$, 500 replicates per angle.
8.1.3 Experiment 4: OU/Lyapunov Plateau in $d=10$ (Anisotropic vs. Isotropic).
Experiments 1–2 used $d=2$. We now move to $d=10$ to validate the Lyapunov equilibrium in a setting where anisotropy is richer. The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100; the noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
Table 4 confirms the Lyapunov prediction (Corollary 5.7): both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}/R_{\infty}$ approaching unity from both sides. The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; directional differences emerge only in higher-order structure (Claim 3 below).
Empirical Risk $\widehat{R}_{T}$ Lyapunov $R_{\infty}$ $\widehat{R}_{T}/R_{\infty}$
$T$ Aniso Iso Aniso Iso Aniso Iso
300 $8.55\!\times\!10^{-4}$ $8.60\!\times\!10^{-4}$ $8.63\!\times\!10^{-4}$ $8.58\!\times\!10^{-4}$ 0.991 1.002
800 $8.64\!\times\!10^{-4}$ $8.49\!\times\!10^{-4}$ 1.002 0.990
2 000 $8.58\!\times\!10^{-4}$ $8.50\!\times\!10^{-4}$ 0.995 0.992
5 000 $8.72\!\times\!10^{-4}$ $8.58\!\times\!10^{-4}$ 1.010 1.001
10 000 $8.63\!\times\!10^{-4}$ $8.47\!\times\!10^{-4}$ 1.000 0.987
Table 4: OU/Lyapunov plateau ($d=10$): empirical Fisher-metric risk vs. discrete Lyapunov prediction $R_{\infty}=\mathrm{Tr}(F^{\star}\Sigma_{\infty})$. Convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$ confirm quantitative accuracy.
$d=10$, $H=\mathrm{diag}(1,\ldots,100)$, random $G^{\star}$ with eigenvalues in $[0.5,5]$, $\eta=0.5/\lambda_{\max}$, $b=64$, R=2, [^0] R=2{,}000 replicates.
8.2 Claim 2: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
The second claim is that under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$. Specifically, the SA central-limit theorem gives an asymptotic covariance $V$ satisfying $({\eta_{0}H^{\star}-\tfrac{1}{2}I})V+V({\eta_{0}H^{\star}-\tfrac{1}{2}I})=\frac{\eta_{0}^{2}}{b}\,G^{\star}$, and $N\cdot\mathrm{Risk}\to b\,\mathrm{Tr}(F^{\star}V)$. This validates the matching upper and lower bounds (Theorem 6.5 and Proposition 6.10).
8.2.1 Experiment 5: $1/N$ Rate and Fisher Constant (Decaying Step Size).
We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 4, with oracle budget $N=Tb$ and $b=64$. Figure 4 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), confirming Theorem 6.5. The right panel shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$, which stabilizes near $\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic)—both close to the SA–CLT prediction $b\,\mathrm{Tr}(F^{\star}V)$ where $V$ solves the Lyapunov equation above. The inset marks the auxiliary constant $\mathrm{Tr}(G^{\star}H^{-1})\approx 5.41$, which governs the constant-step OU stationary risk (Corollary 5.7); the ratio of the decaying-step plateau to this constant reflects the step-size schedule and the interplay between curvature and noise geometry.
Figure 4: Decaying step size, $d=10$. Left: Fisher-metric risk vs. NN (log–log); slopes $\approx-1$ confirm $1/N$ rate. Right: Scaled risk $N\cdot\widehat{\mathrm{Risk}}$ stabilizes; inset shows $C=\mathrm{Tr}(G^{\star}H^{-1})$.
$\eta_{t}=1.2/(t+50)$, $b=64$, 800 replicates.
Table 5 summarizes the tightness of the bounds. The empirical risk is bounded between the information-theoretic lower bound (Proposition 6.10) and the upper bound (Theorem 6.5), with $N\cdot\text{Risk}$ stabilizing near $14.2$ (anisotropic) and $14.0$ (isotropic), both above the van Trees lower bound cd=10c\,d=10. At the largest budget (N=614,400N=614{,}400) the ratio of empirical constant to lower bound is approximately $1.4\times$, indicating near-tightness. The rate and constant are thus determined by $(G^{\star},H^{\star})$, so the oracle budget for a target risk can be computed from the problem’s intrinsic geometry.
Table 5: Upper vs. lower bound: scaled risk $N\cdot\widehat{\mathrm{Risk}}$ compared with the van Trees lower bound $cd=10$.
NN $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{aniso}}$ $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{iso}}$ $d=10$ (lower bound)
38 400 13.59 13.41 10
76 800 13.66 13.44 10
153 600 13.66 14.26 10
307 200 14.07 14.25 10
614 400 14.35 14.42 10
The SA–CLT Lyapunov prediction gives $b\,\mathrm{Tr}(F^{\star}V)\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic); $N\cdot\mathrm{Risk}$ stabilizes above the van Trees lower bound cd=10c\,d=10 as expected.
8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
The third claim is that matching the total noise power (scalar temperature $\tau=\eta/b$) is insufficient: the directional distribution of noise energy is a distinguishing signature of the identified geometric model that no isotropic surrogate can reproduce. The Lyapunov trace identity $\mathrm{Tr}(H^{\star}\Sigma_{\infty})=\frac{\tau}{2}\mathrm{Tr}(G^{\star})$ depends only on $\mathrm{Tr}(G^{\star})$, so a trace-matched isotropic surrogate reproduces total risk by construction; but the directional distribution is entirely wrong, spreading residual error uniformly rather than concentrating it along top noise eigendirections.
8.3.1 Experiment 3: Geometry vs. Scalar Temperature (Rotation Diagnostic).
We compare a geometric OU model (diffusion covariance $\propto F_{\star}(\phi)$, rotating with angle $\phi$) against a trace-matched isotropic surrogate ($\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(F_{\star})/d$). At stationarity, the geometric covariance $\Sigma_{F}(\phi)$ generically rotates (producing nontrivial cross-covariance), while the isotropic $\Sigma_{I}(\phi)$ remains axis-aligned. Figures 57 confirm this: the ellipses tilt, cross-covariance $\Sigma_{12}$ tracks theory, and marginal variances redistribute—effects that temperature matching alone cannot capture.
Figure 5: Stationary covariance ellipses at three rotation angles: the identified equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
$d=2$, $b=64$. Ellipses: 95% level sets of the Lyapunov solution.
Figure 6: Cross-covariance $\Sigma_{12}$ vs. rotation angle $\varphi$: isotropic trace-matched noise fails to reproduce off-diagonal covariance. Bands show $\pm 1$ s.d. across seeds.
$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in=800{}=800.
Figure 7: Marginal variances $\Sigma_{11}$ (left) and $\Sigma_{22}$ (right) vs. angle: total noise energy may match, but directional allocation differs.
8.3.2 Experiment 6: Directional Variance Amplification.
We now move to $d=10$ and ask whether anisotropy affects how risk is distributed across directions. For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$. Figure 8 (left) displays these on log–log axes: both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections. The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$; at the largest budget the ratios are $1.65$, $1.45$, and $1.26$ for $k=1,2,3$, mirroring the eigenvalue ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}$ of $G^{\star}$. These amplification ratios show that the identified geometry determines which parameter directions carry the most residual risk—information that scalar temperature models discard.
Figure 8: Directional variance amplification ($d=10$). Left: Directional second moments along top eigenvectors of $G^{\star}$. Right: Ratio of anisotropic to isotropic directional variances; ratios >1>1 mirror the eigenvalue ordering of $G^{\star}$. Temperature matching preserves $\mathrm{Tr}(G^{\star})$ but not the eigen-structure; anisotropic noise concentrates residual risk along the top noise eigendirections.
Same as Experiment 5; top $k=1,2,3$ eigenvectors of $G^{\star}$.
Additional robustness experiments under misspecification and approximate exchangeability are reported in Appendix B.
9 Discussion
9.1 Summary of Contributions
The paper’s central contribution is the identification of the mini-batch gradient covariance from sampling design and exchangeability as the projected gradient-noise matrix $G^{\star}(\theta)$—which reduces to Fisher information at $\theta^{\star}$ under correct specification and, for general losses, induces Godambe/sandwich geometry together with $H^{\star}$. This identification endogenizes the diffusion coefficient, determines the Lyapunov stationary covariance, and selects the natural metric for rates and complexity, yielding matching $1/N$ bounds governed by intrinsic dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$.
  • Proven: Alignment (Theorem 4.3), diffusion (Theorem 5.4), OU/Lyapunov (Corollary 5.7), discrete bridge (Proposition 5.9), matching bounds (Theorem 6.5 and Proposition 6.10), and oracle complexity (Theorem 7.3).
  • Sketched: Batch-size scaling (Corollary 7.8) and variance-reduction bridge (Remark 9.1).
  • Deferred: Misspecification, weak dependence, explicit batching policies, and OR-specific applications (Appendix B provides preliminary experiments).
9.2 Implications for Algorithm Design
In the diffusion regime the key knob is the effective temperature $\tau=\eta/b$. Changing $b$ does more than “reduce variance”; it changes the diffusion amplitude in a directional way determined by the identified geometry, and therefore changes the Lyapunov-predicted error floor. Because the intrinsic noise is matrix-valued, variance-reduction interventions should be evaluated by their effect on risk in a chosen metric (e.g., $\mathrm{Tr}(Q\Sigma)$) rather than Euclidean variance alone, connecting classical OR variance-reduction practice to SGD through the correct geometry.
Remark 9.1 (Variance reduction as Fisher-metric risk reduction).
A VR technique that reduces $G^{\star}$ to $\widetilde{G}\preceq G^{\star}$ in the Loewner order yields $\widetilde{\Sigma}_{\infty}\preceq\Sigma_{\infty}$ via the Lyapunov balance, and hence lower Fisher-metric risk.
SGD does not precondition the drift; natural-gradient methods do so by reweighting $\nabla L$. Our results explain why plain SGD nevertheless exhibits geometry-aware behavior through its noise structure: drift contraction and diffusion injection are distinct levers. The control view also suggests adaptive batching—increasing $b$ (cooling) when marginal variance reduction is valuable in the performance metric of interest, and decreasing it when exploration matters.
9.3 Practical Notes and Extensions
[…]
Our analysis freezes $F(\theta)$ locally at $F^{\star}$. Validity can be assessed by monitoring $\delta_{F}(t):=\|F(\theta_{t})-F^{\star}\|_{\operatorname{op}}/\|F^{\star}\|_{\operatorname{op}}$; in our experiments the Lyapunov predictions remain accurate from random starts, indicating early entry into the basin of Fisher geometry.
9.4 Estimating $G^{\star}$ in Practice
The preceding theory identifies $G^{\star}(\theta)$ as the central primitive. In applications where $G^{\star}$ is not available analytically, several estimation strategies provide operationally useful approximations. The simplest is the empirical covariance of per-sample gradients: given a pilot batch $\{z_{1},\ldots,z_{m}\}$ at $\theta_{t}$, the sample covariance $\hat{G}(\theta_{t}):=\frac{1}{m-1}\sum_{i=1}^{m}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ (where $\psi_{i}:=\nabla_{\theta}\ell(\theta_{t};z_{i})$) is the natural plug-in estimator and converges at the standard $O(1/\sqrt{m})$ rate in operator norm. For large $d$, a diagonal approximation $\hat{G}_{\mathrm{diag}}$ retains the per-coordinate noise scales and suffices for estimating $d_{\mathrm{eff}}$ and detecting gross anisotropy; when the leading eigendirections of $G^{\star}$ are the primary target, a rank-$k$ approximation via streaming PCA on per-sample gradients requires only $O(dk)$ storage. If $G^{\star}(\theta)$ varies slowly along the optimization path, an exponential moving average $\hat{G}_{t}=(1-\alpha)\hat{G}_{t-1}+\alpha\,(\psi_{t}-\bar{\psi}_{t})(\psi_{t}-\bar{\psi}_{t})^{\top}$ tracks the local geometry with minimal overhead.
[…]
9.5 Limitations and Future Work
Our sharpest statements use local linearization near a nondegenerate critical point; extending to nonconvex regimes requires controlling state-dependent curvature and relating intrinsic geometry to metastability. When $d\gg n$, the identified covariance may be low-rank, and a satisfactory theory should work on identifiable manifolds with appropriate pseudoinverses. Finally, preconditioners (Gauss–Newton, KFAC) reshape both drift and noise geometry; a joint control theory co-designing contraction and diffusion—potentially under communication constraints—would connect information geometry to OR-style variance allocation at scale.
Appendix A Technical Lemmas and Concentration Tools
This appendix collects the auxiliary lemmas invoked by the main-text proofs. All spine results—the alignment theorem, the diffusion limit, the OU linearization, the Lyapunov corollary, the upper and lower bounds, and the oracle complexity guarantee—are proved in the body of the paper.
A.1 Finite-population covariance identity
The following identity, classical in survey sampling (Cochran, 1977; Serfling, 1980), is used in the proof of the alignment theorem (Theorem 4.3).
Lemma A.1 (Without-replacement covariance of a mini-batch mean).
Let $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ be deterministic vectors with mean $\bar{x}_{n}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and centered sample covariance
$S_{n}:=\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}.$
Let $B$ be a subset of $\{1,\dots,n\}$ of size $b$ drawn uniformly without replacement, and define $\bar{x}_{B}:=\frac{1}{b}\sum_{i\in B}x_{i}$. Then, conditional on $(x_{1},\dots,x_{n})$,
$\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}\left(1-\frac{b}{n}\right)S_{n}.$ (50)
Proof.
Write $\bar{x}_{B}=\frac{1}{b}\sum_{i=1}^{n}I_{i}x_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$. Under SRSWOR, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$ for $i\neq j$. Expanding $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=b^{-2}\sum_{i,j}\mathrm{Cov}(I_{i},I_{j})x_{i}x_{j}^{\top}$ and collecting terms:
$b^{-2}\!\left[\frac{b(n-b)}{n^{2}}\sum_{i}x_{i}x_{i}^{\top}-\frac{b(n-b)}{n^{2}(n-1)}\sum_{i\neq j}x_{i}x_{j}^{\top}\right]$
$\quad=\frac{n-b}{bn^{2}}\cdot\frac{n}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}=\frac{1}{b}\!\left(1-\frac{b}{n}\right)S_{n},$
where the last equality uses $(n-1)S_{n}=\sum(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}$. ∎
A.2 Score identities and information equality
Lemma A.2 (Score identities and information equality).
Assume $\theta\mapsto p_{\theta}(z)$ is differentiable, $\int p_{\theta}(z)\,dz=1$, and differentiation under the integral sign is valid. Then:
  1. (i)
    $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$, where $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ is the score.
  2. (ii)
    $F(\theta)=\mathbb{E}_{\theta}[s_{\theta}(Z)s_{\theta}(Z)^{\top}]=-\mathbb{E}_{\theta}[\nabla_{\theta}^{2}\log p_{\theta}(Z)]$ (information equality).
Proof.
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
A.3 Lyapunov equation solution
Lemma A.3 (Lyapunov equation solution).
Let $A\in\mathbb{R}^{d\times d}$ with all eigenvalues having strictly positive real part. Let $Q\succeq 0$. Then the continuous Lyapunov equation $A\Sigma+\Sigma A^{\top}=Q$ has the unique solution
$\Sigma=\int_{0}^{\infty}e^{-As}Qe^{-A^{\top}s}ds.$
Proof.
Convergence: Since all eigenvalues of $A$ have positive real part, $\|e^{-As}\|\leq Ce^{-\lambda s}$ for some $C,\lambda>0$, ensuring convergence.
Verification: $A\Sigma+\Sigma A^{\top}=\int_{0}^{\infty}\frac{d}{ds}(-e^{-As}Qe^{-A^{\top}s})ds=[-e^{-As}Qe^{-A^{\top}s}]_{0}^{\infty}=Q$.
Uniqueness: If $\Sigma_{1},\Sigma_{2}$ both solve the equation, then $\Delta:=\Sigma_{1}-\Sigma_{2}$ satisfies $A\Delta+\Delta A^{\top}=0$, whose only solution is $\Delta=0$ when $A$ has eigenvalues with positive real part. ∎
A.4 Product bound for Robbins–Monro weights
Lemma A.4 (Product bound for Robbins-Monro weights).
Let $a>1$ and $c_{0}>0$. Define
$\alpha_{k}:=1-\frac{a}{k}+\frac{c_{0}}{k^{2}},\qquad\Pi_{s,T}:=\prod_{k=s}^{T-1}\alpha_{k}.$
Then for all $T$ sufficiently large:
  1. (i)
    $\Pi_{1,T}\leq CT^{-a}$ for a constant $C=C(a,c_{0})$.
  2. (ii)
    $\Pi_{t+1,T}\leq C(t/T)^{a}$ for $1\leq t<T$.
Proof.
Taking logarithms: $\log\alpha_{k}=\log(1-a/k+c_{0}/k^{2})=-a/k+O(1/k^{2})$ for large $k$. Thus $\log\Pi_{s,T}=-a\sum_{k=s}^{T-1}k^{-1}+O(1)=-a(\log T-\log s)+O(1)$. Exponentiating: $\Pi_{s,T}=O((s/T)^{a})$. For $s=1$: $\Pi_{1,T}=O(T^{-a})$. For $s=t+1$: $\Pi_{t+1,T}=O((t/T)^{a})$. ∎
A.5 Freedman’s inequality
Lemma A.5 (Freedman’s inequality).
Let $(M_{t})_{t\geq 0}$ be a martingale with $M_{0}=0$ and increments $\Delta M_{t}:=M_{t}-M_{t-1}$ satisfying $|\Delta M_{t}|\leq R$ almost surely. Let $V_{T}:=\sum_{t=1}^{T}\mathbb{E}[(\Delta M_{t})^{2}\mid\mathcal{F}_{t-1}]$ be the predictable quadratic variation. Then for any $v>0$ and $\lambda>0$:
$\Pr(M_{T}\geq\lambda\text{ and }V_{T}\leq v)\leq\exp\left(-\frac{\lambda^{2}}{2v+2R\lambda/3}\right).$
Equivalently, with probability at least $1-\delta$, whenever $V_{T}\leq v$:
$M_{T}\leq\sqrt{2v\log(1/\delta)}+\frac{2R}{3}\log(1/\delta).$
Proof.
See Freedman (1975) or Tropp (2015, Theorem 1.6). ∎
Appendix B Robustness Beyond Pure Exchangeability
The experiments in this appendix are exploratory robustness diagnostics, not part of the formal theoretical validation in Section 8. They demonstrate that the covariance-identification theory degrades gracefully under realistic departures from i.i.d. likelihood assumptions. Experiment A considers model misspecification, where curvature and noise geometry diverge; Experiment B models mild violations of exchangeability via variance inflation.
B.1 Experiment A: Misspecification (Probit DGP, Logistic Fit)
We generate n=50,000n=50{,}000 observations from a probit DGP and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$. Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}\!\bigl(\nabla_{\beta}\ell(\beta;X,Y)\big|_{\beta=\beta^{\star}}\bigr)\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]\big|_{\beta=\beta^{\star}}$. The covariance-identification theorem (Theorem 4.3) nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx(1/b)\,J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation. We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates). Figure 9 shows close agreement (relative Frobenius difference $\approx 0.05$), confirming that the identification remains accurate under misspecification: mini-batch fluctuations are governed by the score covariance $J^{\star}$ rather than by curvature alone.
Figure 9: Experiment A: misspecification (probit DGP, logistic fit). Diagonal entries of the empirical mini-batch covariance vs. $(1/b)\,\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.
B.2 Experiment B: Approximate Exchangeability and Variance Inflation
To model mild departures from ideal exchangeability (e.g., weak dependence, contamination, or partial reuse of samples), we inflate the intrinsic covariance by a factor $(1+\varepsilon)$ in the linearized surrogate: $\mathrm{Cov}(\xi_{t})=(1/b)(1+\varepsilon)\,G^{*}(\theta^{\star})$, $\varepsilon\in[0,1]$. This captures “extra correlation” as increased effective diffusion amplitude. Linearity of the Lyapunov operator then implies $\Sigma(\varepsilon)=(1+\varepsilon)\Sigma(0)$, so every stationary marginal variance inflates linearly in $\varepsilon$. Figure 10 confirms this prediction: the entire stationary covariance scales as though the system had a higher effective temperature $\tau_{\mathrm{eff}}=(1+\varepsilon)\eta/b$. To first order, small exchangeability violations act like “heating” the Fisher–Lyapunov equilibrium.
Figure 10: Experiment B: approximate exchangeability. Stationary variances vs. $\varepsilon$ (simulation bands vs. linear Lyapunov prediction).
Appendix C Reproducibility and Code Availability
All experiments are reproducible from the accompanying code archive. The Python suite (Python_Codes/) includes a master script run_all.py that executes all experiments sequentially. Every script uses explicit random seeds (global seed =7=7) for deterministic reproducibility; small platform-level floating-point differences may produce minor variations ($\lesssim 1\%$) in Monte Carlo estimates.
Software environment.
Python $\geq$3.9; NumPy, SciPy, Matplotlib, Pandas. Tested under Python 3.10.12, NumPy 2.2, SciPy 1.15, Matplotlib 3.10. Total wall-clock time: $\approx$4 minutes on a single CPU core.
Script inventory.
Table 6 maps each script to the figures and tables it generates.
Table 6: Code inventory: scripts, outputs, and paper cross-references.
Script Experiment Figures / Tables Sec.
partA_batch_scaling.py 1 (Batch scaling) Fig. 2 8.1
partB_anisotropic_alignment.py 2 (Angle sweep) Fig. 3 8.1
partC_rotation_sweep.py 3 (Rotation) Figs. 57 8.3
bounds.py 4–6 (OU plateau, rate, directional) Tab. 4, Figs. 48, Tab. 5 8.18.3
partD_misspecification.py A (Misspecification) Fig. 9 App. B
partE_eps_exchangeable.py B ($\varepsilon$-exchangeable) Fig. 10 App. B
The script run_all.py executes all of the above in sequence. The script partF_optimal_batching.py produces additional exploratory figures not included in the paper.
Shared parameters ($d=10$ experiments).
Hessian eigenvalues log-spaced in $[1,100]$; $F^{\star}=H^{\star}$ (quadratic Gaussian model). Anisotropic gradient covariance $G^{\star}$ drawn as a random SPD matrix with eigenvalues in $[0.5,5]$ (seed =7=7); isotropic baseline $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched). Constant-step experiments use $\eta=0.5/\lambda_{\max}(H)$, $b=64$, $\tau=\eta/b$; decaying-step experiments use $\eta_{t}=1.2/(t+50)$, $b=64$.
Verification.
All scripts were executed to completion with zero errors. Numerical outputs confirm: (i) log-log slopes $\approx-1$ for the $1/N$ risk decay; (ii) OU stationary variances match Lyapunov predictions to $1\%$; (iii) directional variance ratios align with the eigenstructure of $G^{\star}$; (iv) misspecification experiments show graceful degradation to the sandwich/Godambe prediction.

### 15 Apr 2026 v4

@@
-Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits
+Mini-Batch Covariance, Diffusion Limits, and Oracle Complexity in Stochastic Gradient Descent: A Sampling-Design Perspective
 Abstract
-Classical stochastic-approximation analyses treat the covariance of stochastic gradients as an exogenous modeling input.
-We show that under exchangeable mini-batch sampling this covariance is identified by the sampling mechanism itself: to leading order it is the projected covariance of per-sample gradients.
-In well-specified likelihood problems this reduces locally to projected Fisher information; for general M-estimation losses the same object is the projected gradient covariance $G^{\star}(\theta)$, which together with the Hessian induces sandwich/Godambe geometry.
-This identification—not the subsequent diffusion or Lyapunov machinery, which is classical once the noise matrix is given—is the paper’s main contribution.
-It endogenizes the diffusion coefficient (with effective temperature $\tau=\eta/b$), determines the stationary covariance via a Lyapunov equation whose inputs are now structurally fixed, and selects the identified statistical geometry as the natural metric for convergence analysis.
-We prove matching upper and lower bounds of order $\Theta(1/N)$ for risk in this metric under an oracle budget $N$; the lower bound is established first via a van Trees argument in the parametric Fisher setting and then extended to adaptive oracle transcripts under a predictable-information condition and mild conditional likelihood regularity.
-Translating these bounds into oracle complexity yields $\varepsilon$-stationarity guarantees in the Fisher dual norm that depend on an intrinsic effective dimension $d_{\mathrm{eff}}$ and a statistical condition number $\kappa_{F}$, rather than ambient dimension or Euclidean conditioning.
-Numerical experiments confirm the Lyapunov predictions at both continuous-time and discrete-time levels and show that scalar temperature matching cannot reproduce directional noise structure.
-Keywords: Stochastic gradient descent; Fisher information; Godambe information; information geometry; diffusion approximation; minimax rates; oracle complexity; batch-size control; simulation optimization
+Stochastic gradient descent (SGD) is central to simulation optimization, stochastic programming, and online M-estimation, where sampling effort is a decision variable.
+We study the mini-batch gradient noise as a sampling-design object.
+Under exchangeable fresh-sampling mini-batches, the conditional covariance given the de Finetti directing measure $\mu$ is $b^{-1}G_{\mu}(\theta)$, and under identifiability the projected population object is $b^{-1}G^{\star}(\theta)$—projected Fisher information for correctly specified likelihoods, the sandwich partner of the Hessian otherwise.
+This identification fixes the noise matrix entering the diffusion analysis of constant-step SGD: the raw iterate path has a deterministic fluid limit, and the $\sqrt{b/\eta}$-scaled fluctuations satisfy a functional CLT with noise covariance $G^{\star}$; near a nondegenerate optimum the limit is Ornstein–Uhlenbeck, and its Lyapunov covariance scaled by $\eta/b$ matches the linearized discrete recursion at leading order.
+Under a curvature–noise compatibility condition $\mu_{F}>0$, we prove $1/N$ mean-square upper bounds and an i.i.d.
+parametric Fisher van Trees lower bound of the same rate order, with oracle-complexity guarantees depending on an effective dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$.
+Numerical experiments verify the identification and confirm the Lyapunov predictions in direct SGD.
+Keywords: Stochastic gradient descent; simulation optimization; sampling-design; mini-batch covariance; diffusion limits; Ornstein–Uhlenbeck approximation; Fisher and Godambe information; oracle complexity; batch-size control
 1 Introduction
@@
 In these settings, sampling effort is a decision variable: each iterate trades off additional replications (or scenarios) against the number of updates that can be executed under a fixed budget.
-Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and plain SGD frequently behaves as if it were curvature-aware despite using only first-order information.
-Classical stochastic approximation (SA) theory provides the analytical machinery—diffusion limits, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—for analyzing such recursions once the gradient-noise covariance is specified.
-We do not claim novelty for that machinery.
-Our contribution is logically prior: we identify a structural property of the stochastic-gradient oracle that SA typically treats as exogenous, namely that the mini-batch gradient covariance is pinned down as $b^{-1}G^{\star}(\theta)$ under exchangeable sampling, where $G^{\star}(\theta)$ is the projected covariance of per-sample gradients.
-In likelihood settings this matrix reduces to Fisher information at the correctly specified point $\theta^{\star}$; for general losses it is the projected gradient covariance that, together with the Hessian $H^{\star}$, induces sandwich/Godambe geometry.
-Once $G^{\star}(\theta)$ is identified, the diffusion approximation for constant-step SGD is determined rather than modeled; the natural metric for rates is statistical rather than Euclidean; and batch size admits principled design rules under a fixed sampling budget.
-The subsequent Lyapunov analysis, rate bounds, and complexity guarantees follow by applying standard tools in this identified geometry.
+Several widely observed phenomena depend on this trade-off: small mini-batches often dominate large batches in wall-clock efficiency; constant-step SGD exhibits steady-state “error floors” and implicit regularization; and the linearized local equilibrium of plain SGD exhibits directional structure related to the local curvature–noise geometry, despite the drift being unpreconditioned.
+Classical stochastic approximation and diffusion-approximation analyses—small-step SDE surrogates, Ornstein–Uhlenbeck linearization, Lyapunov stationary covariance, $\sqrt{t}$-CLT scaling—are well-developed OR tools for analyzing such recursions once the gradient-noise covariance is specified (Kushner and Yin, 2003; Harrison, 1985; Whitt, 2002).
+We do not reprove that machinery.
+Our contribution is a sampling-design reinterpretation of the noise input: under exchangeable fresh-sampling mini-batches, the conditional mini-batch covariance is determined by the sampling mechanism at $b^{-1}$ times the per-sample gradient covariance, a statement that is elementary but carries structural consequences for operational decisions about sampling budgets.
+Writing the identified matrix as $G^{\star}(\theta)$—projected Fisher information at the correctly specified point, and more generally the sandwich partner of the Hessian $H^{\star}$—lets us analyze constant-step SGD in a fixed statistical metric rather than a Euclidean one, with the diffusion coefficient determined by the sampling design rather than postulated.
+Under a curvature–noise compatibility condition $\mu_{F}>0$ (Assumption 6.2; not implied by $H^{\star},F^{\star}\succ 0$ alone), the resulting local rates in the Fisher case match an i.i.d.
+van Trees lower bound in rate order, and batch size becomes a statistical design variable subject to a sampling budget.
 1.1 Motivation: Simulation Optimization as Sampling Control
@@
 In practice, under fixed sampling budgets, small $b$ often wins.
-The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar.
-Its covariance is a matrix that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment.
-Our results show that, to leading order, this covariance scales as $b^{-1}G^{*}(\theta)$, where $G^{*}(\theta)$ is the projected covariance of per-sample gradients.
-For correctly specified likelihood objectives, $G^{*}(\theta^{\star})$ reduces to projected Fisher information at the well-specified point; for general losses $G^{*}(\theta)$ is the projected gradient covariance that, together with $H^{\star}$, induces sandwich/Godambe geometry.
-This provides the statistical foundation for interpreting SGD as a controlled diffusion whose effective temperature is $\tau=\eta/b$.
-This sampling-control perspective is not limited to simulation optimization.
-Three canonical MOR problem classes share the same structure: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency.
-In each case, sampling effort is explicit, $b$ is small relative to $d$, and the question “how many samples per iterate?” is operationally binding.
-The theory developed below applies uniformly to all three classes.
-1.2 Core Insights
-Before turning to formal statements, we summarize the main idea informally.
-When SGD computes a mini-batch gradient by averaging $b$ per-sample gradients, the resulting noise is not generic randomness.
-Its covariance—the matrix describing which directions are noisy and which are quiet—is determined by the loss function and the data-generating process.
-This has three immediate consequences:
-  1. 1.
-    Mini-batch noise is anisotropic.
-Doubling the batch size reduces noise equally in all directions, but the shape of the noise ellipsoid is fixed by the identified covariance $G^{\star}$, with Fisher information recovered locally under correct specification.
-  2. 2.
-    The natural metric for local analysis is statistical, not Euclidean.
-Since the noise covariance is shaped by $G^{\star}$ near $\theta^{\star}$, measuring convergence in the corresponding statistical metric gives tighter and more interpretable bounds than the Euclidean metric.
-  3. 3.
-    Batch size is a design variable, not just a hyperparameter.
-Under a fixed oracle budget $N=Tb$, batch size controls the effective temperature $\tau=\eta/b$ of the induced diffusion.
-It governs the stationary covariance through the Lyapunov balance, making $b$ a design lever for effort allocation.
-The remainder of this section makes these claims precise.
-1.3 What We Show
-(i) Identification from sampling theory.
-Under exchangeable sampling (and simple random mini-batching), the mini-batch gradient satisfies
-$\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mathcal{F}\bigr)\;=\;\frac{1}{b}\,G^{*}(\theta)\;+\;o(b^{-1}),$ (3)
-where $G^{*}(\theta)$ is the projected gradient covariance (Section 4). For correctly specified likelihood losses, $G^{*}(\theta^{\star})=F^{\star}(\theta^{\star})$.
-(ii) Diffusion and Lyapunov structure.
-In the small-stepsize regime with $\tau=\eta/b$ held fixed, the SGD recursion admits a diffusion approximation whose volatility is determined by the identified covariance $G^{*}(\theta)$ (Section 5).
-Near a nondegenerate critical point, the limiting SDE linearizes to an OU process whose stationary covariance $\Sigma_{\infty}$ solves the Lyapunov equation
-$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}\;=\;\tau\,G^{*}(\theta^{\star}),$ (4)
-making explicit that steady-state fluctuations are governed by the ratio of curvature to identified noise.
-(iii) Matching rates and minimax lower bounds.
-We analyze SGD performance in the frozen statistical geometry at $\theta^{\star}$.
-We establish $1/T$ upper bounds for mean-square error in the Fisher metric and match them with information-theoretic lower bounds via van Trees, showing that the rate and its dependence on intrinsic geometry are tight within the stated model class (Sections 6 and 7).
-(iv) Oracle complexity in statistical conditioning.
-The oracle complexity bounds scale with a statistical condition number and an effective dimension, isolating statistical difficulty from Euclidean conditioning.
-This clarifies when classical Hessian-based condition numbers can be misleading for statistical objectives.
-(v) Batch size as an OR design variable.
-The framework above connects batch-size design to the identified statistical geometry; Section 9 discusses extensions to misspecification, dependence, and operational effort allocation.
+The key reason is geometric: the noise in $g_{B}(\theta)$ is not scalar but a matrix-valued covariance that encodes which directions are intrinsically informative and which are weakly identified by the underlying experiment.
+The same sampling-control perspective applies to three canonical MOR problem classes: (i) simulation optimization (queuing, inventory, revenue management), where each mini-batch gradient requires $b$ simulation replications and the replication budget is the binding constraint; (ii) stochastic programming and SAA, where scenario-based gradient estimates are drawn from an empirical distribution and the number of scenarios per iterate is a design variable subject to a total scenario budget; and (iii) online M-estimation in service systems, where per-transaction loss gradients arrive in a stream and the batch window trades off estimation quality against decision latency.
+In each case, sampling effort is explicit, $b$ is typically small relative to $d$, and the question “how many samples per iterate?” is operationally binding.
+1.2 Contributions and scope
+Under exchangeable fresh-sampling mini-batches, the conditional mini-batch covariance given the de Finetti directing measure $\mu$ is $\mathrm{Cov}(g_{B}(\theta)\mid\mu)=b^{-1}G_{\mu}(\theta)$ (Theorem 4.3).
+Under identifiability with a single data-generating distribution, the unconditional covariance reads
+$\mathrm{Cov}\!\bigl(g_{B}(\theta)\bigr)\;=\;\frac{1}{b}\,G(\theta),$ (3)
+and, after projection onto the identifiable tangent space, we denote the resulting population object $G^{\star}(\theta):=\Pi_{\theta}G(\theta)\Pi_{\theta}^{\top}$ (Section 4).
+All downstream results are stated in terms of $G^{\star}$.
+For correctly specified likelihood losses, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$; the finite-dataset-reuse case adds a (1b/n)(1-b/n) correction and is treated separately (Lemma 4.1, Proposition 4.5).
+Once $G^{*}(\theta)$ is identified, three consequences follow within the classical stochastic-approximation toolkit.
+(a) At fixed batch size, constant-step SGD has a deterministic fluid limit (the gradient-flow ODE), and its $\sqrt{b/\eta}$-scaled fluctuations obey a functional CLT converging to a linear SDE with noise covariance $G^{\star}$; near a nondegenerate optimum this is an OU process, and its stationary covariance, multiplied back by $\eta/b$, is the discrete-time Lyapunov benchmark of Proposition 5.13 (Section 5).
+(b) Mean-square error in the Fisher/identified metric admits a $1/N$ upper bound with an intrinsic effective dimension $d_{\mathrm{eff}}$ and statistical condition number $\kappa_{F}$ (Section 6), matched in rate order by an i.i.d.
+parametric Fisher van Trees lower bound (Proposition 6.8); geometry dependence appears only in the upper bound.
+(c) A high-probability oracle-complexity statement follows, with a squared-gradient stationarity rate of order $N^{-1/2}$ (Theorem 7.3).
 Main Results (informal).
-Theorem 4.3 — Identification of mini-batch covariance under exchangeable sampling: $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)+o(b^{-1})$.
-Theorem 5.4 — Diffusion approximation with volatility determined by $G^{\star}$.
+Theorem 4.3 — Covariance identification under exchangeable fresh sampling: $\mathrm{Cov}(g_{B}\mid\mu)=b^{-1}G_{\mu}(\theta)$; under identifiability and after projection onto the identifiable tangent space, the population covariance is $\mathrm{Cov}(g_{B})=b^{-1}G^{\star}(\theta)$.
+Theorem 5.7 — Fluid limit for the raw iterates and functional CLT for $\sqrt{b/\eta}$-scaled fluctuations, with noise covariance determined by $G^{\star}$.
 Thm.
-5.6, Cor.
-5.7 — OU linearization at $\theta^{\star}$; stationary covariance solves a Lyapunov equation.
+5.10, Cor.
+5.11 — OU equilibrium at $\theta^{\star}$ on the fluctuation scale; linearized discrete-recursion stationary covariance on the original scale, $\Sigma_{\eta}^{\mathrm{lin}}=(\eta/b)\Sigma_{U}+o(\eta/b)$ (Prop.
+5.13).
 Thm.
 6.5, Prop.
-6.10 — Matching $1/(Tb)$ upper and lower bounds in the frozen Fisher metric.
-Theorem 7.3 — Oracle complexity $N\!=\!\widetilde{\Theta}(\kappa_{F}\,d_{\mathrm{eff}}/\varepsilon^{2})$ for Fisher-dual stationarity.
+6.8 — $1/(Tb)$ upper bound and rate-order-matching i.i.d.
+Fisher lower bound.
+Theorem 7.3 — Squared-gradient stationarity at rate $N^{-1/2}$; norm stationarity at $N^{-1/4}$.
 1.3 Organization
@@
 Section 5 builds the diffusion approximation, OU linearization, and discrete-time bridge.
-Section 6 proves matching upper and lower bounds in Fisher geometry, and Section 7 states oracle complexity.
+Section 6 proves an upper bound and a rate-order-matching i.i.d.
+Fisher lower bound, and Section 7 states oracle complexity.
 Section 8 provides numerical validation, and Section 9 discusses extensions and practical implications.
@@
 In those works the diffusion covariance is either postulated, estimated from running iterates, or justified heuristically through moment matching.
-Our covariance identification removes this modeling degree of freedom: under exchangeable mini-batching, the diffusion matrix is determined as $G^{\star}(\theta)$ (Theorem 5.4), and the local OU regime inherits its Lyapunov balance at temperature $\tau$ (Corollary 5.7).
-Because the identification is logically prior to the discretization corrections of Li et al.
-(2017) or the Hessian-scaling regime of Smith et al.
-(2021), it is what enables the subsequent rate-matching bounds.
+Our covariance identification removes this modeling degree of freedom.
+At fixed batch size, we do not claim a nondegenerate diffusion limit of the raw iterate path—its martingale variation is of order $\eta/b$ and vanishes—but we prove a fluid limit plus a functional CLT on the $\sqrt{b/\eta}$-rescaled fluctuations, with noise covariance $C^{\star}{C^{\star}}^{\top}=G^{\star}$ (Theorem 5.7).
+The local OU regime and the Lyapunov balance then appear as the equilibrium specialization on the fluctuation scale, with $\tau=\eta/b$ recovered on the original iterate scale (Theorem 5.10, Corollary 5.11).
+This is the standard stochastic-approximation functional CLT architecture (Kushner and Yin, 2003) applied to the fresh-sampling covariance identified in Section 4, and is what fixes the noise input to the subsequent rate analysis.
 Scaling regimes.
-Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2024).
+Recent rigorous diffusion-limit work studies joint regimes where $b$ grows with $d$ (Paquette et al., 2025).
 Our analysis is complementary and closer to OR operating constraints: we keep $d$ fixed and treat $b$ as a design variable subject to a budget, with particular attention to the common regime $b\ll d$ where the matrix geometry of $G^{\star}(\theta)$ controls transient behavior and risk.
@@
 Classical oracle lower bounds are stated in Euclidean norms under smoothness and strong convexity (Nemirovski et al., 2009; Agarwal et al., 2012), which may not reflect the intrinsic difficulty of statistical objectives.
-We state both upper and lower bounds in the identified statistical metric via a van Trees argument (Theorem 6.5 and Proposition 6.10), with complexity governed by statistical curvature and $d_{\operatorname{eff}}$ rather than ambient dimension.
+We give an upper bound in the identified statistical metric that surfaces dependence on $\kappa_{F}$ and $d_{\operatorname{eff}}$ (Theorem 6.5), and match its $1/N$ rate with an i.i.d.
+van Trees lower bound in the parametric Fisher setting (Proposition 6.8); the lower bound is at the level of rate order and does not certify the geometry dependence.
 OR context: sampling budgets.
@@
 Mini-batch noise covariance Exogenous / isotropic / heuristic $\Gamma$ Identified from sampling: $\mathrm{Cov}(g_{B})\propto b^{-1}G^{\star}(\theta)$; Fisher is a special case Sec. 4; Thm. 4.3
-Diffusion matrix for SGD Postulated $\mathbf{C}(\theta)$ in SDEs Diffusion coefficient determined by sampling theory: $G^{\star}(\theta)$ Sec. 5; Thm. 5.4
-OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU linearization with Lyapunov balance driven by $G^{\star}$ at temperature $\tau=\eta/b$ Sec. 5; Cor. 5.7
+Continuous-time limit Postulated SDE with ad hoc $\mathbf{C}(\theta)$ Fluid limit (ODE) for raw iterates; functional CLT for $\sqrt{b/\eta}$-scaled fluctuations with diffusion $G^{\star}(\theta)$ Sec.
+5; Thm.
+5.7
+OU / Lyapunov equilibrium Heuristic OU with ad hoc covariance OU fluctuation limit at $\theta^{\star}$; Lyapunov balance $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$ on fluctuation scale; linearized original-scale benchmark $\Sigma_{\eta}^{\mathrm{lin}}=(\eta/b)\Sigma_{U}+o(\eta/b)$ (Prop.
+5.13 / Cor.
+5.11) Sec.
+5; Cor.
+5.11
 Metric for rates Euclidean distance; $\kappa_{H}$ dominates Identified statistical metric is the natural risk Sec. 6; Thm. 6.5
-Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.10
+Lower bounds Euclidean oracle lower bounds Information-theoretic lower bound in Fisher metric (van Trees) Sec. 6; Prop. 6.8
 Dimension notion Ambient $d$ Effective dimension $d_{\operatorname{eff}}$ induced by $F^{\star}$ / $G^{\star}$ Sec. 6; Def. (40)
@@
 given a directing measure).
-All lower bounds in Section 6 are stated for a general oracle transcript with bounded predictable information (Assumption 6.9); the i.i.d.
-sampling model is recovered as a special case.
+The lower bound in Section 6 is stated for the i.i.d.
+parametric Fisher setting; adaptive-oracle extensions are discussed in Remark 6.11.
 At iteration $t$, the algorithm selects a mini-batch $B_{t}\subseteq\{1,\dots,n\}$ of size $b_{t}$ uniformly without replacement and forms the mini-batch gradient
@@
 When $T_{\theta}=\mathbb{R}^{d}$, $\Pi_{\theta}=I$ and $G^{\star}=G$, $F^{\star}=F$.
-Remark 3.6 (Terminological convention: “Fisher/Godambe”).
-Throughout the paper we write “Fisher/Godambe” as shorthand for the following identification: in the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ (Fisher information); for general losses, $G^{\star}(\theta^{\star})$ is the projected gradient covariance whose inverse is the Godambe information.
+Remark 3.6 (Terminological convention).
+Throughout the paper, the identified noise matrix is $G^{\star}(\theta)$ (the projected gradient covariance).
+In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ (projected Fisher information).
+For general losses, $G^{\star}(\theta^{\star})$ is not itself the Godambe information; rather, paired with the Hessian $H^{\star}$ it induces the classical sandwich covariance $H^{\star-1}G^{\star}H^{\star-1}$ and, by inversion, the Godambe information $J_{\mathrm{God}}=H^{\star\top}G^{\star-1}H^{\star}$.
 Every theorem statement holds with $G^{\star}$ in the general case and specializes to $F^{\star}$ under correct specification.
-When we write “Fisher metric” or “Fisher-dual norm,” the general-loss reader should substitute the corresponding Godambe/sandwich object.
+When we write “Fisher metric” or “Fisher-dual norm,” the general-loss reader should substitute the corresponding object defined by $G^{\star}$; the paired Godambe/sandwich interpretation always requires $H^{\star}$ as well.
 Remark 3.7 (What enters later sections).
@@
 $d_{\operatorname{eff}}(F^{\star})$ Effective dimension / stable rank of $F^{\star}$
-4 Fisher/Godambe Alignment of Mini-Batch Noise
+4 Covariance Identification for Mini-Batch Noise
 Mini-batch gradient noise is not an arbitrary covariance: it is determined by (i) the sampling design used to form mini-batches and (ii) the intrinsic covariance of the per-sample estimating equation $\psi(\theta;Z)$ (Fisher information in the well-specified likelihood case, the projected gradient covariance more generally).
@@
 Lemma 4.1 is the clean MOR-style decomposition: sampling design contributes the scalar finite-population correction 1b/n1-b/n, while the geometry lives in $S_{n}(\theta)$.
-4.2 From exchangeability to Godambe/Fisher alignment (population geometry)
+4.2 From exchangeability to the population covariance
 We now convert the dataset-dependent covariance $S_{n}(\theta)$ into a population-level object.
@@
 and, when identifiability requires it, the projected version $G_{\mu}^{\star}(\theta):=\Pi_{\theta}G_{\mu}(\theta)\Pi_{\theta}^{\top}$.
-Theorem 4.3 (Godambe alignment under exchangeability).
+Theorem 4.3 (Covariance identification under exchangeability).
 Assume $\mathbb{E}_{\mu}\|\psi(\theta;Z)\|_{2}^{2}<\infty$.
@@
 $\mathrm{Cov}\!\bigl(g_{B}(\theta)\mid\mu\bigr)\;=\;\frac{1}{b}\,G_{\mu}(\theta).$ (21)
+When the conditional mean $\mathbb{E}[\psi(\theta;Z)\mid\mu]$ does not vary with $\mu$ (identifiability under a single data-generating distribution), the unconditional covariance reads $\mathrm{Cov}(g_{B}(\theta))=b^{-1}G(\theta)$ where $G(\theta):=\mathbb{E}_{\mu}[G_{\mu}(\theta)]$; after projection onto the identifiable tangent space (Section 3), we denote this $G^{\star}(\theta)=\Pi_{\theta}G(\theta)\Pi_{\theta}^{\top}$.
+The remainder of the paper works with $G^{\star}$.
 Interpretation.
@@
 […]
-Figure 1: Isotropic noise (left) vs. Fisher-aligned noise (right) at the same total variance. Batch size rescales temperature but not shape.
+Figure 1: Isotropic noise (left, dashed circle) vs. Fisher-aligned noise (right, solid ellipse) at the same total variance. Batch size rescales temperature but not shape.
 Remark 4.7 (Specialization to Fisher alignment).
@@
 Neither Lemma 4.1 nor Theorem 4.3 assumes Gaussian noise. Gaussianity enters later only as a diffusion approximation (martingale CLT) when we pass from the discrete recursion to an OU limit.
-5 Continuous-Time Diffusion Approximation and the OU Regime
-This section links the discrete-time SGD recursion to a continuous-time SDE whose diffusion covariance is determined by the mini-batch sampling model and the intrinsic noise geometry derived in Section 4.
-Specifically, $G^{\star}(\theta)$ was defined in Section 3 and shown in Theorem 4.3 to govern the leading-order mini-batch covariance.
-Near a nondegenerate critical point, the SDE linearizes to an Ornstein–Uhlenbeck (OU) process, and its stationary covariance is characterized by a Lyapunov equation.
-The resulting “temperature” parameter $\tau=\eta/b$ is the scalar that converts sampling effort into diffusion strength.
-Diffusion approximations are a standard tool in OR for reducing complex stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
-The stationary covariance and risk expressions derived below depend only on the local second-moment structure of the oracle noise and do not require independence across oracle calls; a martingale-difference oracle (Assumption 6.9) suffices.
-5.1 SGD decomposition and effective temperature
-Fix constant stepsize $\eta>0$ and batch size $b$. Let
+5 Diffusion Limits: Fluid Scale, Fluctuation Scale, and the OU Regime
+This section links the discrete-time SGD recursion to a continuous-time stochastic process whose diffusion coefficient is determined by the mini-batch covariance $G^{\star}(\theta)$ identified in Section 4.
+The link is not a raw diffusion limit at fixed batch size: the martingale quadratic variation of the raw interpolation is of order $\eta/b$ and vanishes as $\eta\downarrow 0$, so the raw process converges to the deterministic gradient-flow ODE.
+The nondegenerate stochastic limit appears only after centering around that ODE and rescaling by $\sqrt{b/\eta}$, yielding a fluctuation process whose limiting covariance structure is exactly $G^{\star}$.
+Near a nondegenerate optimum, this fluctuation process is an Ornstein–Uhlenbeck (OU) process whose stationary covariance $\Sigma_{U}$ satisfies a fluctuation-scale Lyapunov equation.
+On the original iterate scale, the linearized discrete recursion of Proposition 5.13 admits a unique stationary covariance $\Sigma_{\eta}^{\mathrm{lin}}$ that matches $(\eta/b)\Sigma_{U}$ at leading order in $\eta$, recovering the paper’s “effective temperature” $\tau=\eta/b$ on the correct scale; we do not claim an invariant law for the nonlinear SGD iterates.
+Diffusion approximations are a standard tool in OR for reducing stochastic systems to tractable Brownian models (Harrison, 1985; Whitt, 2002); the present setting is analogous, with the sampling mechanism playing the role of the arrival process.
+Throughout this section we work in the fresh-sampling / population conditioning regime of Theorem 4.3: at each iterate the mini-batch is a simple random sample drawn with fresh randomness from the population (equivalently, conditional on the de Finetti directing measure $\mu$).
+The fixed-dataset reuse case with (1b/n)(1-b/n) correction belongs to Lemma 4.1 and Proposition 4.5 and is not used below.
+5.1 SGD recursion and the martingale-noise decomposition
+Fix a batch size $b\in\mathbb{N}$ and a step size $\eta>0$. Let
 $h(\theta):=\nabla L(\theta)=\mathbb{E}[\psi(\theta;Z)],\qquad\psi(\theta;z):=\nabla_{\theta}\ell(\theta;z),$
-and write the mini-batch gradient estimator as
-$g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t}(\theta_{t}),\qquad\mathbb{E}[\xi_{t}(\theta_{t})\mid\mathcal{F}_{t}]=0,$
-where $\mathcal{F}_{t}$ is the natural filtration. Then SGD is
-$\theta_{t+1}=\theta_{t}-\eta\,h(\theta_{t})-\eta\,\xi_{t}(\theta_{t}).$ (24)
-By Theorem 4.3 (and its Fisher specialization in Remark 4.7), the conditional covariance of the noise obeys
-$\mathrm{Cov}\!\bigl(\xi_{t}(\theta)\mid\mathcal{F}_{t}\bigr)=\frac{1}{b}\,G^{\star}(\theta)+r_{b,n}(\theta),\qquad\|r_{b,n}(\theta)\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1}),$ (25)
-uniformly in a local neighborhood (the $O(n^{-1})$ term accounts for finite-population corrections and covariance-estimation bias; it is negligible whenever $b=o(n)$).
-We call $\tau:=\eta/b$ the effective temperature:
+and write the mini-batch gradient estimator as $g_{t}(\theta_{t})=h(\theta_{t})+\xi_{t}(\theta_{t})$ with $\mathbb{E}[\xi_{t}(\theta_{t})\mid\mathcal{F}_{t}]=0$, where $\mathcal{F}_{t}$ is the natural filtration.
+The SGD recursion is
+$\theta_{k+1}^{\eta}=\theta_{k}^{\eta}-\eta\,h(\theta_{k}^{\eta})-\eta\,\xi_{k+1}^{\eta},$ (23)
+with the superscript $\eta$ emphasizing the dependence of the iterate path on the step size. Under the fresh-sampling conditioning,
+$\mathbb{E}\!\bigl[\xi_{k+1}^{\eta}(\xi_{k+1}^{\eta})^{\top}\mid\mathcal{F}_{k}^{\eta}\bigr]\;=\;\frac{1}{b}\,G^{\star}(\theta_{k}^{\eta})+r_{k+1}^{\eta},$ (24)
+with remainder $r_{k+1}^{\eta}$ satisfying the uniform bound stated in Assumption 5.2 below. We call $\tau:=\eta/b$ the effective temperature:
 $\tau:=\eta/b,$ (25)
-which will appear as the single scalar factor multiplying the diffusion covariance in the continuous-time limit.
-5.2 A diffusion approximation (Euler–Maruyama viewpoint)
-Let $s=t\eta$ denote algorithmic time and define the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ and linearity on $t\eta,(t+1)\eta]$.
-Assumption 5.1 (Local regularity and localization).
-On a neighborhood $\mathcal{U}$ of $\theta^{\star}$:
+the scalar that sets the size of the stationary fluctuations on the original iterate scale (Corollary 5.11).
+The covariance identification of Section 4 determines the noise geometry, but at fixed batch size $b$ the raw interpolation of SGD on algorithmic time $t=k\eta$ has martingale quadratic variation of order $\eta/b$, which vanishes as $\eta\downarrow 0$.
+Accordingly, the raw process converges to the deterministic gradient-flow ODE.
+The correct nondegenerate stochastic limit is obtained by centering around that ODE and rescaling by $\sqrt{b/\eta}$, yielding a fluctuation process whose limiting covariance structure is exactly $G^{\star}$.
+5.2 Assumptions for the fluid and fluctuation limits
+Assumption 5.1 (Local regularity and ODE flow).
+There is an open neighborhood $\mathcal{U}\subseteq\mathbb{R}^{d}$ such that:
   1. (i)
-    $h(\theta):=\nabla L(\theta)$ is locally Lipschitz;
+    $h\in C^{1}(\mathcal{U};\mathbb{R}^{d})$ and $\nabla h$ is locally Lipschitz on $\mathcal{U}$;
   2. (ii)
-    $\nabla^{2}L(\theta)$ is locally Lipschitz (i.e., there exists $L_{H}<\infty$ such that $\|\nabla^{2}L(\theta)-\nabla^{2}L(\theta^{\prime})\|_{\operatorname{op}}\leq L_{H}\|\theta-\theta^{\prime}\|_{2}$ for all $\theta,\theta^{\prime}\in\mathcal{U}$);
+    $G^{\star}:\mathcal{U}\to\mathbb{S}_{+}^{d}$ is continuous;
   3. (iii)
-    $\sup_{\theta\in\mathcal{U}}\mathbb{E}\|\psi(\theta;Z)\|^{4}<\infty$ (bounded fourth moment; implied by Assumption 3.8(b));
-  4. (iv)
-    $H^{\star}:=\nabla^{2}L(\theta^{\star})\succ 0$ (local strong convexity; Assumption 3.9);
-  5. (v)
-    for each fixed horizon $T>0$, $\Pr(\theta_{t}\in\mathcal{U}\ \forall\,t\leq T/\eta)\to 1$ as $\eta\downarrow 0$ (localization; standard in stochastic approximation).
-Assumption 5.2 (Asymptotic conditional covariance and Lindeberg).
-Uniformly on $\mathcal{U}$, $\mathrm{Cov}(\xi_{t}(\theta)\mid\mathcal{F}_{t})=b^{-1}G^{\star}(\theta)+o(b^{-1})$ and the conditional Lindeberg condition holds for the martingale increments $\eta\xi_{t}(\theta_{t})$ (e.g., via a conditional CLT under exchangeability as in Section 4).
-Remark 5.3 (Verifiable sufficient conditions).
-Assumption 5.2 holds if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|_{2}^{2+\delta}]<\infty$ for some $\delta>0$ (guaranteed by Assumption 3.8(b)), $G^{\star}(\theta)$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence.
-Theorem 5.4 (Diffusion approximation).
-Suppose Assumptions 5.15.2 hold.
-For fixed batch size $b\geq 1$ and constant step size $\eta>0$, define the effective temperature $\tau:=\eta/b$.
-For sufficiently small $\eta$, the SGD recursion admits a diffusion approximation: the piecewise-linear interpolation $\bar{\theta}^{\eta}(\cdot)$ converges weakly on $D([0,T];\mathbb{R}^{d})$ as $\eta\downarrow 0$ to the solution $\theta(\cdot)$ of the SDE
-$d\theta_{s}=-h(\theta_{s})\,ds+\sqrt{\tau}\,C^{\star}(\theta_{s})\,dW_{s},\qquad C^{\star}(\theta)\,{C^{\star}(\theta)}^{\top}=G^{\star}(\theta),$ (27)
-where $C^{\star}$ is, e.g., the symmetric PSD square root of $G^{\star}$,
-where $W_{s}$ is standard $d$-dimensional Brownian motion.
-In particular, the diffusion has instantaneous covariance $\tau\,G^{\star}(\theta_{s})\,ds$.
-In the correctly specified likelihood case, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ at $\theta^{\star}$, so the diffusion covariance is locally Fisher-structured near the optimum.
+    the ODE $\dot{\bar{\theta}}(t)=-h(\bar{\theta}(t))$ with $\bar{\theta}(0)=\theta_{0}$ has a unique solution on $[0,T]$ remaining in a compact set $K\Subset\mathcal{U}$.
+Assumption 5.2 (Martingale noise and conditional covariance).
+The sequence $(\xi_{k+1}^{\eta},\mathcal{F}_{k+1}^{\eta})$ is a martingale difference, and the conditional covariance admits the expansion (24) with
+$\sup_{k\leq T/\eta}\|r_{k+1}^{\eta}\|_{\operatorname{op}}\;\xrightarrow{\,\mathbb{P}\,}\;0\qquad\text{as }\eta\downarrow 0.$
+Assumption 5.3 (Conditional fourth-moment bound on $K$).
+There is a constant $C_{K}<\infty$ such that
+$\sup_{\eta>0}\,\sup_{k\leq T/\eta}\,\mathbb{E}\!\left[\|\xi_{k+1}^{\eta}\|^{4}\,\mathbf{1}_{\{\theta_{k}^{\eta}\in K\}}\,\big|\,\mathcal{F}_{k}^{\eta}\right]\;\leq\;C_{K}\quad\text{a.s.}$
+Assumption 5.4 (Localization).
+$\Pr\!\bigl(\theta_{k}^{\eta}\in K\text{ for all }k\leq T/\eta\bigr)\to 1$ as $\eta\downarrow 0$.
+Assumption 5.5 (Initial fluctuations converge).
+$\theta_{0}^{\eta}\to\theta_{0}$, and $U_{0}^{\eta}:=\sqrt{b/\eta}\,(\theta_{0}^{\eta}-\theta_{0})\Rightarrow U_{0}$ for some random variable $U_{0}$.
+Remark 5.6 (Verifiability).
+Assumptions 5.25.3 hold if $\sup_{\theta\in\mathcal{U}}\mathbb{E}[\|\psi(\theta;Z)\|^{4}]<\infty$ (guaranteed by Assumption 3.8(b)), $G^{\star}$ is continuous on $\mathcal{U}$, and the mini-batch is drawn by simple random sampling from an exchangeable sequence.
+Assumption 5.4 is standard in stochastic approximation (Kushner–Yin style) and holds, e.g., when the ODE is globally stable and the initial iterate is in the basin.
+5.3 Fluid limit and fluctuation theorem
+Theorem 5.7 (Fluid limit and fluctuation limit at fixed batch size).
+Fix $T>0$ and $b\in\mathbb{N}$, and suppose Assumptions 5.15.5 hold.
+Let $\hat{\theta}^{\eta}(t):=\theta_{\lfloor t/\eta\rfloor}^{\eta}$ be the step interpolation of the SGD iterates (23), and let $\bar{\theta}$ solve the ODE in Assumption 5.1(iii).
+Then:
+  1. (a)
+    Fluid limit.
+    $\sup_{t\in[0,T]}\bigl\|\hat{\theta}^{\eta}(t)-\bar{\theta}(t)\bigr\|\;\xrightarrow{\,\mathbb{P}\,}\;0.$
+  2. (b)
+    Fluctuation FCLT. The rescaled fluctuation process
+    $U^{\eta}(t)\;:=\;\sqrt{\tfrac{b}{\eta}}\,\bigl(\hat{\theta}^{\eta}(t)-\bar{\theta}(t)\bigr)$
+    converges weakly in $D([0,T];\mathbb{R}^{d})$ (equivalently, in $C([0,T];\mathbb{R}^{d})$, as the limit is continuous) to the unique weak solution $U$ of the linear SDE
+[…]
+    where $W$ is a standard $d$-dimensional Brownian motion and $C^{\star}(\theta){C^{\star}(\theta)}^{\top}=G^{\star}(\theta)$ (e.g., the symmetric PSD square root of $G^{\star}$).
+Equivalently, on the original iterate scale,
+$\hat{\theta}^{\eta}(t)\;=\;\bar{\theta}(t)\;+\;\sqrt{\tfrac{\eta}{b}}\,U_{t}\;+\;o_{\mathbb{P}}\!\Bigl(\sqrt{\tfrac{\eta}{b}}\Bigr),$ (27)
+uniformly on $[0,T]$.
 Proof.
-Step 1 (scaled interpolation).
-Write the recursion in signed-noise form $\theta_{t+1}=\theta_{t}-\eta h(\theta_{t})+\eta\,\zeta_{t+1}$ with $\zeta_{t+1}:=-\xi_{t}$, so $(\zeta_{t+1},\mathcal{F}_{t+1})$ is a martingale difference.
-Define the piecewise-linear interpolation $\bar{\theta}^{\eta}(s)$ on algorithmic time $s=t\eta$ by $\bar{\theta}^{\eta}(t\eta)=\theta_{t}$ with linear interpolation.
-Step 2 (drift convergence).
-By local Lipschitzness of $h$ (Assumption 5.1) and the localization condition, $\sum_{k<s/\eta}\eta\,h(\theta_{k})\to\int_{0}^{s}h(\theta(u))\,du$ uniformly on compacts, by standard ODE-approximation theory (Kushner and Yin, 2003, Chapter 5).
-Step 3 (predictable quadratic variation).
-Define $M^{\eta}(s):=\sum_{k<s/\eta}\eta\,\zeta_{k+1}$.
-By the alignment theorem (Theorem 4.3), $\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\frac{1}{b}\,G^{\star}(\theta_{k})+R_{k}$ with $\|R_{k}\|_{\operatorname{op}}\leq(c/b)\rho_{k}$ and $\rho_{k}\to 0$ locally.
-Using $\eta^{2}/b=\eta\tau$:
-$\langle M^{\eta}\rangle(s)=\sum_{k<s/\eta}\eta^{2}\mathbb{E}[\zeta_{k+1}\zeta_{k+1}^{\top}\mid\mathcal{F}_{k}]=\tau\sum_{k<s/\eta}\eta\,G^{\star}(\theta_{k})+o(1)\;\to\;\tau\!\int_{0}^{s}\!G^{\star}(\theta(u))\,du$
-in probability, uniformly on $[0,S]$.
-Step 4 (Lindeberg and tightness).
-The fourth moment bound (Assumption 3.8(b)) yields $\sum_{k<S/\eta}\mathbb{E}[\|\eta\zeta_{k+1}\|^{2}\mathbf{1}\{\|\eta\zeta_{k+1}\|>\varepsilon\}\mid\mathcal{F}_{k}]\leq CS\eta^{3}m_{4}/(\varepsilon^{2}b^{2})\to 0$, verifying the conditional Lindeberg condition.
-Tightness of $\{\bar{\theta}^{\eta}\}$ in $D([0,S];\mathbb{R}^{d})$ follows from Aldous’ criterion and the Lipschitz-controlled drift.
-Step 5 (martingale FCLT).
-By the martingale central limit theorem (Ethier and Kurtz, 1986, Theorem 7.1.4), $M^{\eta}(\cdot)\Rightarrow M(\cdot)$ weakly in $D([0,S];\mathbb{R}^{d})$ where $M$ is a continuous martingale with quadratic variation $\tau\int_{0}^{s}G^{\star}(\theta(u))\,du$.
-Lévy’s characterization gives a Brownian motion $W$ such that $M(s)=\int_{0}^{s}\sqrt{\tau}\,C^{\star}(\theta(u))\,dW_{u}$.
-Combining drift and martingale convergence identifies the limit as the unique solution of (27).
+We follow the standard stochastic-approximation FCLT architecture (cf.
+Kushner and Yin, 2003, Ch.
+8), tailored to the fresh-sampling covariance (24).
+Write $t_{k}:=k\eta$ and $\bar{\theta}_{k}:=\bar{\theta}(t_{k})$; on the localization event of Assumption 5.4, which has probability tending to $1$, the iterates stay in $K$ for all $k\leq T/\eta$, and we work under this event without further comment.
+Step 1 (fluid limit).
+Define the raw error $e_{k}^{\eta}:=\theta_{k}^{\eta}-\bar{\theta}_{k}$ and the ODE local truncation error $\rho_{k+1}^{\eta}:=\bar{\theta}_{k+1}-\bar{\theta}_{k}+\eta\,h(\bar{\theta}_{k})$.
+Local Lipschitzness of $h$ on $K$ gives $\sup_{k\leq T/\eta}\|\rho_{k+1}^{\eta}\|\leq C_{T}\eta^{2}$ for some $C_{T}<\infty$.
+Subtracting the Euler step for $\bar{\theta}$ from (23),
+$e_{k+1}^{\eta}\;=\;e_{k}^{\eta}\;-\;\eta\bigl(h(\theta_{k}^{\eta})-h(\bar{\theta}_{k})\bigr)\;-\;\eta\,\xi_{k+1}^{\eta}\;-\;\rho_{k+1}^{\eta}.$ (28)
+Let $N_{n}^{\eta}:=\eta\sum_{j=0}^{n-1}\xi_{j+1}^{\eta}$; this is an $\mathcal{F}_{k}^{\eta}$-martingale.
+Set $\bar{G}_{K}:=\sup_{\theta\in K}\mathrm{Tr}G^{\star}(\theta)<\infty$ (finite by continuity of $G^{\star}$ on the compact set $K$).
+By Doob’s maximal inequality and Assumptions 5.25.3,
+$\mathbb{E}\!\Bigl[\sup_{n\leq T/\eta}\|N_{n}^{\eta}\|^{2}\Bigr]\;\leq\;4\,\eta^{2}\sum_{j<T/\eta}\mathbb{E}\|\xi_{j+1}^{\eta}\|^{2}\;\leq\;4\,\eta^{2}\cdot\frac{T}{\eta}\cdot\frac{\bar{G}_{K}+o(1)}{b}\;=\;O(\eta/b).$
+Hence $\sup_{n\leq T/\eta}\|N_{n}^{\eta}\|\xrightarrow{\mathbb{P}}0$. Using local Lipschitzness of $h$ on $K$ (constant $L_{K}$) in (28),
+$\|e_{n}^{\eta}\|\;\leq\;\|e_{0}^{\eta}\|\;+\;L_{K}\eta\sum_{j=0}^{n-1}\|e_{j}^{\eta}\|\;+\;\sup_{m\leq n}\|N_{m}^{\eta}\|\;+\;C_{T}\,T\,\eta,$
+and discrete Grönwall gives
+$\sup_{n\leq T/\eta}\|e_{n}^{\eta}\|\;\leq\;e^{L_{K}T}\Bigl(\|e_{0}^{\eta}\|+\sup_{m\leq T/\eta}\|N_{m}^{\eta}\|+C_{T}T\eta\Bigr)\;\xrightarrow{\,\mathbb{P}\,}\;0,$
+since $\|e_{0}^{\eta}\|\to 0$ by Assumption 5.5. This establishes part (a).
+Step 2 (rescaled recursion).
+Set $U_{k}^{\eta}:=\sqrt{b/\eta}\,e_{k}^{\eta}$.
+Local Lipschitzness of $\nabla h$ on $K$ gives the Taylor expansion $h(\theta_{k}^{\eta})-h(\bar{\theta}_{k})=A_{k}^{\eta}e_{k}^{\eta}+q_{k}^{\eta}$ with $A_{k}^{\eta}:=\nabla h(\bar{\theta}_{k})$ and $\|q_{k}^{\eta}\|\leq L^{\prime}_{K}\|e_{k}^{\eta}\|^{2}$.
+Multiplying (28) by $\sqrt{b/\eta}$ and using $\eta\sqrt{b/\eta}=\sqrt{\eta b}$,
+$U_{k+1}^{\eta}\;=\;U_{k}^{\eta}\;-\;\eta\,A_{k}^{\eta}\,U_{k}^{\eta}\;+\;\Delta M_{k+1}^{\eta}\;+\;\varepsilon_{k+1}^{\eta},$ (29)
+with
+$\Delta M_{k+1}^{\eta}\;:=\;-\sqrt{\eta b}\,\xi_{k+1}^{\eta},\qquad\varepsilon_{k+1}^{\eta}\;:=\;-\sqrt{\eta b}\,q_{k}^{\eta}\;-\;\sqrt{\tfrac{b}{\eta}}\,\rho_{k+1}^{\eta}.$ (30)
+Step 3 (tightness scale and remainders). From Step 1, $\sup_{n}\|e_{n}^{\eta}\|=e^{L_{K}T}(\|e_{0}^{\eta}\|+\sup\|N^{\eta}\|+C_{T}T\eta)$. Multiplying by $\sqrt{b/\eta}$,
+$\sup_{n}\|U_{n}^{\eta}\|\;\leq\;e^{L_{K}T}\Bigl(\|U_{0}^{\eta}\|+\sqrt{\tfrac{b}{\eta}}\sup_{m}\|N_{m}^{\eta}\|+C_{T}T\sqrt{b\eta}\Bigr).$
+The first term is tight by Assumption 5.5; $\mathbb{E}[(b/\eta)\sup_{m}\|N_{m}^{\eta}\|^{2}]\leq 4b/\eta\cdot O(\eta/b)=O(1)$, so the second is $O_{\mathbb{P}}(1)$; the third tends to 0. Hence
+$\sup_{n\leq T/\eta}\|U_{n}^{\eta}\|=O_{\mathbb{P}}(1).$ (31)
+Using (31) in the Taylor-remainder bound,
+$\Bigl\|\sum_{k<n}\sqrt{\eta b}\,q_{k}^{\eta}\Bigr\|\;\leq\;\sqrt{\eta b}\cdot L^{\prime}_{K}\sum_{k<n}\|e_{k}^{\eta}\|^{2}\;\leq\;L^{\prime}_{K}\,T\,\sqrt{\tfrac{\eta}{b}}\,\sup_{k}\|U_{k}^{\eta}\|^{2}\xrightarrow{\mathbb{P}}0,$
+and, since $\|\rho_{k+1}^{\eta}\|\leq C_{T}\eta^{2}$,
+$\Bigl\|\sum_{k<n}\sqrt{\tfrac{b}{\eta}}\,\rho_{k+1}^{\eta}\Bigr\|\;\leq\;\frac{T}{\eta}\,C_{T}\eta^{2}\sqrt{\tfrac{b}{\eta}}\;=\;C_{T}T\sqrt{b\eta}\to 0.$
+Hence $\sup_{t\leq T}\bigl\|\sum_{k<t/\eta}\varepsilon_{k+1}^{\eta}\bigr\|\xrightarrow{\mathbb{P}}0$.
+Step 4 (martingale FCLT). Let $M^{\eta}(t):=\sum_{k<t/\eta}\Delta M_{k+1}^{\eta}=-\sqrt{\eta b}\sum_{k<t/\eta}\xi_{k+1}^{\eta}$. By (24) its predictable quadratic variation is
+$\langle M^{\eta}\rangle(t)\;=\;\eta b\sum_{k<t/\eta}\!\mathbb{E}\!\bigl[\xi_{k+1}^{\eta}(\xi_{k+1}^{\eta})^{\top}\mid\mathcal{F}_{k}^{\eta}\bigr]\;=\;\eta\sum_{k<t/\eta}G^{\star}(\theta_{k}^{\eta})\;+\;\eta b\sum_{k<t/\eta}r_{k+1}^{\eta}.$
+The remainder is bounded uniformly in $t\leq T$ by $Tb\cdot\sup_{k}\|r_{k+1}^{\eta}\|\xrightarrow{\mathbb{P}}0$ under Assumption 5.2.
+For the leading term, continuity of $G^{\star}$ on $K$ and the fluid limit $\sup_{k}\|\theta_{k}^{\eta}-\bar{\theta}_{k}\|\to 0$ (Step 1) give $\eta\sum_{k<t/\eta}\|G^{\star}(\theta_{k}^{\eta})-G^{\star}(\bar{\theta}_{k})\|\to 0$ in probability uniformly in $t$; and $\eta\sum_{k<t/\eta}G^{\star}(\bar{\theta}_{k})\to\int_{0}^{t}G^{\star}(\bar{\theta}(s))\,ds$ is a Riemann-sum approximation with continuous integrand.
+Hence
+$\langle M^{\eta}\rangle(t)\;\Rightarrow\;\int_{0}^{t}G^{\star}(\bar{\theta}(s))\,ds\qquad\text{uniformly on }[0,T].$
+For the conditional Lindeberg condition, using $\mathbf{1}_{\{x>\varepsilon\}}\leq x^{2}/\varepsilon^{2}$,
+$\sum_{k<T/\eta}\!\mathbb{E}\!\bigl[\|\Delta M_{k+1}^{\eta}\|^{2}\mathbf{1}_{\{\|\Delta M_{k+1}^{\eta}\|>\varepsilon\}}\mid\mathcal{F}_{k}^{\eta}\bigr]\;\leq\;\frac{(\eta b)^{2}}{\varepsilon^{2}}\sum_{k<T/\eta}\mathbb{E}[\|\xi_{k+1}^{\eta}\|^{4}\mid\mathcal{F}_{k}^{\eta}]\;\leq\;\frac{\eta\,b^{2}\,T\,C_{K}}{\varepsilon^{2}}\to 0.$
+The martingale FCLT (Ethier and Kurtz, 1986, Theorem 7.1.4) then gives $M^{\eta}\Rightarrow M$ in $D([0,T];\mathbb{R}^{d})$, where $M$ is a continuous Gaussian martingale with quadratic variation $\int_{0}^{\cdot}G^{\star}(\bar{\theta}(s))\,ds$.
+Lévy’s characterization yields a Brownian motion $W$ with $M_{t}=\int_{0}^{t}C^{\star}(\bar{\theta}(s))\,dW_{s}$, and $C^{\star}(\bar{\theta}(\cdot))$ can be taken as the (continuous) symmetric PSD square root of $G^{\star}(\bar{\theta}(\cdot))$.
+Step 5 (identify the limiting SDE). Summing (29) and using Step 3,
+$U^{\eta}(t)\;=\;U_{0}^{\eta}\;-\;\int_{0}^{t}\!\nabla h\!\bigl(\bar{\theta}(\lfloor s/\eta\rfloor\eta)\bigr)\,U^{\eta}(s)\,ds\;+\;M^{\eta}(t)\;+\;o_{\mathbb{P}}(1),$
+uniformly on $[0,T]$.
+Continuity of $s\mapsto\nabla h(\bar{\theta}(s))$ and tightness $\sup_{t}\|U^{\eta}(t)\|=O_{\mathbb{P}}(1)$ give $\int_{0}^{t}\nabla h(\bar{\theta}(\lfloor s/\eta\rfloor\eta))U^{\eta}(s)ds-\int_{0}^{t}\nabla h(\bar{\theta}(s))U^{\eta}(s)ds\to 0$ uniformly in probability.
+Tightness of $\{U^{\eta}\}$ in $D([0,T];\mathbb{R}^{d})$ follows from (31) and the controlled drift together with Aldous’ criterion applied to the martingale term (whose FCLT is in hand).
+Any subsequential weak limit $U$ therefore satisfies the stochastic integral equation
+$U_{t}\;=\;U_{0}\;-\;\int_{0}^{t}\nabla h(\bar{\theta}(s))\,U_{s}\,ds\;+\;\int_{0}^{t}C^{\star}(\bar{\theta}(s))\,dW_{s},$
+which is (26) in integrated form. The linear SDE (26) with continuous coefficients admits a unique weak solution, so the entire sequence converges: $U^{\eta}\Rightarrow U$.
+Step 6 (expansion).
+By construction $\hat{\theta}^{\eta}(t)-\bar{\theta}(t)=\sqrt{\eta/b}\,U^{\eta}(t)$; combining (a) and (b) gives (27).
+The asymptotic equivalence of step and piecewise-linear interpolations under Assumption 5.3 is standard and yields the same limit for either choice.
 ∎
-5.3 The OU regime near a nondegenerate critical point
-Assume $\theta^{\star}$ is a stationary point: $h(\theta^{\star})=0$, and define the local curvature $H^{\star}:=\nabla h(\theta^{\star})=\nabla^{2}L(\theta^{\star})$.
-Let $\Delta_{s}:=\theta_{s}-\theta^{\star}$ denote the centered diffusion.
+Remark 5.8 (Why the raw process is deterministic in the limit).
+At fixed $b$, the martingale variation of the raw interpolation is $\sum_{k<t/\eta}\eta^{2}\,b^{-1}G^{\star}(\theta_{k}^{\eta})\sim(\eta/b)\int_{0}^{t}G^{\star}(\bar{\theta}(s))\,ds$, which vanishes as $\eta\downarrow 0$.
+Hence the raw stochastic fluctuation is invisible at the original scale, and a nondegenerate Gaussian limit is recovered only after centering and rescaling by $\sqrt{b/\eta}$.
+This is the standard stochastic-approximation functional CLT regime; the diffusion lives on the fluctuation scale, not in the raw iterate path.
+5.4 The OU regime as the equilibrium specialization
+Suppose $\theta^{\star}$ is a stationary point: $h(\theta^{\star})=0$, and $\theta_{0}=\theta^{\star}$ so that the ODE flow is trivial, $\bar{\theta}(t)\equiv\theta^{\star}$.
+Then the drift coefficient in (26) is the constant matrix $H^{\star}:=\nabla h(\theta^{\star})=\nabla^{2}L(\theta^{\star})$, and the diffusion coefficient is the constant matrix $C^{\star}(\theta^{\star})$.
 Assumption 5.9 (Nondegeneracy at $\theta^{\star}$).
 $H^{\star}\succ 0$ and $G^{\star}(\theta^{\star})\succ 0$.
-Theorem 5.6 (OU linearization).
-Under Assumption 5.5 and the conditions of Theorem 5.4, the diffusion (27) admits the local expansion
-$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s}+\rho_{1}(\Delta_{s})\,ds+\sqrt{\tau}\,\rho_{2}(\Delta_{s})\,dW_{s},$ (28)
-where $\|\rho_{1}(\Delta)\|=o(\|\Delta\|)$ and $\|\rho_{2}(\Delta)\|_{\operatorname{op}}=o(1)$ as $\Delta\to 0$. Neglecting the higher-order remainder terms yields the OU SDE
-$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C^{\star}(\theta^{\star})\,dW_{s},\qquad C^{\star}(\theta^{\star}){C^{\star}(\theta^{\star})}^{\top}=G^{\star}(\theta^{\star}).$ (29)
+Theorem 5.10 (OU fluctuation limit at a stationary point).
+Under Assumption 5.9 and the conditions of Theorem 5.7 with $\theta_{0}=\theta^{\star}$, the fluctuation limit $U$ of Theorem 5.7 solves the Ornstein–Uhlenbeck SDE
+$dU_{t}\;=\;-H^{\star}\,U_{t}\,dt\;+\;C^{\star}(\theta^{\star})\,dW_{s},\qquad C^{\star}(\theta^{\star}){C^{\star}(\theta^{\star})}^{\top}=G^{\star}(\theta^{\star}).$ (32)
+Equivalently, on the original iterate scale, the centered iterate $\hat{\theta}^{\eta}(t)-\theta^{\star}$ admits the asymptotic expansion
+$\hat{\theta}^{\eta}(t)-\theta^{\star}\;=\;\sqrt{\tfrac{\eta}{b}}\,U_{t}\;+\;o_{\mathbb{P}}\!\bigl(\sqrt{\tfrac{\eta}{b}}\bigr).$
 Proof.
-A Taylor expansion gives $h(\theta^{\star}+\Delta)=H^{\star}\Delta+r(\Delta)$ with $\|r(\Delta)\|=o(\|\Delta\|)$.
-Continuity of $G^{\star}$ implies we can choose a continuous square root $C^{\star}(\theta)$ locally (e.g., the symmetric PSD square root), so
-$C^{\star}(\theta^{\star}+\Delta)=C^{\star}(\theta^{\star})+R_{C}(\Delta),\qquad\|R_{C}(\Delta)\|_{\operatorname{op}}=o(1)\quad(\Delta\to 0).$
-Substituting into (27) and collecting terms yields (28) with $\rho_{1}(\Delta):=-r(\Delta)$ and $\rho_{2}(\Delta):=R_{C}(\Delta)$. ∎
-Corollary 5.7 (Stationary covariance and the Lyapunov equation).
-If $H^{\star}\succ 0$, the OU process (29) is ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$.
-Since $H^{\star}\succ 0$, the matrix $H^{\star}$ is Hurwitz, so the continuous Lyapunov equation
-$H^{\star}\Sigma_{\infty}+\Sigma_{\infty}(H^{\star})^{\top}=\tau\,G^{\star}(\theta^{\star})$ (30)
-admits a unique symmetric positive definite solution $\Sigma_{\infty}$. In particular, the stationary covariance scales linearly with temperature $\tau=\eta/b$.
-Because $H^{\star}$ is symmetric positive definite, left-multiplying (30) by $(H^{\star})^{-1}$ and taking traces gives the trace identity
-$\mathrm{Tr}(\Sigma_{\infty})+\mathrm{Tr}\!\bigl((H^{\star})^{-1}\Sigma_{\infty}H^{\star}\bigr)=\tau\,\mathrm{Tr}\!\bigl((H^{\star})^{-1}G^{\star}(\theta^{\star})\bigr),$ (31)
-which reduces to $2\,\mathrm{Tr}(\Sigma_{\infty})=\tau\,\mathrm{Tr}(G^{\star}H^{\star-1})$ when $H^{\star}$ and $\Sigma_{\infty}$ commute (e.g., when the eigenbases of curvature and noise align).
+Specialize (26) to $\bar{\theta}(\cdot)\equiv\theta^{\star}$: the drift becomes $H^{\star}U_{t}$ and the diffusion becomes $C^{\star}(\theta^{\star})dW_{t}$.
+The original-scale expansion is (27) with $\bar{\theta}\equiv\theta^{\star}$.
+∎
+Corollary 5.11 (Fluctuation-scale Lyapunov equation and discrete-scale benchmark).
+Under Assumption 5.9, the OU process (32) is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{U})$, where $\Sigma_{U}$ is the unique symmetric positive definite solution of the fluctuation-scale Lyapunov equation
+$H^{\star}\,\Sigma_{U}\;+\;\Sigma_{U}\,(H^{\star})^{\top}\;=\;G^{\star}(\theta^{\star}).$ (33)
+On the original iterate scale, Proposition 5.13 provides a discrete Lyapunov covariance $\Sigma_{\eta}^{\mathrm{lin}}$ for the linearized recursion, which at leading order in $\eta$ equals $(\eta/b)\Sigma_{U}$ and satisfies
+$H^{\star}\,\Sigma_{\eta}^{\mathrm{lin}}+\Sigma_{\eta}^{\mathrm{lin}}\,(H^{\star})^{\top}\;=\;\frac{\eta}{b}\,G^{\star}(\theta^{\star})+O(\eta)\cdot\Sigma_{\eta}^{\mathrm{lin}}.$ (34)
+This is the linearized-recursion benchmark on the original iterate scale; we do not claim an invariant law for the nonlinear SGD iterates here.
+Because $H^{\star}$ is symmetric positive definite, cyclicity of trace gives the unconditional trace identity
+$2\,\mathrm{Tr}(\Sigma_{U})\;=\;\mathrm{Tr}\!\bigl((H^{\star})^{-1}G^{\star}(\theta^{\star})\bigr).$ (35)
 Proof.
-Write $B:=\sqrt{\tau}\,(G^{\star}(\theta^{\star}))^{1/2}$, so $BB^{\top}=\tau G^{\star}(\theta^{\star})$.
-Since $^{\star}\succ 0, H-H^{\star}$ is Hurwitz and the OU process is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{\infty})$.
-Apply Itô’s formula to $\Delta_{s}\Delta_{s}^{\top}$: the drift contributes $(H^{\star}\Sigma_{s}+\Sigma_{s}H^{\star\top})$ and the quadratic variation contributes $BB^{\top}$, giving $\frac{d}{ds}\Sigma_{s}=-H^{\star}\Sigma_{s}-\Sigma_{s}H^{\star\top}+BB^{\top}$.
-Setting $\frac{d}{ds}\Sigma_{s}=0$ yields (30).
-Uniqueness follows because for Hurwitz $H^{\star}$ the integral $\Sigma_{\infty}=\int_{0}^{\infty}e^{-H^{\star}s}\,(\tau G^{\star})\,e^{-H^{\star\top}s}\,ds$ converges and is the unique solution (Lemma A.3).
+Write $B:=(G^{\star}(\theta^{\star}))^{1/2}$, so $BB^{\top}=G^{\star}(\theta^{\star})$.
+Since $^{\star}\succ 0, H-H^{\star}$ is Hurwitz and the OU process is exponentially ergodic with stationary law $\mathcal{N}(0,\Sigma_{U})$.
+Applying Itô’s formula to $U_{t}U_{t}^{\top}$ and taking stationary expectations yields $(H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top})+BB^{\top}=0$, which is (33); the integral representation $\Sigma_{U}=\int_{0}^{\infty}e^{-H^{\star}s}G^{\star}(\theta^{\star})e^{-H^{\star\top}s}ds$ converges and gives uniqueness (Lemma A.1).
+Equation (34) follows from Proposition 5.13 after dividing out leading powers of $\eta$ (Remark 5.14).
+The trace identity follows by left-multiplying (33) by $(H^{\star})^{-1}$ and using $\mathrm{Tr}(H^{-1}\Sigma H)=\mathrm{Tr}(\Sigma)$.
 ∎
-Remark 5.8 (Likelihood case (Fisher-aligned diffusion)).
-If $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, then $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$ and the OU diffusion reads
-$d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,C_{F}^{\star}\,dW_{s},\qquad C_{F}^{\star}{C_{F}^{\star}}^{\top}=F^{\star}(\theta^{\star}),$
-making Fisher geometry explicit in both transient fluctuations and in the stationary Lyapunov balance (30).
+Remark 5.12 (Likelihood specialization).
+When $\ell(\theta;z)=-\log p_{\theta}(z)$ and the model is correctly specified, $G^{\star}(\theta^{\star})=F^{\star}(\theta^{\star})$.
+The fluctuation-scale Lyapunov equation (33) then reads $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=F^{\star}(\theta^{\star})$; the linearized discrete-recursion benchmark of Proposition 5.13 satisfies $\Sigma_{\eta}^{\mathrm{lin}}=(\eta/b)\Sigma_{U}+o(\eta/b)$ at leading order in $\eta$ (Corollary 5.11).
+Fisher geometry is therefore explicit both in the transient fluctuations and in the linearized stationary Lyapunov balance; we do not assert an invariant law for the nonlinear SGD iterates here.
 5.5 Discrete-time stationary covariance: an exact bridge
-The OU viewpoint is not the only consequence of the covariance identification.
-The same $G^{\star}$ governs the exact linearized discrete-time stationary covariance; the continuous-time Lyapunov equation emerges as the small-stepsize asymptotic reduction.
-This subsection makes the bridge explicit, so that the paper’s conclusions do not rest on the diffusion approximation alone.
+The same $G^{\star}$ that governs the fluctuation-scale Lyapunov equation (33) also governs the exact linearized discrete-time stationary covariance.
+The continuous-time identity (34) is the small-$\eta$ reduction of the discrete Lyapunov equation below.
 Proposition 5.13 (Discrete-time linearized stationary covariance).
@@
 $\Delta_{t+1}=(I-\eta H^{\star})\,\Delta_{t}+\eta\,\zeta_{t+1},\qquad\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]=\frac{1}{b}\,G^{\star}(\theta^{\star})+r_{t},$ (36)
-where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})+O(n^{-1})$ as in (25).
-If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), then the iterates are geometrically ergodic and the stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfies the discrete Lyapunov equation
+where $\|r_{t}\|_{\operatorname{op}}=o(b^{-1})$ as in (24).
+If $\rho(I-\eta H^{\star})<1$ (equivalently, $\eta<2/\lambda_{\max}(H^{\star})$), the linearized covariance recursion is stable and admits a unique stationary covariance $\Sigma_{\eta}:=\lim_{t\to\infty}\mathrm{Cov}(\Delta_{t})$ satisfying the discrete Lyapunov equation
 $\Sigma_{\eta}=(I-\eta H^{\star})\,\Sigma_{\eta}\,(I-\eta H^{\star})^{\top}+\frac{\eta^{2}}{b}\,G^{\star}(\theta^{\star}),$ (37)
-up to the same local remainder terms already used in the diffusion analysis.
+up to the same local remainder terms already used in the fluctuation analysis.
 Proof.
-From (32), compute $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$, using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$.
-Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the iteration converges to the unique fixed point (33).
-The spectral radius condition $\rho(I-\eta H^{\star})<1$ is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$.
+From (36), $\mathrm{Cov}(\Delta_{t+1})=(I-\eta H^{\star})\mathrm{Cov}(\Delta_{t})(I-\eta H^{\star})^{\top}+\eta^{2}\mathbb{E}[\zeta_{t+1}\zeta_{t+1}^{\top}\mid\mathcal{F}_{t}]$ using $\mathbb{E}[\zeta_{t+1}\mid\mathcal{F}_{t}]=0$.
+Under $\rho(I-\eta H^{\star})<1$, the map $\Sigma\mapsto(I-\eta H^{\star})\Sigma(I-\eta H^{\star})^{\top}+Q$ is a contraction on positive semidefinite matrices, so the covariance iteration converges to the unique fixed point (37).
+The spectral radius condition is equivalent to $0<\eta<2/\lambda_{\max}(H^{\star})$, since $H^{\star}\succ 0$.
+The statement is about the second-moment recursion for the linearized iterates; geometric ergodicity of the nonlinear SGD iterates would require additional assumptions (drift + minorization) that we do not impose.
 ∎
-Remark 5.10 (Continuous-time reduction and discrete-time corrections).
-Expanding (33) and dividing by $\eta$, one recovers $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=\frac{\eta}{b}G^{\star}+\eta H^{\star}\Sigma_{\eta}H^{\star\top}$.
-Sending $\eta\downarrow 0$ with $\tau=\eta/b$ fixed drops the $O(\eta)$ correction term and yields the continuous Lyapunov equation (30).
-The correction is governed by $\eta\lambda_{\max}(H^{\star})$: when this product is small the two benchmarks nearly coincide; for moderate stepsizes the discrete equation is the correct comparator (Section 8).
-The covariance identification $\mathrm{Cov}(\xi_{t})\approx b^{-1}G^{\star}$ is logically prior to both analyses: the discrete Lyapunov equation is its exact linearized consequence, and the OU equation is the small-$\eta$ reduction.
+Remark 5.14 (Small-stepsize reduction and the continuous limit).
+Expanding (37), $H^{\star}\Sigma_{\eta}+\Sigma_{\eta}H^{\star\top}=(\eta/b)G^{\star}+\eta\,H^{\star}\Sigma_{\eta}H^{\star\top}$.
+When $\eta\lambda_{\max}(H^{\star})$ is small, the $O(\eta)$ discretization correction is negligible, and the discrete equation reduces to the continuous-time, original-scale identity (34).
+The covariance identification $\mathbb{E}[\xi\xi^{\top}\mid\mathcal{F}]\approx b^{-1}G^{\star}$ is logically prior to both: the discrete Lyapunov equation is its exact linearized consequence, and the continuous equation is the small-$\eta$ reduction that matches the fluctuation-scale limit of Theorem 5.7 after multiplying back by $\eta/b$.
 6 Convergence Rates in Fisher Geometry
@@
 Convention.
-All rates in this section are stated in Fisher/Godambe risk $\mathbb{E}[\theta_{T}^{\top}F^{\star}\theta_{T}]$ unless explicitly noted otherwise; any Euclidean norm that appears is immediately converted.
+All rates in this section are stated in the identified statistical risk $\mathbb{E}[\theta_{T}^{\top}F^{\star}\theta_{T}]$ unless explicitly noted otherwise; any Euclidean norm that appears is immediately converted.
 6.1 Local Fisher Geometry and a Frozen-Metric Regime
@@
 Take total expectations in (41) to obtain $\mathbb{E}[V_{t+1}]\leq a_{t}\,\mathbb{E}[V_{t}]+u_{t}$ with $a_{t}:=1-2\mu_{F}\eta_{t}+c_{0}\eta_{t}^{2}$ and $u_{t}:=\eta_{t}^{2}\frac{1}{b}\mathrm{Tr}((F^{\star})^{2})+\eta_{t}^{2}\frac{c_{1}}{b}\mathrm{Tr}(F^{\star})\rho_{t}$.
-For $\eta_{t}=\eta_{0}/t$, there exists $t_{0}$ such that for $t\geq t_{0}$, $a_{t}\leq 1-\mu_{F}\eta_{0}/t$.
-Define $A_{t,T}:=\prod_{s=t}^{T-1}a_{s}$.
+Fix any $\epsilon\in(0,2\mu_{F})$.
+For $\eta_{t}=\eta_{0}/t$ and $t$ large enough that $c_{0}\eta_{0}^{2}/t\leq\epsilon\eta_{0}$, we have $a_{t}\leq 1-(2\mu_{F}-\epsilon)\eta_{0}/t$; set $\alpha:=(2\mu_{F}-\epsilon)\eta_{0}$, which satisfies $\alpha>1$ whenever $\eta_{0}>1/(2\mu_{F})$ and $\epsilon$ is chosen small enough.
+Define $A_{t,T}:=\prod_{s=t}^{T-1}a_{s}$ (Lemma A.2).
 Unrolling gives $\mathbb{E}[V_{T}]\leq A_{1,T}V_{1}+\sum_{t=1}^{T-1}A_{t+1,T}u_{t}$.
-Standard product bounds yield $A_{1,T}\leq CT^{-2\mu_{F}\eta_{0}}$ and $A_{t+1,T}\leq C(t/T)^{\mu_{F}\eta_{0}}$.
-Then, since $\mu_{F}\eta_{0}>1/2$, $\sum_{t\leq T}A_{t+1,T}\eta_{t}^{2}=O(1/T)$, giving the leading $(\mathrm{Tr}((F^{\star})^{2})/b)(1/T)$ term.
-The remainder term is controlled because $\sum_{t}\eta_{t}^{2}\rho_{t}<\infty$ and $A_{t+1,T}\leq 1$, yielding an additional $O((\mathrm{Tr}(F^{\star})/b)(1/T))$ contribution.
-This proves (38).
+With the consistent exponent $\alpha$, the product bounds are $A_{1,T}\leq C\,T^{-\alpha}$ and $A_{t+1,T}\leq C\,(t/T)^{\alpha}$.
+Since $\alpha>1$, the standard Robbins–Monro calculation gives $\sum_{t=1}^{T-1}A_{t+1,T}\,\eta_{t}^{2}=O(1/T)$, yielding the leading $(\mathrm{Tr}((F^{\star})^{2})/b)(1/T)$ term.
+The remainder contribution is controlled because $\sum_{t}\eta_{t}^{2}\rho_{t}<\infty$ and $A_{t+1,T}\leq 1$, adding $O((\mathrm{Tr}(F^{\star})/b)(1/T))$.
+The transient term $A_{1,T}V_{1}=O(T^{-\alpha})$ appears in the bound (42) with exponent $\alpha\leq 2\mu_{F}\eta_{0}$; sending $\epsilon\downarrow 0$ gives the exponent $2\mu_{F}\eta_{0}$ stated in (42) up to an arbitrarily small loss.
 The reduction to (6.5) follows from $\mathrm{Tr}((F^{\star})^{2})\leq\lambda_{\max}(F^{\star})\mathrm{Tr}(F^{\star})$.
@@
 6.3 Lower Bound
-The lower bound proceeds in three layers of increasing generality.
-We first state the cleanest result in the parametric likelihood/Fisher case, then interpret it through oracle information, and finally discuss the extension to general losses.
-6.3.1 Layer 1: Parametric Fisher-case lower bound.
+We state the clean lower bound in the local parametric Fisher setting with i.i.d.
+observations.
+Extensions to adaptive oracles and general losses require additional machinery and are left as discussion (Remarks 6.116.12).
 Assumption 6.7 (Local parametric regularity for van Trees).
@@
     there exists a smooth prior density $\pi$ supported on $\Theta_{\operatorname{loc}}$ with finite Fisher information $J_{\pi}:=\int\|\nabla\log\pi(\theta)\|^{2}\pi(\theta)\,d\theta<\infty$.
-Theorem 6.8 (Fisher-case lower bound via van Trees).
-Suppose Assumption 6.7 holds and the estimator $\hat{\theta}_{N}$ is based on $N$ i.i.d. observations from $P_{\theta^{\star}}$. Then
-$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (40)
-for a universal constant $c>0$.
+Proposition 6.8 (Local i.i.d. Fisher-case lower bound via van Trees).
+Suppose Assumption 6.7 holds.
+Consider the local estimation model in which, for each $\theta\in\Theta_{\operatorname{loc}}$, one observes $N$ i.i.d.
+samples from $P_{\theta}$ and forms an estimator $\hat{\theta}_{N}$ (a measurable function of the samples).
+Then
+$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c_{\mathrm{loc}}\,d}{N},$ (44)
+for a constant $c_{\mathrm{loc}}>0$ that depends on the local regularity of $F(\theta)$ on $\Theta_{\operatorname{loc}}$ and on the prior regularity $J_{\pi}$ from Assumption 6.7(iv).
+In particular, if each SGD iteration uses a fresh mini-batch of size $b$ (so $N=Tb$ fresh-sample gradient calls), the Fisher-metric risk is bounded below by $c_{\mathrm{loc}}\,d/(Tb)$ on the i.i.d.
+population oracle.
 Proof.
-This is the special case of Proposition 6.10 below with i.i.d.
-observations (so the oracle information bound holds with equality).
-We give the self-contained argument in the proof of Proposition 6.10.
+Standard van Trees argument (Gill and Levit, 1995).
+By Assumption 6.7(ii), the per-coordinate local Fisher scale $\bar{f}:=\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathrm{Tr}F(\theta)/d<\infty$ is $N$-independent.
+Fix a smooth prior $\pi$ on $\Theta_{\operatorname{loc}}$ with finite $J_{\pi}$ (Assumption 6.7(iv)); the van Trees inequality for the Fisher-weighted Bayes risk gives $r_{\pi}(\hat{\theta}_{N})\geq d/(N\bar{f}+J_{\pi}/d)$, and for $N\geq N_{0}:=J_{\pi}/(d\bar{f})$ the denominator is at most $2N\bar{f}$, yielding $r_{\pi}(\hat{\theta}_{N})\geq c_{\mathrm{loc}}\,d/N$ with $c_{\mathrm{loc}}:=1/(2\bar{f})>0$ depending only on the local regularity of $F(\theta)$ on $\Theta_{\operatorname{loc}}$ and independent of both $N$ and $d$.
+Minimax risk dominates Bayes risk for any such prior, proving (44).
 ∎
-6.3.2 Layer 2: Oracle-information interpretation.
-The heart of the lower bound is not i.i.d. sampling per se, but bounded predictable information growth. We now restate the result in terms of an abstract oracle transcript.
-Assumption 6.9 (Oracle–information structure).
-The oracle produces observations $Y_{t}=\psi(\theta_{t};Z_{t})$, where:
-  1. (i)
-    $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, with filtration $\mathcal{F}_{t}:=\sigma(Y_{1},\dots,Y_{t})$;
-  2. (ii)
-    Conditional unbiasedness: $\mathbb{E}_{\theta^{\star}}[Y_{t}\mid\mathcal{F}_{t-1}]=0$;
-  3. (iii)
-    Predictable information bound: $\mathbb{E}_{\theta^{\star}}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq\mathcal{I}_{1}(\theta^{\star})$, a.s., where $\mathcal{I}_{1}(\theta^{\star}):=F(\theta^{\star})=F^{\star}$ in the likelihood case;
-  4. (iv)
-    Uniform moment control: $\sup_{t}\mathbb{E}_{\theta^{\star}}\|Y_{t}\|^{4}<\infty$.
-Under Assumption 6.9, the predictable quadratic variation of the oracle transcript satisfies $\sum_{t=1}^{N}\mathbb{E}[Y_{t}Y_{t}^{\top}\mid\mathcal{F}_{t-1}]\preceq N\mathcal{I}_{1}(\theta^{\star})$, so the total information available to any $\mathcal{F}_{N}$-measurable estimator is bounded by $N\mathcal{I}_{1}$.
-Proposition 6.10 (Lower bound under predictable information budget).
-Suppose Assumption 6.7 holds for the parametric family $\{P_{\theta}\}$ and the oracle satisfies Assumption 6.9 with $\mathcal{I}_{1}(\theta^{\star})=F^{\star}$.
-Additionally, assume that each conditional law of $Z_{t}$ given $\mathcal{F}_{t-1}$ has a density differentiable in $\theta$, so that the conditional Fisher information $I_{t}(\theta\mid\mathcal{F}_{t-1})$ exists and satisfies $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$ a.s.
-Then every $\mathcal{F}_{N}$-measurable estimator $\hat{\theta}_{N}$ satisfies
-$\inf_{\hat{\theta}_{N}}\;\sup_{\theta\in\Theta_{\operatorname{loc}}}\mathbb{E}_{\theta}\!\left[\|\hat{\theta}_{N}-\theta\|_{F(\theta)}^{2}\right]\;\geq\;\frac{c\,d}{N},$ (41)
-for a universal constant $c>0$.
-In particular, if one SGD iteration uses a mini-batch of size $b$ (so $N=Tb$ oracle calls), then $\inf_{\hat{\theta}_{T}}\sup_{\theta}\mathbb{E}[\|\hat{\theta}_{T}-\theta\|_{F(\theta)}^{2}]\geq c\,d/(Tb)$.
-Proof.
-The argument extends Theorem 6.8 from i.i.d.
-observations to the adaptive oracle transcript.
-The van Trees bound requires an upper bound on the Fisher information $I_{N}(\theta)$ of the joint likelihood of $(Y_{1},\ldots,Y_{N})$ with respect to $\theta$.
-Three ingredients deliver this bound.
-Step 1 (Adaptivity does not inflate per-step information).
-At step $t$, the query $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, but $Z_{t}\sim P_{\theta}$ is drawn independently of $\mathcal{F}_{t-1}$ given $\theta$.
-Hence the conditional distribution of $Z_{t}$ given $\mathcal{F}_{t-1}$ depends on $\theta$ only through $\{P_{\theta}\}$, not through the past transcript.
-Step 2 (Second-moment bound == per-step Fisher bound).
-Under correct specification, $Y_{t}=-s_{\theta}(Z_{t})$ is the negative score.
-The conditional Fisher information of $Z_{t}$ given $\mathcal{F}_{t-1}$ about $\theta$ is $I_{t}(\theta\mid\mathcal{F}_{t-1})=\mathbb{E}_{\theta}[s_{\theta}s_{\theta}^{\top}\mid\mathcal{F}_{t-1}]=F(\theta)$—exactly the score second-moment matrix.
-So Assumption 6.9(iii) directly bounds $I_{t}(\theta^{\star}\mid\mathcal{F}_{t-1})\preceq F^{\star}$.
-To extend from $\theta^{\star}$ to all $\theta\in\Theta_{\operatorname{loc}}$, continuity of $F(\theta)$ (Assumption 6.7(ii)) gives $F(\theta)\preceq(1+\epsilon)F^{\star}$ on a sufficiently small $\Theta_{\operatorname{loc}}$; the factor $(1+\epsilon)$ is absorbed into the universal constant $c$.
-Step 3 (Chain rule and conclusion). The chain rule for conditional Fisher information applied to the sequential factorization of the joint likelihood gives
-$I_{N}(\theta)=\sum_{t=1}^{N}\mathbb{E}_{\theta}\!\bigl[I_{t}(\theta\mid\mathcal{F}_{t-1})\bigr]\preceq N\,(1{+}\epsilon)\,F^{\star}.$
-Substituting into van Trees with prior $\pi$ and sending $\epsilon\to 0$ yields (41). ∎
-Remark 6.11 (Beyond i.i.d. oracles).
-The classical i.i.d.
-setting satisfies Assumption 6.9 with equality in (iii).
-The assumption also accommodates exchangeable or conditionally i.i.d.
-sampling schemes (e.g., de Finetti mixtures) provided the per-step conditional information is bounded by $\mathcal{I}_{1}$.
-What is required is a predictable information bound on the oracle transcript, not independence across calls.
-The critical logical step is that the van Trees inequality applies to the joint transcript likelihood, whose total Fisher information is controlled by summing the per-step conditional bounds.
-Remark 6.12 (What enters the lower bound).
-The lower bound (41) involves $d$ (ambient dimension) and $N$ (oracle budget).
-The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ enter the upper bound and oracle complexity (Theorems 6.5 and 7.3) but not the lower bound, which is stated in terms of $d$ to give the sharpest minimax comparison.
-When $d_{\mathrm{eff}}\ll d$, the gap between upper and lower bounds tightens correspondingly.
-6.3.3 Layer 3: General-loss / Godambe extension.
-Remark 6.13 (Extension to general losses).
-In the general-loss case, the same lower-bound form $\Omega(d/N)$ is recovered under local asymptotic conditions that identify the relevant sandwich geometry.
-Specifically, if the loss satisfies a local asymptotic normality (LAN) condition at $\theta^{\star}$ with information matrix $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$, then the van Trees argument carries through with $F^{\star}$ replaced by $G^{\star}(\theta^{\star})$ and the dual norm adjusted accordingly.
-The LAN condition is standard for smooth M-estimators under regularity conditions comparable to those already imposed in Section 3; see Le Cam (1986) and van der Vaart (1998).
-We state this extension as a remark rather than a theorem because the full verification of LAN for the adaptive oracle transcript of Assumption 6.9 requires additional technical conditions (uniform local asymptotic equicontinuity) that we do not develop here.
-Corollary 6.14 (Matching rates (up to constants)).
-Under Theorem 6.5 and the oracle model of Assumption 6.9,
-$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]=\Theta\!\left(\frac{1}{Tb}\right)\quad\text{equivalently}\quad\Theta\!\left(\frac{1}{N}\right)\text{ per oracle call},\quad N=Tb.$
+Remark 6.9 (Noise-covariance condition used in the upper bounds).
+The upper bounds of Theorem 6.5 and the oracle complexity of Theorem 7.3 use the conditional-covariance bound $\mathbb{E}[\xi_{t}\xi_{t}^{\top}\mid\mathcal{F}_{t}]\preceq b^{-1}F^{\star}$, which is exactly the fresh-sampling specialization of Theorem 4.3 in the correctly specified likelihood case.
+No structure beyond this covariance identification is required.
+Remark 6.10 (Rate matching, not geometry matching).
+Proposition 6.8 establishes a lower bound of rate order $\Omega(d/N)$ in the i.i.d.
+parametric Fisher setting.
+The quantities $d_{\mathrm{eff}}$ and $\kappa_{F}$ appear only in the upper bound and oracle complexity (Theorems 6.5 and 7.3), not in the lower bound.
+Accordingly, we claim rate-order matching in the local Fisher case; we do not claim that the lower bound matches the geometry dependence of the upper bound.
+Remark 6.11 (Adaptive-oracle and general-loss extensions).
+Two natural extensions of Proposition 6.8 are left to future work.
+(a) Adaptive oracles.
+A sequential oracle transcript $Y_{t}=\psi(\theta_{t};Z_{t})$ with a predictable per-step Fisher-information bound should yield (44) through a sequential van Trees argument, provided one establishes a uniform chain-rule bound on the joint transcript likelihood (requiring regularity conditions such as uniform local asymptotic equicontinuity that we do not develop here).
+(b) General losses.
+A local asymptotic normality (LAN) condition at $\theta^{\star}$ with Godambe information $J_{\mathrm{God}}(\theta^{\star})=H^{\star\top}G^{\star-1}H^{\star}$ would produce a sandwich analogue of (44); see Le Cam (1986) and van der Vaart (1998).
+Remark 6.12 (Legacy labels for cross-references).
+These labels are retained for backward compatibility; the i.i.d. Fisher lower bound is Proposition 6.8, and the general-loss extension is discussed in Remark 6.11.
+Corollary 6.13 (Matching rate order in the local Fisher case).
+Under Theorem 6.5 and the noise identification of Theorem 4.3 in the correctly specified likelihood setting,
+$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\right]\;=\;\Theta\!\left(\frac{1}{N}\right),\qquad N=Tb,$
+at the level of rate order. Dependence on intrinsic geometry ($d_{\mathrm{eff}}$, $\kappa_{F}$) appears only in the upper bound.
 7 Oracle Complexity
-This section turns the Fisher-geometry mean-square bounds of Section 6 into an explicit oracle-complexity guarantee for achieving $\varepsilon$-stationarity.
-The steps are standard—smoothness descent, averaged-gradient bound, martingale concentration (cf.
-Nesterov, 2004, Chapter 2)—but we express the result in the natural Fisher dual norm and keep the statistical conditioning explicit.
+This section records a standard high-probability conversion of the Fisher-geometry mean-square bounds of Section 6 into an oracle-complexity guarantee for $\varepsilon$-stationarity in the Fisher dual norm.
+The proof is a textbook descent-lemma plus martingale concentration argument (cf.
+Nesterov, 2004; Wainwright, 2019, Ch.
+2); the novelty is not the complexity theorem itself but the fact that, once the mini-batch covariance is identified as $b^{-1}F^{\star}$ (Theorem 4.3), the resulting constants read as intrinsic statistical quantities ($\kappa_{F}$, $d_{\mathrm{eff}}$) rather than as Euclidean conditioning.
 Throughout, we work locally near $\theta^{\star}$ and freeze the metric $F^{\star}:=F(\theta^{\star})\succ 0$ (for general losses, replace $F^{\star}$ by $G^{\star}(\theta^{\star})$; see Remark 7.5).
@@
 7.3 High-probability oracle complexity
-Theorem 7.3 (Oracle complexity for Fisher-dual stationarity).
+Theorem 7.3 (High-probability oracle complexity under the identified noise covariance).
 Assume the local regularity conditions of Section 3 hold on a neighborhood $\mathcal{U}$ of $\theta^{\star}$, and suppose that on $\mathcal{U}$ the mini-batch noise is Fisher-aligned in the sense that, for some batch size $b\geq 1$,
@@
 $\min_{0\leq t\leq T-1}\;\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C\left(\frac{L(\theta_{0})-L(\theta^{\star})}{\eta\,T\,\lambda_{\min}(F^{\star})}\;+\;\frac{\eta}{b}\,\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\right)\log\!\frac{1}{\delta}.$ (48)
-In particular, choosing $(\eta,T)$ so that the right-hand side is $\leq\varepsilon^{2}$ implies $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}\leq\varepsilon$ with probability at least $1-\delta$.
-Equivalently, in terms of oracle calls $N:=Tb$ (sample-gradient evaluations), one sufficient scaling (up to constants and ignoring the transient term $\frac{L(\theta_{0})-L(\theta^{\star})}{\eta T\lambda_{\min}(F^{\star})}$) is
-$N\;=\;\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})}{\varepsilon^{2}}\,\log\!\frac{1}{\delta}\right).$ (46)
+In particular, balancing the two terms on the right-hand side over $\eta$ yields
+$\min_{0\leq t\leq T-1}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C^{\prime}\sqrt{\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\,(L(\theta_{0})-L(\theta^{\star}))}{N\,\lambda_{\min}(F^{\star})}}\,\log\!\frac{1}{\delta},$
+where $N=Tb$.
+Hence the squared Fisher-dual stationarity gap decays as $N^{-1/2}$, and to achieve $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}\leq\varepsilon$ with probability at least $1-\delta$ it suffices that
+$N\;=\;\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})\,(L(\theta_{0})-L(\theta^{\star}))}{\lambda_{\min}(F^{\star})\,\varepsilon^{4}}\,\log^{2}\!\frac{1}{\delta}\right).$ (49)
+The $\varepsilon^{-4}$ rate is the standard nonconvex-SGD rate for squared-gradient stationarity at level $\varepsilon^{2}$; it is not the $\varepsilon^{-2}$ rate of (strongly) convex stochastic optimization.
 Proof.
-Step 1 (descent inequality).
-By $M$-smoothness and $\eta\leq 1/(4M)$, expand $L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L,\theta_{t+1}-\theta_{t}\rangle+\frac{M}{2}\|\theta_{t+1}-\theta_{t}\|^{2}$ with $\theta_{t+1}-\theta_{t}=-\eta(\nabla L(\theta_{t})+\xi_{t})$.
-After expanding and using $\eta\leq 1/(4M)$, so $1-M\eta/2\geq 1-1/8=7/8$ and $\eta M\leq 1/4$:
-$L(\theta_{t+1})\leq L(\theta_{t})-\tfrac{3\eta}{4}\|\nabla L(\theta_{t})\|_{2}^{2}+\eta(1+M\eta)\langle\nabla L(\theta_{t}),\xi_{t}\rangle+\tfrac{M\eta^{2}}{2}\|\xi_{t}\|_{2}^{2}.$
+Step 1 (descent inequality, correct sign).
+By $M$-smoothness, $L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\tfrac{M}{2}\|\theta_{t+1}-\theta_{t}\|^{2}$ with $\theta_{t+1}-\theta_{t}=-\eta(\nabla L(\theta_{t})+\xi_{t})$.
+Expanding $\|\nabla L(\theta_{t})+\xi_{t}\|^{2}=\|\nabla L(\theta_{t})\|^{2}+2\langle\nabla L(\theta_{t}),\xi_{t}\rangle+\|\xi_{t}\|^{2}$,
+$L(\theta_{t+1})\;\leq\;L(\theta_{t})\;-\;\eta\bigl(1-\tfrac{M\eta}{2}\bigr)\|\nabla L(\theta_{t})\|^{2}\;-\;\eta(1-M\eta)\,\langle\nabla L(\theta_{t}),\xi_{t}\rangle\;+\;\tfrac{M\eta^{2}}{2}\|\xi_{t}\|^{2}.$
+Under $\eta\leq 1/(4M)$, $1-M\eta/2\geq 7/8$ and $1-M\eta\geq 3/4$. Denote $\alpha:=1-M\eta\in[3/4,1)$.
 Step 2 (define martingale).
-Let $\Delta M_{t+1}:=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ and $M_{T}:=\sum_{t=0}^{T-1}\Delta M_{t+1}$; since $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$, this is a martingale.
-Step 3 (sum the descent). Summing from $t=0$ to T1T-1:
-$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}}{2}\,Q_{T},\quad\Delta L:=L(\theta_{0})-L(\theta^{\star}),$
-where $Q_{T}:=\sum_{t<T}\|\xi_{t}\|_{2}^{2}$.
-To replace the random $Q_{T}$ by a deterministic bound we need concentration, not just conditional moments.
-Under conditional sub-Gaussianity, $\|\xi_{t}\|_{2}^{2}$ is conditionally sub-exponential with $\mathbb{E}[\|\xi_{t}\|_{2}^{2}\mid\mathcal{F}_{t}]\leq b^{-1}\mathrm{Tr}(F^{\star})$; Bernstein’s inequality for sums of conditionally sub-exponential random variables (see, e.g., Wainwright 2019, Cor.
-2.21) gives, with probability at least $1-\delta$, $Q_{T}\leq(T/b)\mathrm{Tr}(F^{\star})+c\sqrt{T\,d}\,(\|F^{\star}\|_{\operatorname{op}}/b)\sqrt{\log(1/\delta)}.$ Since $M\eta^{2}/2\leq\eta/8$ (as $\eta\leq 1/(4M)$), the deviation is $O(\eta\sqrt{T/b}\,\sqrt{d}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)}\,/b)$, which is lower-order relative to the bias term $\Delta L/(\eta T)$ in the final bound.
-Absorbing it into the constant $C$ (and taking a union bound over the two $\delta$-events, i.e.
-replacing $\delta$ by $\delta/2$ throughout):
-$\tfrac{3\eta}{4}\,S_{T}\;\leq\;\Delta L+C_{1}M_{T}+\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})\log\!\tfrac{2}{\delta},$
-with probability at least $1-\delta$.
-Step 4 (sub-Gaussian martingale concentration).
-Each increment satisfies $\Delta M_{t+1}=\eta\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ where $\nabla L(\theta_{t})$ is $\mathcal{F}_{t}$-measurable.
-By the conditional sub-Gaussianity assumption, the scalar $\langle v,\xi_{t}\rangle$ with $v:=\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|_{2}$ has sub-Gaussian parameter $\sigma_{t}^{2}\leq\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $c_{t}^{2}:=\eta^{2}\|\nabla L(\theta_{t})\|_{2}^{2}\|F^{\star}\|_{\operatorname{op}}/b$.
-By the standard exponential supermartingale bound for sums of conditionally sub-Gaussian increments (see, e.g., Wainwright 2019, Theorem 2.19), with probability at least $1-\delta$, $M_{T}\leq\sqrt{2\sum_{t<T}c_{t}^{2}\,\log(1/\delta)}=\eta\sqrt{2(\|F^{\star}\|_{\operatorname{op}}/b)\,S_{T}\,\log(1/\delta)}$, where $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}$.
-Step 5 (solve for gradient bound). Substituting the Step 4 tail bound into Step 3 gives a quadratic inequality in $\sqrt{S_{T}}$; solving:
-$\min_{t<T}\|\nabla L(\theta_{t})\|_{2}^{2}\;\leq\;\frac{S_{T}}{T}\;\leq\;C\!\left(\frac{\Delta L}{\eta T}+\frac{\eta}{b}\mathrm{Tr}(F^{\star})\right)\!\log\!\frac{1}{\delta}.$
+Let $\Delta M_{t+1}:=-\eta\alpha\,\langle\nabla L(\theta_{t}),\xi_{t}\rangle$ and $M_{T}:=\sum_{t=0}^{T-1}\Delta M_{t+1}$; since $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t}]=0$, $(M_{t})$ is a martingale adapted to $(\mathcal{F}_{t})$.
+Write $\Delta L:=L(\theta_{0})-L(\theta^{\star})$, $S_{T}:=\sum_{t<T}\|\nabla L(\theta_{t})\|^{2}$, and $Q_{T}:=\sum_{t<T}\|\xi_{t}\|^{2}$.
+Summing the Step-1 inequality from $t=0$ to T1T-1,
+$\tfrac{7\eta}{8}\,S_{T}\;\leq\;\Delta L\;+\;M_{T}\;+\;\tfrac{M\eta^{2}}{2}\,Q_{T}.$ (50)
+Step 3 ($Q_{T}$ tail via a Hanson–Wright-type bound).
+Under conditional sub-Gaussianity of $\xi_{t}$ with variance proxy $\Sigma_{t}\preceq b^{-1}F^{\star}$, the Hanson–Wright inequality for sub-Gaussian vectors (Vershynin, 2018, Thm.
+6.2.1) gives, for an absolute constant $c_{0}$, $\Pr(\,|\|\xi_{t}\|^{2}-\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t}]|>s\mid\mathcal{F}_{t})\leq 2\exp(-c_{0}\min\{s^{2}b^{2}/\|F^{\star}\|_{F}^{2},\,sb/\|F^{\star}\|_{\operatorname{op}}\})$; so $\|\xi_{t}\|^{2}-\mathbb{E}[\|\xi_{t}\|^{2}\mid\mathcal{F}_{t}]$ is conditionally sub-exponential with parameters $(\|F^{\star}\|_{F}^{2}/b^{2},\|F^{\star}\|_{\operatorname{op}}/b)$.
+Applying Bernstein’s inequality (Wainwright, 2019, Ch.
+2) over $t<T$ gives, with probability at least $1-\delta/2$,
+$Q_{T}\;\leq\;\frac{T}{b}\,\mathrm{Tr}(F^{\star})\;+\;c_{1}\Bigl(\tfrac{\|F^{\star}\|_{F}}{b}\sqrt{T\log(2/\delta)}+\tfrac{\|F^{\star}\|_{\operatorname{op}}}{b}\log(2/\delta)\Bigr).$ (51)
+Since $\|F^{\star}\|_{F}^{2}\leq d\,\|F^{\star}\|_{\operatorname{op}}^{2}$, the deviation term is $O(b^{-1}\sqrt{d\,T}\,\|F^{\star}\|_{\operatorname{op}}\sqrt{\log(1/\delta)})$, lower order than the leading $(T/b)\mathrm{Tr}(F^{\star})$ when $T\gtrsim d\log(1/\delta)$.
+Step 4 (two-sided martingale concentration).
+The scalar $\langle\nabla L(\theta_{t})/\|\nabla L(\theta_{t})\|,\,\xi_{t}\rangle$ is conditionally sub-Gaussian with parameter at most $\|F^{\star}\|_{\operatorname{op}}/b$, so $\Delta M_{t+1}$ is conditionally sub-Gaussian with parameter $v_{t}^{2}:=\eta^{2}\alpha^{2}\,\|\nabla L(\theta_{t})\|^{2}\,\|F^{\star}\|_{\operatorname{op}}/b$.
+The standard two-sided Azuma–Hoeffding/Freedman-type bound for conditionally sub-Gaussian martingale sums (Wainwright, 2019, Thm.
+2.19) gives, with probability at least $1-\delta/2$,
+$|M_{T}|\;\leq\;\sqrt{2\sum_{t<T}v_{t}^{2}\,\log(4/\delta)}\;=\;\eta\alpha\,\sqrt{\frac{2\|F^{\star}\|_{\operatorname{op}}}{b}\,S_{T}\,\log(4/\delta)}.$ (52)
+Step 5 (combine). On the intersection of the events in (51)–(52) (probability $\geq 1-\delta$ by union bound), substitute into (50) and use $|M_{T}|$ in place of $M_{T}$:
+$\tfrac{7\eta}{8}\,S_{T}$ $\;\leq\;\Delta L\;+\;\eta\sqrt{\tfrac{2\|F^{\star}\|_{\operatorname{op}}}{b}\,S_{T}\,\log(4/\delta)}\;+\;\tfrac{M\eta^{2}T}{2b}\,\mathrm{Tr}(F^{\star})$
+$\qquad\;+\;c_{2}\,\tfrac{M\eta^{2}\|F^{\star}\|_{\operatorname{op}}}{b}\Bigl(\sqrt{T\log(2/\delta)}+\log(2/\delta)\Bigr).$ (53)
+The last term scales as $\eta^{2}\sqrt{T}/b$ (up to log factors), strictly lower order than the leading variance term $\eta^{2}T\mathrm{Tr}(F^{\star})/b$ once $T\gtrsim\log(1/\delta)$, and is absorbed into the constant $C$ below.
+Solving the resulting quadratic inequality in $\sqrt{S_{T}}$ for $S_{T}/T\geq\min_{t<T}\|\nabla L(\theta_{t})\|^{2}$ yields
+$\min_{t<T}\|\nabla L(\theta_{t})\|^{2}\;\leq\;C\!\left(\frac{\Delta L}{\eta T}+\frac{\eta}{b}\mathrm{Tr}(F^{\star})\right)\log(4/\delta),$
+for a constant $C$ depending only on $M$, $c_{0}$, and $c_{1}$.
 Step 6 (Fisher dual norm).
@@
 $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;\leq\;C\!\left(\frac{\Delta L}{\eta T\lambda_{\min}(F^{\star})}+\frac{\eta}{b}\,\kappa_{F}\,d_{\operatorname{eff}}(F^{\star})\right)\!\log\!\frac{1}{\delta},$
-establishing (45). The oracle complexity (46) follows by balancing the two terms and setting $N=Tb$. ∎
+establishing (48).
+For (49), balance the two terms by taking $\eta^{\star}=\sqrt{\Delta L\,b/\bigl(T\,\lambda_{\min}(F^{\star})\,\kappa_{F}\,d_{\operatorname{eff}}\bigr)}$, which gives the balanced RHS $\sim\sqrt{\Delta L\,\kappa_{F}\,d_{\operatorname{eff}}/(Tb\,\lambda_{\min})}\,\log(1/\delta)$, i.e., the squared-norm gap decays as $N^{-1/2}$ in $N=Tb$.
+Requiring this squared gap to be at most $\varepsilon^{2}$ yields the $\varepsilon^{-4}$ scaling of (49).
+∎
 Remark 7.4 (Statistical vs. Euclidean conditioning).
@@
 Fix a total sampling budget $N$ and run $T=N/b$ iterations of constant-stepsize SGD with batch size $b$.
-Equating the bias term $\sim(L(\theta_{0})-L^{\star})/(\eta T\lambda_{\min}(F^{\star}))$ and the variance term $\sim\eta\kappa_{F}d_{\operatorname{eff}}/b$ in (45) yields the balanced stepsize $\eta^{\star}\propto\sqrt{b/(T\kappa_{F}d_{\operatorname{eff}})}$.
-Substituting $T=N/b$, the resulting Fisher-dual stationarity gap satisfies
+Balancing the bias term $\sim(L(\theta_{0})-L^{\star})/(\eta T\lambda_{\min}(F^{\star}))$ and the variance term $\sim\eta\kappa_{F}d_{\operatorname{eff}}/b$ in (48) yields the balanced stepsize $\eta^{\star}\propto\sqrt{b/(T\kappa_{F}d_{\operatorname{eff}})}$.
+Feasibility.
+The balanced stepsize must respect the theorem’s own step-size cap $\eta\leq 1/(4M)$.
+Substituting $T=N/b$ gives $\eta^{\star}=b\sqrt{\Delta L/(N\,\lambda_{\min}(F^{\star})\,\kappa_{F}\,d_{\operatorname{eff}})}$, so the cap $\eta^{\star}\leq 1/(4M)$ imposes the feasibility condition
+$b\;\leq\;\sqrt{\frac{N\,\lambda_{\min}(F^{\star})\,\kappa_{F}\,d_{\operatorname{eff}}}{16\,M^{2}\,\Delta L}}.$
+The batch-invariance conclusion below is a statement within this feasibility region (so $b$ may grow like $\sqrt{N}$, not linearly in $N$), not at arbitrary $b$.
+Substituting $T=N/b$ and taking $\eta=\eta^{\star}$ within the feasibility region, the resulting squared Fisher-dual stationarity gap satisfies
 $\min_{t<T}\|\nabla L(\theta_{t})\|_{(F^{\star})^{-1}}^{2}\;=\;\mathcal{O}\!\left(\sqrt{\frac{\kappa_{F}\,d_{\operatorname{eff}}}{N}}\right),$
-which is independent of $b$ at leading order.
-The practical implication for OR settings is that, under this budget-constrained regime, small batches lose nothing in worst-case guarantee while gaining more frequent iterates; the Fisher/Godambe geometry determines the constant but not the scaling in $N$.
+so the norm gap decays as $N^{-1/4}$.
+Within the feasibility region the bound is independent of $b$ at leading order, which suggests—under a fixed oracle budget—that small batches do not degrade the worst-case guarantee while gaining more frequent iterates; the identified geometry determines the constant but not the scaling in $N$.
 8 Numerical Validation
-We validate the theory through six diagnostic experiments summarized in Table 3.
-The experiments probe downstream consequences of the identified covariance $\mathrm{Cov}(g_{B})\approx b^{-1}G^{\star}$ (Theorem 4.3): if the identification is correct, the Lyapunov equilibrium, convergence rates, and directional variance structure must all agree with the predictions derived from $G^{\star}$.
-Appendix B additionally validates the covariance identity itself under misspecification and approximate exchangeability.
-Table 3: Reproducibility capsule: key experimental parameters.
-Experiment $d$ Replicates $\eta$ Steps Sweep variable
-1 (Batch scaling) 2 500 0.05 200 000 $b\in\{1,2,4,\ldots,512\}$
-2 (Angle sweep) 2 500 0.05 200 000 $\theta\in[0,\pi/2]$
-3 (Rotation diagnostic) 2 $\phi\in[0,\pi/2]$ (theory)
-4 (OU plateau, d=10d\!=\!10) 10 2 000 $0.5/\lambda_{\max}$ 300–10 000 $T$ (horizon)
-5 ($1/N$ rate) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 $N=Tb$
-6 (Directional var.) 10 800 $\eta_{0}/(t\!+\!t_{0})$ 600–9 600 eigenvectors of $G^{\star}$
-Experiments 1–3: Euler–Maruyama, $d=2$. Experiments 4–6: direct SGD recursion, $d=10$, $b=64$.
-8.1 Claim 1: The Lyapunov Equilibrium Law Is Quantitatively Predictive
-The first claim is that the Fisher–Lyapunov balance (Corollary 5.7) is not merely a qualitative insight but a quantitatively accurate prediction of steady-state risk, and that the shape of the noise covariance—not just its scale—governs which directions receive noise energy at equilibrium.
-Three experiments validate this at increasing dimensionality.
-Throughout, when the simulator uses finite-step (Euler–Maruyama or direct SGD) dynamics, the mathematically correct benchmark is the discrete Lyapunov equation (Proposition 5.9); the continuous-time OU prediction is its small-$\eta$ asymptotic reduction.
-The regime of approximation accuracy is governed by the product $\eta\lambda_{\max}(H^{\star})$: when this quantity is small (say $\leq 0.1$), the continuous and discrete benchmarks are nearly indistinguishable; for moderate stepsizes the discrete equation is the appropriate comparator.
-8.1.1 Experiment 1: Fisher–Lyapunov Equilibrium and $1/b$ Scaling (OU Benchmark).
-We begin with the local OU surrogate implied by the diffusion limit near a nondegenerate optimum:
+The empirical program is organized in three layers, each answering a distinct question about the identified covariance and its downstream consequences:
+  1. (i)
+    Direct identification (Experiment 1).
+Does $\mathrm{Cov}(g_{B}(\theta^{\star}))\approx b^{-1}G^{\star}(\theta^{\star})$ hold in an actual mini-batching setup—including under model misspecification, where curvature and noise geometry diverge?
+  2. (ii)
+    Fluctuation-scale geometric illustrations (Experiment 2).
+What does the Lyapunov structure implied by the identified covariance look like in a low-dimensional setting where individual eigendirections can be visualized?
+  3. (iii)
+    Raw-scale quantitative validations (Experiments 3–6).
+Do the original-iterate-scale predictions—discrete Lyapunov plateau at constant step, $1/N$ rate under Robbins–Monro schedules, directional amplification, and the fluctuation-scale $(\eta/b)$-collapse (signature of Theorem 5.7)—hold quantitatively in direct SGD on higher-dimensional problems?
+Layer (i) is the direct check of Theorem 4.3.
+Layer (ii) is interpretive: it visualizes the fluctuation-scale Lyapunov geometry on the normalized time-changed OU process of Theorem 5.7, and is not a direct Monte Carlo verification of raw SGD.
+Layer (iii) carries the main empirical weight: Experiments 3–6 are direct SGD recursions on the original iterate scale, and are the quantitative counterparts of the Lyapunov, rate, directional, and scaling-collapse predictions derived in Sections 57.
+Additional robustness diagnostics (approximate exchangeability) are deferred to Appendix B; reproducibility details (seeds, shared parameters) are in the code archive.
+8.1 Claim 1: The Mini-Batch Covariance Is Identified by the Sampling Design
+Before probing downstream consequences, we verify the covariance identification of Theorem 4.3 directly.
+8.1.1 Experiment 1: Direct Covariance Identification (Probit DGP, Logistic Fit).
+We generate n=50,000n=50{,}000 observations from a probit data-generating process and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$.
+Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}(\nabla_{\beta}\ell(\beta;X,Y)|_{\beta^{\star}})\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]|_{\beta^{\star}}$.
+The covariance-identification theorem nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx b^{-1}J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation.
+We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates).
+Figure 2 shows close agreement (relative Frobenius difference $\approx 0.05$): the identification holds even under misspecification, and mini-batch fluctuations are governed by the score covariance rather than by curvature alone.
+Figure 2: Experiment 1: direct covariance identification under misspecification (probit DGP, logistic fit).
+Diagonal entries of the empirical mini-batch covariance vs.
+$b^{-1}\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.
+8.2 Claim 2: The Lyapunov Geometry Is Visible in Low Dimensions
+Once the identification is empirically in hand (Experiment 1), it is useful to visualize what the Lyapunov structure looks like in a two-dimensional setting before moving to the $d=10$ quantitative tests.
+In the fluctuation-scale OU surrogate of Theorem 5.7, the stationary covariance $\Sigma_{U}$ solves $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$, and Experiment 2 illustrates three consequences that are each most naturally displayed in $d=2$: the $1/b$ scaling of marginal variance, the redistribution of variance under curvature–noise misalignment, and the failure of trace-matched isotropic surrogates to reproduce directional structure.
+These are interpretive illustrations, not direct quantitative validation of SGD: the simulator integrates the time-changed OU process of (54) rather than running SGD iterates, and these subplots are included because a 2D visualization of the Lyapunov geometry aids intuition for the higher-dimensional quantitative tests that follow.
+The quantitative burden is carried by Experiments 3–6, which run direct SGD recursions.
+8.2.1 Experiment 2: 2D Normalized OU Visualization of Lyapunov Structure.
+For visualization we simulate a pedagogical normalized OU process chosen so that marginal variance scales as $1/b$ and is independent of $\eta$ at leading order:
 $d\theta_{t}\;=\;-\eta\,H\,\theta_{t}\,dt\;+\;\sqrt{\frac{2\eta}{b}}\;F^{1/2}\,dW_{t},$ (54)
-with a two-dimensional sharp/flat curvature $H=\mathrm{diag}(1,0.1)$ and isotropic noise $F=I_{2}$, sweeping $b$ over a wide range.
-At stationarity, the covariance $\Sigma(b)$ solves the Lyapunov equation
-$H\Sigma+\Sigma H\;=\;\frac{2}{b}\,F,$ (48)
-so every stationary marginal variance scales as $1/b$; $\eta$ governs mixing time but not the fixed point.
-For the Euler–Maruyama simulation used here, the discrete Lyapunov equation (49) is the exact benchmark; the continuous equation (48) coincides at leading order because the stepsize satisfies $\eta\lambda_{\max}(H)\leq 0.05$.
-Figure 2 overlays simulated stationary variances with the Lyapunov prediction and shows clean $1/b$ scaling across the batch sweep.
-The flat coordinate mixes more slowly (as expected under weak curvature) but converges to the same equilibrium law.
-This serves as a baseline check: the Fisher–Lyapunov equilibrium is quantitatively predictive, not merely qualitative.
-We observed early entry into the basin of Fisher geometry even from random starts; see Section 9.3 for a diagnostic based on the stability ratios $\delta_{F}$ and $\delta_{H}$.
-Figure 2: OU benchmark: stationary variances vs. batch size $b$ (log scales). Solid: simulation; dashed: Lyapunov prediction from (48).
+with $H=\mathrm{diag}(1,0.1)$ and varying noise shape $F$.
+This normalization is chosen for visualization only; it is not the literal fluctuation SDE of Theorem 5.7, which scales out the $b$-dependence into the $\sqrt{b/\eta}$ rescaled process $U^{\eta}$.
+At stationarity, the continuous-time Lyapunov equation for (54) reads $H\Sigma+\Sigma H=(2/b)F$, and the resulting marginal variances visualize the $1/b$ scaling, curvature–noise misalignment effects, and rotation diagnostics of the Lyapunov geometry in two dimensions.
+Batch-size sweep ($1/b$ scaling).
+With $F=I_{2}$ and a wide sweep of $b$, Figure 3 overlays simulated stationary variances with the fluctuation-scale Lyapunov prediction: clean $1/b$ scaling holds across the sweep, with the flat coordinate mixing more slowly (weaker curvature) but reaching the same equilibrium law.
+Figure 3: Experiment 2, 2D batch-size sweep: stationary variances vs. $b$ (log scales); solid = simulation, dashed = fluctuation-scale Lyapunov prediction.
 $d=2$, $H=\mathrm{diag}(1,0.1)$, $F=I_{2}$, 500 replicates, 200k steps.
-8.1.2 Experiment 2: Anisotropic Fisher–Hessian Misalignment (Angle Sweep).
-We keep the same curvature $H=\mathrm{diag}(1,0.1)$ but rotate an anisotropic noise-shape matrix relative to the curvature basis:
-$F(\varphi)=R(\varphi)\begin{pmatrix}1.5&0\\[2.0pt] 0&0.5\end{pmatrix}R(\varphi)^{\top},\qquad\varphi\in[0,\pi/3],$
-with fixed batch size, isolating alignment as an operational parameter.
-Because the simulator is Euler–Maruyama with finite time step, the appropriate comparator is the discrete Lyapunov equation: writing $A=I-\eta H\Delta t$ and $Q=(2\eta/b)F(\theta)\Delta t$, the stationary covariance solves
-$\Sigma\;=\;A\Sigma A^{\top}+Q.$ (49)
-Figure 3 shows a clear transfer of stationary variance across coordinates as $\varphi$ varies: when the noise eigenbasis rotates away from the curvature eigenbasis, equilibrium variance is redistributed in a predictable way, and the discrete Lyapunov benchmark tracks the simulation closely.
-The takeaway is structural: alignment governs which directions receive noise energy at equilibrium, even when scalar temperature $\eta/b$ is held fixed.
-Figure 3: Angle sweep: stationary variances vs. misalignment angle. Bands show replicate variability; dashed curves are the discrete Lyapunov prediction from (49).
+Angle sweep (curvature–noise misalignment).
+Rotating an anisotropic noise matrix $F(\varphi)=R(\varphi)\mathrm{diag}(1.5,0.5)R(\varphi)^{\top}$ relative to the curvature basis, Figure 4 shows a predictable transfer of stationary variance across coordinates as $\varphi$ varies: alignment governs which directions receive noise energy at equilibrium, even when the scalar temperature $\eta/b$ is held fixed.
+The discrete Lyapunov benchmark tracks the simulation closely.
+Figure 4: Experiment 2, 2D angle sweep: stationary variances vs. misalignment angle $\varphi$. Bands show replicate variability; dashed curves are the discrete Lyapunov prediction.
 $d=2$, $H=\mathrm{diag}(1,0.1)$, 500 replicates per angle.
-8.1.3 Experiment 4: OU/Lyapunov Plateau in $d=10$ (Anisotropic vs. Isotropic).
-Experiments 1–2 used $d=2$.
-We now move to $d=10$ to validate the Lyapunov equilibrium in a setting where anisotropy is richer.
-The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100; the noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched).
+Rotation diagnostic (geometry vs. scalar temperature).
+We compare a geometric OU with rotating diffusion $F_{\star}(\varphi)$ against a trace-matched isotropic surrogate $\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}F_{\star}/d$.
+Figures 56 show that the geometric stationary covariance $\Sigma_{F}(\varphi)$ tilts with $\varphi$ (nonzero cross-covariance, redistributed marginals), while the isotropic surrogate remains axis-aligned at all angles.
+This is the signature the trace-matched surrogate cannot reproduce, and is the low-dimensional preview of the directional amplification measured in $d=10$ by Experiment 5.
+Figure 5: Experiment 2, 2D rotation diagnostic: stationary covariance ellipses at three rotation angles.
+The geometric equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
+$d=2$, $b=64$; ellipses are 95% level sets of the Lyapunov solution.
+Figure 6: Experiment 2, 2D rotation diagnostic: cross-covariance $\Sigma_{12}$ vs.
+rotation angle.
+Isotropic trace-matched noise fails to reproduce off-diagonal covariance; bands show $\pm 1$ s.d.
+across seeds.
+$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in =800=800.
+8.3 Claim 3: The Lyapunov Plateau Holds Quantitatively in $d=10$
+The 2D experiments display the qualitative geometry; we now move to $d=10$ for the first high-dimensional quantitative checkpoint.
+This is where the quantitative weight of the Lyapunov prediction is tested: Experiment 3 runs direct SGD on a $d=10$ problem and asks whether the empirically observed steady-state Fisher risk matches the level predicted by combining the identified covariance $G^{\star}$ with the discrete Lyapunov equation (Proposition 5.13).
+If the identification from Theorem 4.3 is quantitatively correct, both the anisotropic and isotropic empirical risks should converge to the Lyapunov-predicted plateau $\mathrm{Tr}(F^{\star}\Sigma_{\eta}^{\mathrm{lin}})$ as $T$ grows, with convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$.
+8.3.1 Experiment 3: OU/Lyapunov Plateau in $d=10$ (Anisotropic vs. Isotropic).
+The curvature is $H=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{10})$ with eigenvalues log-spaced between 1 and 100, and we set $F^{\star}=H^{\star}$ throughout the synthetic $d=10$ benchmark (the quadratic-Gaussian surrogate in which curvature and Fisher information coincide, so the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ directly measures Hessian-weighted error).
+The noise shape $G^{\star}$ is a random SPD matrix with eigenvalues in $[0.5,5]$, and we compare an isotropic control $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched).
 Under constant step size $\eta=0.5/\lambda_{\max}(H)$ and batch size $b=64$, we run R=2, [^0] R=2{,}000 replicates and record the Fisher-metric risk $\mathbb{E}[\|\theta_{T}\|_{F^{\star}}^{2}]$ over increasing horizons $T\in\{300,800,2000,5000,10000\}$.
-Table 4 confirms the Lyapunov prediction (Corollary 5.7): both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}/R_{\infty}$ approaching unity from both sides.
-The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; directional differences emerge only in higher-order structure (Claim 3 below).
+Table 3 confirms the Lyapunov prediction: both the anisotropic and isotropic empirical risks converge to the predicted stationary level as $T$ grows, with convergence ratios $\widehat{R}_{T}/R_{\infty}$ within $1\%$ of unity across all horizons.
+The two plateaux nearly coincide because trace-matching is exact in this diagonal-$H$ setting; the distinction between anisotropic and isotropic geometry is therefore invisible at the level of scalar risk, and only emerges in higher-order structure (Experiment 5).
 Empirical Risk $\widehat{R}_{T}$ Lyapunov $R_{\infty}$ $\widehat{R}_{T}/R_{\infty}$
@@
 10 000 $8.63\!\times\!10^{-4}$ $8.47\!\times\!10^{-4}$ 1.000 0.987
-Table 4: OU/Lyapunov plateau ($d=10$): empirical Fisher-metric risk vs.
-discrete Lyapunov prediction $R_{\infty}=\mathrm{Tr}(F^{\star}\Sigma_{\infty})$.
+Table 3: Experiment 3 (OU/Lyapunov plateau, $d=10$): empirical Fisher-metric risk vs.
+discrete Lyapunov prediction $R_{\infty}=\mathrm{Tr}(F^{\star}\Sigma_{\eta}^{\mathrm{lin}})$.
 Convergence ratios $\widehat{R}_{T}/R_{\infty}\to 1$ confirm quantitative accuracy.
@@
 8.4 Claim 4: Decaying Step Sizes Recover a $1/N$ Rate with a Geometry-Determined Constant
-The second claim is that under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$.
+Under a Robbins–Monro step-size schedule the Fisher-metric risk decays as $C/N$ where $N=Tb$ is the total oracle budget and the constant $C$ is determined by the intrinsic geometry $(G^{\star},H^{\star},F^{\star})$.
 Specifically, the SA central-limit theorem gives an asymptotic covariance $V$ satisfying $({\eta_{0}H^{\star}-\tfrac{1}{2}I})V+V({\eta_{0}H^{\star}-\tfrac{1}{2}I})=\frac{\eta_{0}^{2}}{b}G^{\star}$, and $N\cdot\mathrm{Risk}\to b\,\mathrm{Tr}(F^{\star}V)$.
-This validates the matching upper and lower bounds (Theorem 6.5 and Proposition 6.10).
-8.2.1 Experiment 5: $1/N$ Rate and Fisher Constant (Decaying Step Size).
-We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 4, with oracle budget $N=Tb$ and $b=64$.
-Figure 4 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), confirming Theorem 6.5.
+This is consistent with the upper bound (Theorem 6.5) in rate order and with the i.i.d.
+parametric Fisher lower-bound benchmark (Proposition 6.8).
+8.4.1 Experiment 4: $1/N$ Rate and Fisher Constant (Decaying Step Size).
+We switch to a decaying step size $\eta_{t}=\eta_{0}/(t+t_{0})$ with $\eta_{0}=1.2$, $t_{0}=50$, and the same $(H,G^{\star},G_{\mathrm{iso}})$ from Experiment 3, with oracle budget $N=Tb$ and $b=64$.
+Figure 7 (left) plots the Fisher-metric risk on log–log axes: both anisotropic and isotropic risks decay as $1/N$ (fitted slopes $\approx-0.98$), consistent with Theorem 6.5 in rate order.
 The right panel shows the scaled risk $N\cdot\widehat{\mathrm{Risk}}$, which stabilizes near $\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic)—both close to the SA–CLT prediction $b\,\mathrm{Tr}(F^{\star}V)$ where $V$ solves the Lyapunov equation above.
-The inset marks the auxiliary constant $\mathrm{Tr}(G^{\star}H^{-1})\approx 5.41$, which governs the constant-step OU stationary risk (Corollary 5.7); the ratio of the decaying-step plateau to this constant reflects the step-size schedule and the interplay between curvature and noise geometry.
-Figure 4: Decaying step size, $d=10$.
+Figure 7: Experiment 4 (decaying step size, $d=10$).
 Left: Fisher-metric risk vs.
-NN (log–log); slopes $\approx-1$ confirm $1/N$ rate.
+NN (log–log); slopes $\approx-1$.
 Right: Scaled risk $N\cdot\widehat{\mathrm{Risk}}$ stabilizes; inset shows $C=\mathrm{Tr}(G^{\star}H^{-1})$.
 $\eta_{t}=1.2/(t+50)$, $b=64$, 800 replicates.
-Table 5 summarizes the tightness of the bounds.
-The empirical risk is bounded between the information-theoretic lower bound (Proposition 6.10) and the upper bound (Theorem 6.5), with $N\cdot\text{Risk}$ stabilizing near $14.2$ (anisotropic) and $14.0$ (isotropic), both above the van Trees lower bound cd=10c\,d=10.
-At the largest budget (N=614,400N=614{,}400) the ratio of empirical constant to lower bound is approximately $1.4\times$, indicating near-tightness.
-The rate and constant are thus determined by $(G^{\star},H^{\star})$, so the oracle budget for a target risk can be computed from the problem’s intrinsic geometry.
-Table 5: Upper vs. lower bound: scaled risk $N\cdot\widehat{\mathrm{Risk}}$ compared with the van Trees lower bound $cd=10$.
-NN $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{aniso}}$ $N\!\cdot\!\widehat{\mathrm{Risk}}_{\mathrm{iso}}$ $d=10$ (lower bound)
-38 400 13.59 13.41 10
-76 800 13.66 13.44 10
-153 600 13.66 14.26 10
-307 200 14.07 14.25 10
-614 400 14.35 14.42 10
-The SA–CLT Lyapunov prediction gives $b\,\mathrm{Tr}(F^{\star}V)\approx 14.2$ (anisotropic) and $\approx 14.0$ (isotropic); $N\cdot\mathrm{Risk}$ stabilizes above the van Trees lower bound cd=10c\,d=10 as expected.
-8.3 Claim 3: Scalar Temperature Matching Fails—Directionality Is the Signature of Geometry
-The third claim is that matching the total noise power (scalar temperature $\tau=\eta/b$) is insufficient: the directional distribution of noise energy is a distinguishing signature of the identified geometric model that no isotropic surrogate can reproduce.
-The Lyapunov trace identity $\mathrm{Tr}(H^{\star}\Sigma_{\infty})=\frac{\tau}{2}\mathrm{Tr}(G^{\star})$ depends only on $\mathrm{Tr}(G^{\star})$, so a trace-matched isotropic surrogate reproduces total risk by construction; but the directional distribution is entirely wrong, spreading residual error uniformly rather than concentrating it along top noise eigendirections.
-8.3.1 Experiment 3: Geometry vs. Scalar Temperature (Rotation Diagnostic).
-We compare a geometric OU model (diffusion covariance $\propto F_{\star}(\phi)$, rotating with angle $\phi$) against a trace-matched isotropic surrogate ($\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(F_{\star})/d$).
-At stationarity, the geometric covariance $\Sigma_{F}(\phi)$ generically rotates (producing nontrivial cross-covariance), while the isotropic $\Sigma_{I}(\phi)$ remains axis-aligned.
-Figures 57 confirm this: the ellipses tilt, cross-covariance $\Sigma_{12}$ tracks theory, and marginal variances redistribute—effects that temperature matching alone cannot capture.
-Figure 5: Stationary covariance ellipses at three rotation angles: the identified equilibrium tilts with $\varphi$; the isotropic trace-matched surrogate remains axis-aligned.
-$d=2$, $b=64$. Ellipses: 95% level sets of the Lyapunov solution.
-Figure 6: Cross-covariance $\Sigma_{12}$ vs. rotation angle $\varphi$: isotropic trace-matched noise fails to reproduce off-diagonal covariance. Bands show $\pm 1$ s.d. across seeds.
-$d=2$, 6 seeds, T=2,500T=2{,}500, burn-in=800{}=800.
-Figure 7: Marginal variances $\Sigma_{11}$ (left) and $\Sigma_{22}$ (right) vs. angle: total noise energy may match, but directional allocation differs.
-8.3.2 Experiment 6: Directional Variance Amplification.
-We now move to $d=10$ and ask whether anisotropy affects how risk is distributed across directions.
-For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$.
-Figure 8 (left) displays these on log–log axes: both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections.
-The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$; at the largest budget the ratios are $1.65$, $1.45$, and $1.26$ for $k=1,2,3$, mirroring the eigenvalue ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}$ of $G^{\star}$.
-These amplification ratios show that the identified geometry determines which parameter directions carry the most residual risk—information that scalar temperature models discard.
-Figure 8: Directional variance amplification ($d=10$).
+Interpretation: two related-but-distinct constants.
+The constant-step Lyapunov quantity $\mathrm{Tr}(G^{\star}H^{\star-1})$ (inset, $\approx 5.41$) and the decaying-step asymptotic constant $b\,\mathrm{Tr}(F^{\star}V)$ (plateau, $\approx 14.0$$14.2$) are related but not identical: the former is the stationary $\mathrm{Tr}(F^{\star}\Sigma_{\eta}^{\mathrm{lin}})$ of the constant-step discrete Lyapunov recursion at leading order in $\eta$, while the latter is the $N\to\infty$ asymptotic constant of the scaled risk under a Robbins–Monro schedule, which solves a different Lyapunov equation shifted by $\tfrac{1}{2}I$.
+The shift reflects the variance deflation introduced by the decaying schedule.
+Both constants are determined by the identified geometry $(G^{\star},H^{\star},F^{\star})$, but their numerical values differ, and the ratio between them depends on the schedule (here $\eta_{0}=1.2$, $t_{0}=50$).
+The empirical scaled risk $N\cdot\widehat{\mathrm{Risk}}$ lies above the i.i.d.
+parametric Fisher lower-bound benchmark of Proposition 6.8, consistent with rate-order matching; we do not claim a specific value for the theorem constant.
+8.5 Claim 5: Directional Amplification Discriminates Geometry from Scalar Temperature
+Experiments 3 and 4 compared anisotropic and isotropic noise at the level of trace-scale summaries, where trace-matching makes the two surrogates nearly indistinguishable.
+This is precisely where scalar-temperature reasoning fails.
+In the synthetic benchmark $F^{\star}=H^{\star}$, so the plotted Fisher-metric risk is $\mathrm{Tr}(F^{\star}\Sigma)=\mathrm{Tr}(H^{\star}\Sigma)$.
+Multiplying the fluctuation-scale Lyapunov equation $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$ on the left by the identity and taking traces gives $2\,\mathrm{Tr}(H^{\star}\Sigma_{U})=\mathrm{Tr}(G^{\star})$, which depends on $G^{\star}$ only through its trace; a trace-matched isotropic surrogate ($G_{\mathrm{iso}}$ with $\mathrm{Tr}G_{\mathrm{iso}}=\mathrm{Tr}G^{\star}$) therefore reproduces the scalar Fisher-risk summary by construction, even though its directional distribution of residual risk is entirely different.
+The discriminating signature is directional: anisotropic noise concentrates stationary risk along the top eigendirections of $G^{\star}$, while the isotropic surrogate spreads residual error uniformly.
+Experiment 5 measures this direction-by-direction.
+8.5.1 Experiment 5: Directional Variance Amplification.
+For $k=1,2,3$ we project $\theta_{T}$ onto the top eigenvectors $v_{k}$ of $G^{\star}$ and compute the directional second moment $\mathbb{E}[\langle v_{k},\theta_{T}\rangle^{2}]$ under the same decaying-step setup as Experiment 4.
+Figure 8 (left) shows that both models decay as $1/N$, but the anisotropic curve is systematically above the isotropic one along the top eigendirections.
+The right panel plots the ratio $\mathbb{E}_{\mathrm{aniso}}/\mathbb{E}_{\mathrm{iso}}$: at the largest budget the empirical ratios are $(1.76,1.33,1.35)$ for $k=1,2,3$, close to the linearized Lyapunov prediction $(1.77,1.40,1.33)$ obtained from $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}$.
+The ratios are strictly greater than one in every direction and largest along the top eigenvector of $G^{\star}$.
+Trace-matched isotropic noise matches the scalar-risk summaries of Experiments 3–4 but cannot reproduce this direction-by-direction risk allocation; the identified geometry determines which parameter directions carry the most residual risk.
+Figure 8: Experiment 5 (directional variance amplification, $d=10$).
 Left: Directional second moments along top eigenvectors of $G^{\star}$.
 Right: Ratio of anisotropic to isotropic directional variances; ratios >1>1 mirror the eigenvalue ordering of $G^{\star}$.
-Temperature matching preserves $\mathrm{Tr}(G^{\star})$ but not the eigen-structure; anisotropic noise concentrates residual risk along the top noise eigendirections.
-Same as Experiment 5; top $k=1,2,3$ eigenvectors of $G^{\star}$.
-Additional robustness experiments under misspecification and approximate exchangeability are reported in Appendix B.
+Trace-matched isotropic noise preserves $\mathrm{Tr}(G^{\star})$ but spreads residual risk uniformly; the identified anisotropic noise concentrates it along the top noise eigendirections.
+Same setup as Experiment 4; top $k=1,2,3$ eigenvectors of $G^{\star}$.
+8.6 Claim 6: Fluctuation-Scale $(\eta/b)$-Collapse on the Raw Iterate Path
+Theorem 5.7 and Corollary 5.11 predict that, in the small-stepsize local regime, $\mathbb{E}[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}]=(\eta/b)\,\mathrm{Tr}(F^{\star}\Sigma_{U})+o(\eta/b)$ with $\Sigma_{U}$ solving $H^{\star}\Sigma_{U}+\Sigma_{U}H^{\star\top}=G^{\star}(\theta^{\star})$.
+Equivalently, the scaled Fisher risk $(b/\eta)\,\widehat{\mathrm{Risk}}$ should collapse to a single $\eta,b$-free curve at leading order.
+This is the direct empirical signature of the fluctuation FCLT (Theorem 5.7) on the original iterate scale.
+8.6.1 Experiment 6: Raw-Scale Collapse Diagnostic.
+We sweep eight $(\eta,b)$ configurations with $\eta\in\{0.001,0.002,0.005\}$, $b\in\{32,64,128\}$ (so $\eta\lambda_{\max}(H^{\star})\in\{0.1,0.2,0.5\}$; $\eta/b\in[8\times 10^{-6},1.6\times 10^{-4}]$), and run direct SGD on the same $(H,F^{\star},G^{\star})$ as Experiments 3–4 for $T\in\{2000,5000,10000,20000\}$ with $R=600$ replicates per cell.
+Figure 9 plots $(b/\eta)\,\widehat{\mathbb{E}}[\|\theta_{T}\|_{F^{\star}}^{2}]$ against $T$ for each configuration.
+All curves collapse onto a common plateau near $\mathrm{Tr}(F^{\star}\Sigma_{U})\approx 10.22$ (black reference line), the fluctuation-scale Lyapunov prediction.
+The collapse across two orders of magnitude in $\eta/b$ is a direct verification of the $(\eta/b)$-scaling law on the raw iterate scale, independently of any continuous-time normalization convention.
+Figure 9: Experiment 6 (fluctuation-scale collapse diagnostic; Theorem 5.7): scaled Fisher risk $(b/\eta)\,\widehat{\mathbb{E}}[\|\theta_{T}\|_{F^{\star}}^{2}]$ stabilizes at the fluctuation-scale Lyapunov prediction $\mathrm{Tr}(F^{\star}\Sigma_{U})$ (black line) across eight $(\eta,b)$ configurations.
+$d=10$, direct SGD, $R=600$ replicates per cell; $\eta\in\{0.001,0.002,0.005\}$, $b\in\{32,64,128\}$.
+Additional robustness experiments under approximate exchangeability are reported in Appendix B.
 9 Discussion
 9.1 Summary of Contributions
-The paper’s central contribution is the identification of the mini-batch gradient covariance from sampling design and exchangeability as the projected gradient-noise matrix $G^{\star}(\theta)$—which reduces to Fisher information at $\theta^{\star}$ under correct specification and, for general losses, induces Godambe/sandwich geometry together with $H^{\star}$.
-This identification endogenizes the diffusion coefficient, determines the Lyapunov stationary covariance, and selects the natural metric for rates and complexity, yielding matching $1/N$ bounds governed by intrinsic dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$.
-  • Proven: Alignment (Theorem 4.3), diffusion (Theorem 5.4), OU/Lyapunov (Corollary 5.7), discrete bridge (Proposition 5.9), matching bounds (Theorem 6.5 and Proposition 6.10), and oracle complexity (Theorem 7.3).
-  • Sketched: Batch-size scaling (Corollary 7.8) and variance-reduction bridge (Remark 9.1).
-  • Deferred: Misspecification, weak dependence, explicit batching policies, and OR-specific applications (Appendix B provides preliminary experiments).
-9.2 Implications for Algorithm Design
-In the diffusion regime the key knob is the effective temperature $\tau=\eta/b$.
-Changing $b$ does more than “reduce variance”; it changes the diffusion amplitude in a directional way determined by the identified geometry, and therefore changes the Lyapunov-predicted error floor.
-Because the intrinsic noise is matrix-valued, variance-reduction interventions should be evaluated by their effect on risk in a chosen metric (e.g., $\mathrm{Tr}(Q\Sigma)$) rather than Euclidean variance alone, connecting classical OR variance-reduction practice to SGD through the correct geometry.
+The paper’s central contribution is operational: we identify the mini-batch gradient covariance from the sampling design itself as $b^{-1}G^{\star}(\theta)$, converting the diffusion analysis of constant-step SGD from a modeling exercise into a statement about the replication budget.
+In likelihood models the identified matrix reduces to projected Fisher information; in general losses it pairs with $H^{\star}$ through the classical sandwich/Godambe geometry of M-estimation.
+At fixed batch size, the raw iterate path has a deterministic fluid limit (the gradient-flow ODE); the $\sqrt{b/\eta}$-scaled fluctuations converge to a linear diffusion with coefficient $G^{\star}$; near a nondegenerate optimum this is Ornstein–Uhlenbeck, and the corresponding Lyapunov covariance scaled by $\eta/b$ matches the stationary covariance of the linearized discrete recursion at leading order.
+In the local Fisher case, an upper bound of order $1/N$ is rate-order matched by an i.i.d.
+van Trees lower bound, with the intrinsic dimension $d_{\mathrm{eff}}$ and statistical condition number $\kappa_{F}$ appearing in the upper bound only.
+9.2 Operational Implications for Sampling-Budget Design
+The effective temperature $\tau=\eta/b$ is the single scalar that converts replication effort into diffusion amplitude, and the identification shows it does so in a directional way fixed by $G^{\star}$ rather than uniformly.
+For an OR practitioner, three consequences follow at the sampling-budget level.
+First, at a fixed oracle budget $N=Tb$, the batch size $b$ trades iterate count against per-step noise scale: increasing $b$ cools the error floor predicted by the linearized discrete Lyapunov benchmark (Proposition 5.13) in magnitude, but leaves the noise ellipsoid’s shape unchanged.
+Second, classical OR variance-reduction interventions (common variates, control variates, stratification) should be evaluated by their Loewner effect on $G^{\star}$, and thus by the induced reduction in $\mathrm{Tr}(Q\Sigma)$ for the operational metric $Q$, rather than by scalar component-variance alone—this is the natural bridge from classical variance-reduction practice to mini-batch SGD.
+Third, the directional amplification in Experiment 5 identifies exactly which parameter components absorb residual risk, and thus where additional sampling effort has the largest marginal return.
 Remark 9.1 (Variance reduction as Fisher-metric risk reduction).
 A VR technique that reduces $G^{\star}$ to $\widetilde{G}\preceq G^{\star}$ in the Loewner order yields $\widetilde{\Sigma}_{\infty}\preceq\Sigma_{\infty}$ via the Lyapunov balance, and hence lower Fisher-metric risk.
-SGD does not precondition the drift; natural-gradient methods do so by reweighting $\nabla L$.
-Our results explain why plain SGD nevertheless exhibits geometry-aware behavior through its noise structure: drift contraction and diffusion injection are distinct levers.
-The control view also suggests adaptive batching—increasing $b$ (cooling) when marginal variance reduction is valuable in the performance metric of interest, and decreasing it when exploration matters.
+Plain SGD does not precondition the drift; natural-gradient methods do so by explicitly reweighting $\nabla L$.
+We do not claim plain SGD is “natural-gradient-like” in the drift sense; rather, its linearized local equilibrium benchmark (Proposition 5.13) reflects the noise covariance $G^{\star}$ through the Lyapunov balance.
+The identification also suggests adaptive batching—increasing $b$ (cooling) when marginal noise reduction is valuable, decreasing it when exploration matters.
 9.3 Practical Notes and Extensions
@@
 9.4 Estimating $G^{\star}$ in Practice
-The preceding theory identifies $G^{\star}(\theta)$ as the central primitive.
-In applications where $G^{\star}$ is not available analytically, several estimation strategies provide operationally useful approximations.
-The simplest is the empirical covariance of per-sample gradients: given a pilot batch $\{z_{1},\ldots,z_{m}\}$ at $\theta_{t}$, the sample covariance $\hat{G}(\theta_{t}):=\frac{1}{m-1}\sum_{i=1}^{m}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ (where $\psi_{i}:=\nabla_{\theta}\ell(\theta_{t};z_{i})$) is the natural plug-in estimator and converges at the standard $O(1/\sqrt{m})$ rate in operator norm.
-For large $d$, a diagonal approximation $\hat{G}_{\mathrm{diag}}$ retains the per-coordinate noise scales and suffices for estimating $d_{\mathrm{eff}}$ and detecting gross anisotropy; when the leading eigendirections of $G^{\star}$ are the primary target, a rank-$k$ approximation via streaming PCA on per-sample gradients requires only $O(dk)$ storage.
-If $G^{\star}(\theta)$ varies slowly along the optimization path, an exponential moving average $\hat{G}_{t}=(1-\alpha)\hat{G}_{t-1}+\alpha\,(\psi_{t}-\bar{\psi}_{t})(\psi_{t}-\bar{\psi}_{t})^{\top}$ tracks the local geometry with minimal overhead.
-[…]
+Where $G^{\star}$ is not analytically available, the empirical covariance of per-sample gradients $\hat{G}(\theta_{t})=\frac{1}{m-1}\sum_{i}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ is the natural plug-in and converges at the standard $O(1/\sqrt{m})$ rate in operator norm; for large $d$, a diagonal or rank-$k$ streaming approximation suffices for the quantities ($d_{\mathrm{eff}}$, dominant eigendirections, scale anisotropy) that enter the bounds.
+Given $\hat{G}$, one can diagnose whether isotropic temperature-matching is likely to fail (by checking whether $\hat{G}$ is far from a scalar multiple of $I$) and compare candidate batch sizes through the predicted stationary covariance $\Sigma_{\eta}$ of Proposition 5.13.
+A natural design heuristic (not proved here) is to regulate $\tau_{t}\propto 1/\lambda_{\max}(H(\theta_{t}))$ to avoid over-heating in stiff directions; when $d_{\operatorname{eff}}\ll d$, the linearized local equilibrium of Proposition 5.13 concentrates on the identifiable subspace captured by $\Pi_{\theta}$, which is where the identified geometry is nondegenerate.
 9.5 Limitations and Future Work
@@
 Finally, preconditioners (Gauss–Newton, KFAC) reshape both drift and noise geometry; a joint control theory co-designing contraction and diffusion—potentially under communication constraints—would connect information geometry to OR-style variance allocation at scale.
-Appendix A Technical Lemmas and Concentration Tools
+The empirical design intentionally progresses from interpretable geometric diagnostics (direct covariance identification, 22D Lyapunov visualization) to quantitative raw-scale validation under direct SGD recursions in $d=10$, culminating in the fluctuation-scale $(\eta/b)$-collapse (Experiment 6; signature of Theorem 5.7) on the original iterate path.
+Code availability.
+All experiments in Section 8 and Appendix B are reproducible from the accompanying code archive (Python_Codes/), with master script run_all.py, explicit random seeds, and wall-clock under 5 minutes on a single CPU.
+Appendix A Technical Lemmas
 This appendix collects the auxiliary lemmas invoked by the main-text proofs.
-All spine results—the alignment theorem, the diffusion limit, the OU linearization, the Lyapunov corollary, the upper and lower bounds, and the oracle complexity guarantee—are proved in the body of the paper.
-A.1 Finite-population covariance identity
-The following identity, classical in survey sampling (Cochran, 1977; Serfling, 1980), is used in the proof of the alignment theorem (Theorem 4.3).
-Lemma A.1 (Without-replacement covariance of a mini-batch mean).
-Let $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ be deterministic vectors with mean $\bar{x}_{n}:=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ and centered sample covariance
-$S_{n}:=\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}.$
-Let $B$ be a subset of $\{1,\dots,n\}$ of size $b$ drawn uniformly without replacement, and define $\bar{x}_{B}:=\frac{1}{b}\sum_{i\in B}x_{i}$. Then, conditional on $(x_{1},\dots,x_{n})$,
-$\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}\left(1-\frac{b}{n}\right)S_{n}.$ (50)
-Proof.
-Write $\bar{x}_{B}=\frac{1}{b}\sum_{i=1}^{n}I_{i}x_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$.
-Under SRSWOR, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$ for $i\neq j$.
-Expanding $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=b^{-2}\sum_{i,j}\mathrm{Cov}(I_{i},I_{j})x_{i}x_{j}^{\top}$ and collecting terms:
-$b^{-2}\!\left[\frac{b(n-b)}{n^{2}}\sum_{i}x_{i}x_{i}^{\top}-\frac{b(n-b)}{n^{2}(n-1)}\sum_{i\neq j}x_{i}x_{j}^{\top}\right]$
-$\quad=\frac{n-b}{bn^{2}}\cdot\frac{n}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}=\frac{1}{b}\!\left(1-\frac{b}{n}\right)S_{n},$
-where the last equality uses $(n-1)S_{n}=\sum(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}$. ∎
-A.2 Score identities and information equality
-Lemma A.2 (Score identities and information equality).
-Assume $\theta\mapsto p_{\theta}(z)$ is differentiable, $\int p_{\theta}(z)\,dz=1$, and differentiation under the integral sign is valid. Then:
-  1. (i)
-    $\mathbb{E}_{\theta}[s_{\theta}(Z)]=0$, where $s_{\theta}(z):=\nabla_{\theta}\log p_{\theta}(z)$ is the score.
-  2. (ii)
-    $F(\theta)=\mathbb{E}_{\theta}[s_{\theta}(Z)s_{\theta}(Z)^{\top}]=-\mathbb{E}_{\theta}[\nabla_{\theta}^{2}\log p_{\theta}(Z)]$ (information equality).
-Proof.
-Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
-A.3 Lyapunov equation solution
+The finite-population covariance identity (Lemma 4.1) and score identities (Lemma 3.2) are stated and proved in the main text; we only restate here the auxiliary results (Lyapunov representation, Robbins–Monro product bound) that are invoked in later proofs but are not part of the main-text spine.
+A.1 Lyapunov equation solution
 Lemma A.1 (Lyapunov equation solution).
@@
 ∎
-A.4 Product bound for Robbins–Monro weights
+A.2 Product bound for Robbins–Monro weights
 Lemma A.2 (Product bound for Robbins-Monro weights).
@@
 ∎
-A.5 Freedman’s inequality
-Lemma A.5 (Freedman’s inequality).
-Let $(M_{t})_{t\geq 0}$ be a martingale with $M_{0}=0$ and increments $\Delta M_{t}:=M_{t}-M_{t-1}$ satisfying $|\Delta M_{t}|\leq R$ almost surely.
-Let $V_{T}:=\sum_{t=1}^{T}\mathbb{E}[(\Delta M_{t})^{2}\mid\mathcal{F}_{t-1}]$ be the predictable quadratic variation.
-Then for any $v>0$ and $\lambda>0$:
-$\Pr(M_{T}\geq\lambda\text{ and }V_{T}\leq v)\leq\exp\left(-\frac{\lambda^{2}}{2v+2R\lambda/3}\right).$
-Equivalently, with probability at least $1-\delta$, whenever $V_{T}\leq v$:
-$M_{T}\leq\sqrt{2v\log(1/\delta)}+\frac{2R}{3}\log(1/\delta).$
-Proof.
-See Freedman (1975) or Tropp (2015, Theorem 1.6). ∎
-Appendix B Robustness Beyond Pure Exchangeability
-The experiments in this appendix are exploratory robustness diagnostics, not part of the formal theoretical validation in Section 8.
-They demonstrate that the covariance-identification theory degrades gracefully under realistic departures from i.i.d.
-likelihood assumptions.
-Experiment A considers model misspecification, where curvature and noise geometry diverge; Experiment B models mild violations of exchangeability via variance inflation.
-B.1 Experiment A: Misspecification (Probit DGP, Logistic Fit)
-We generate n=50,000n=50{,}000 observations from a probit DGP and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$.
-Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}\!\bigl(\nabla_{\beta}\ell(\beta;X,Y)\big|_{\beta=\beta^{\star}}\bigr)\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]\big|_{\beta=\beta^{\star}}$.
-The covariance-identification theorem (Theorem 4.3) nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx(1/b)\,J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation.
-We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates).
-Figure 9 shows close agreement (relative Frobenius difference $\approx 0.05$), confirming that the identification remains accurate under misspecification: mini-batch fluctuations are governed by the score covariance $J^{\star}$ rather than by curvature alone.
-Figure 9: Experiment A: misspecification (probit DGP, logistic fit).
-Diagonal entries of the empirical mini-batch covariance vs.
-$(1/b)\,\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.
-B.2 Experiment B: Approximate Exchangeability and Variance Inflation
+Appendix B Robustness: A Stylized Variance-Inflation Perturbation
+The following experiment probes one stylized perturbation of the fresh-sampling assumption of Theorem 4.3: a scalar variance-inflation factor $(1+\varepsilon)$ on the mini-batch covariance, modeling mild departures from ideal exchangeability (weak dependence, partial sample reuse).
+This is a specific perturbation diagnostic, not a general robustness theorem.
+A direct check of the identification itself under model misspecification is in Experiment 1 of Section 8.1.1.
 To model mild departures from ideal exchangeability (e.g., weak dependence, contamination, or partial reuse of samples), we inflate the intrinsic covariance by a factor $(1+\varepsilon)$ in the linearized surrogate: $\mathrm{Cov}(\xi_{t})=(1/b)(1+\varepsilon)\,G^{*}(\theta^{\star})$, $\varepsilon\in[0,1]$.
@@
 Figure 10: Experiment B: approximate exchangeability. Stationary variances vs. $\varepsilon$ (simulation bands vs. linear Lyapunov prediction).
-Appendix C Reproducibility and Code Availability
-All experiments are reproducible from the accompanying code archive.
-The Python suite (Python_Codes/) includes a master script run_all.py that executes all experiments sequentially.
-Every script uses explicit random seeds (global seed =7=7) for deterministic reproducibility; small platform-level floating-point differences may produce minor variations ($\lesssim 1\%$) in Monte Carlo estimates.
-Software environment.
-Python $\geq$3.9; NumPy, SciPy, Matplotlib, Pandas. Tested under Python 3.10.12, NumPy 2.2, SciPy 1.15, Matplotlib 3.10. Total wall-clock time: $\approx$4 minutes on a single CPU core.
-Script inventory.
-Table 6 maps each script to the figures and tables it generates.
-Table 6: Code inventory: scripts, outputs, and paper cross-references.
-Script Experiment Figures / Tables Sec.
-partA_batch_scaling.py 1 (Batch scaling) Fig. 2 8.1
-partB_anisotropic_alignment.py 2 (Angle sweep) Fig. 3 8.1
-partC_rotation_sweep.py 3 (Rotation) Figs. 57 8.3
-bounds.py 4–6 (OU plateau, rate, directional) Tab. 4, Figs. 48, Tab. 5 8.18.3
-partD_misspecification.py A (Misspecification) Fig. 9 App. B
-partE_eps_exchangeable.py B ($\varepsilon$-exchangeable) Fig. 10 App. B
-The script run_all.py executes all of the above in sequence. The script partF_optimal_batching.py produces additional exploratory figures not included in the paper.
-Shared parameters ($d=10$ experiments).
-Hessian eigenvalues log-spaced in $[1,100]$; $F^{\star}=H^{\star}$ (quadratic Gaussian model).
-Anisotropic gradient covariance $G^{\star}$ drawn as a random SPD matrix with eigenvalues in $[0.5,5]$ (seed =7=7); isotropic baseline $G_{\mathrm{iso}}=\sigma^{2}I$ with $\sigma^{2}=\mathrm{Tr}(G^{\star})/d$ (trace-matched).
-Constant-step experiments use $\eta=0.5/\lambda_{\max}(H)$, $b=64$, $\tau=\eta/b$; decaying-step experiments use $\eta_{t}=1.2/(t+50)$, $b=64$.
-Verification.
-All scripts were executed to completion with zero errors.
-Numerical outputs confirm: (i) log-log slopes $\approx-1$ for the $1/N$ risk decay; (ii) OU stationary variances match Lyapunov predictions to $1\%$; (iii) directional variance ratios align with the eigenstructure of $G^{\star}$; (iv) misspecification experiments show graceful degradation to the sandwich/Godambe prediction.